### v1

# Heterogeneous Intertemporal Treatment Effects via Dynamic Panel Data Models

Abstract
We study the identification and estimation of heterogeneous, intertemporal treatment effects (TE) when potential outcomes depend on past treatments. First, applying a dynamic panel data model to observed outcomes, we show that instrument-based GMM estimators, such as Arellano and Bond (1991), converge to a non-convex (negatively weighted) aggregate of TE plus non-vanishing trends. We then provide restrictions on sequential exchangeability (SE) of treatment and TE heterogeneity that reduce the GMM estimand to a convex (positively weighted) aggregate of TE. Second, we introduce an adjusted inverse-propensity-weighted (IPW) estimator for a new notion of average treatment effect (ATE) over past observed treatments. Third, we show that when potential outcomes are generated by dynamic panel data models with homogeneous TE, such GMM estimators converge to causal parameters (even when SE is generically violated without conditioning on individual fixed effects). Finally, we motivate SE and compare it with parallel trends (PT) in various settings with observational data (when treatments are dynamic, rational choices under learning) or experimental data (when treatments are sequentially randomized).
1 Introduction
Consider a panel data setting in which a researcher observes individual outcomes $Y_{it}$ and treatments $D_{it}$ over multiple periods $t=1,...,T$. At any given time $t$, potential outcomes depend on a full sequence of potential treatments in the past, i.e., $Y_{it}(d^{t})$ where $d^{t}\equiv(d_{s})_{s=0}^{t}$ with $d_{s}\in\{0,1\}$. The objective of causal inference in this context is to learn about certain moments of $Y_{it}(d^{t})$ using the joint distribution of $\{D_{it},Y_{it}\}_{t\leq T}$, where the observed outcomes are determined as $Y_{it}=Y_{it}(D_{i}^{t})$.
This paper considers the question of causal inference through the lens of (linear) panel data models that also include a lagged outcome as a regressor, e.g.,
$Y_{it}=\beta D_{it}+\gamma Y_{it-1}+\theta_{t}+\alpha_{i}+\varepsilon_{it}.$ (1)
There are at least two reasons one might argue for the necessity of such a dynamic panel data model (DPDM). First, treatment may only be randomly assigned conditional on the past history. For example, when estimating the effect of a training program on earnings or employment outcomes, enrollment into the training program may depend on previous earnings and employment (Ashenfelter, 1978). In such cases, the lagged outcome is a necessary control for the exogeneity of treatment. Even though (1) appears to deviate from a static panel data model only by an additional control, the fact that this additional control is a lagged dependent variable raises concerns about the strict exogeneity assumption that is necessary for two-way fixed-effect (TWFE) regressions. Second, outcomes may exhibit state-dependence. For example, past employment may have a causal effect on present employment (Heckman, 1981). In such cases, the lagged outcome also has a causal effect and may furthermore mediate the effects of past treatment. A DPDM such as (1) allows researchers to disentangle the partial effects of $D_{it}$ from the state dependence in $Y_{it-1}$.
One of our main questions is: when and how can we use existing dynamic panel data estimators (DPDE) of observed outcomes, particularly the GMM/2SLS estimators of Anderson and Hsiao (1982) and Arellano and Bond (1991), to shed light on causal effects? These estimators use lagged regressors as instruments to consistently estimate $\beta,\gamma$ in the first-differenced transformation of (1):
$\Delta Y_{it}=\beta\Delta D_{it}+\gamma\Delta Y_{it-1}+\Delta\theta_{t}+\Delta\varepsilon_{it},$ (2)
where $\Delta$ denotes the first-difference operator. To answer this question, we begin by deriving a general causal decomposition of the aforementioned DPDEs (Lemma 1). Implicit in this decomposition, and more generally in our goal of causal inference, is the possibility that (1) and (2) are not structural models of the outcome-generating process, for example because of the possibility of treatment effect heterogeneity. This attempt to link such instrumented first-difference estimators of a reduced-form model to interpretations in terms of aggregated heterogeneous treatment effects (TE) is one contribution of this paper.
The preceding exercise bears resemblance to, but differs qualitatively from, the work of De Chaisemartin and d’Haultfoeuille (2020) and Goodman-Bacon (2021) in decomposing and interpreting TWFE regressions. In a non-dynamic setting where potential outcomes are indexed only by contemporary treatment $Y_{it}(d_{t})$, they showed the TWFE estimand in static panel data models (a special case of (1) with $\gamma=0$) is a weighted average of heterogeneous TE $Y_{it}(1)-Y_{it}(0)$. [^1] The use of TWFE in static panel data models requires that $D_{it}$ be strict exogenous in order to identify $\beta$ and $\gamma$. However, the difference-in-differences approach in De Chaisemartin and d’Haultfoeuille (2020) does not impose that the true data-generating process satisfies the parametric, reduced-form $Y_{it}=\beta D_{it}+\theta_{t}+\alpha_{i}+\varepsilon_{it}$. Their focus is instead on relating the TWFE estimand to aggregated heterogeneous treatment effects under parallel trends; they also consider first-difference regressions under the same assumptions. This literature has also considered the role and the complications of covariates (Heckman, Ichimura, and Todd, 1997; Abadie, 2005; Caetano and Callaway, 2023), as well as proposing alternative estimators under parallel trends (PT) and other assumptions on the outcome or treatment processes, like staggered designs (e.g.,De Chaisemartin and d’Haultfoeuille (2020); Callaway and Sant’Anna (2021); Sun and Abraham (2021)). In comparison, we face different challenges, because the treatment effects in our setting vary with the full sequence of past potential treatments, and the estimator in question uses instruments designed to address the failure of strict exogeneity of $D_{it}$ and $Y_{it-1}$ in the DPDM (1). In the context of papers imposing variants of parallel trends, the setting of de Chaisemartin and D’Haultfoeuille (2024) is most similar in that it also allows for dynamic potential outcomes and non-staggered designs; in contrast, however, our approach is based on a set of assumptions different than parallel trends.
Building on our decomposition in Lemma 1, we derive sufficient conditions under which a 2SLS first-differenced estimator as in Anderson and Hsiao (1982) is guaranteed to be a positively weighted average of heterogeneous TE (Proposition 2). These conditions include the sequential exchangeability assumption of Robins (1986) — which posits that the present potential outcomes are (mean) independent from contemporary treatment, once conditioned on a full history of observed outcomes and treatments — as well as restrictions on the heterogeneity of TE. We extend and strengthen the result by introducing an adjusted inverse-(treament) propensity-weighted (IPW) estimator and show that a sample analogue is consistent for a particular average treatment effect. This new estimator dispenses with the restrictions on heterogeneous TE in Proposition 2; it only requires sequential exchangeability and some mild conditions on treatment assignment. (See Proposition 3.)
By relying on an assumption of sequential exchangeability, our results relate closely to a broad suite of treatment effect estimators collectively referred to as $g$-methods that have been developed for causal inference in this setting [e.g., see Robins (1986, 1997) and Robins and Hernán (2008) for a survey]. Most closely related, Robins (1997) proposes an IPW estimator that identifies $E[Y_{it}(d^{t})]$ using the observed outcomes $Y_{it}$ in stratum $\{D_{i}^{t}=d^{t}\}$; these potential outcomes can in turn be used to identify history-dependent treatment effects such as $E[Y_{it}(d^{t-1},1)-Y_{it}(d^{t-1},0)]$. In contrast, our adjusted IPW estimator identifies an aggregate treatment effect $E[Y_{it}(D_{i}^{t-1},1)-Y_{it}(D_{i}^{t-1},0)]$ using the first-differenced outcomes $\Delta Y_{it}$ among the entire sample. We conjecture that such an aggregation of treatment effects may be useful both from the standpoint of estimation and interpretation. In particular, our target parameter provides a potentially useful summary causal effect in multi-period settings where the number of potential contrasts grows large.
The sequential exchangeability condition has played a central role in our results above (i.e., causal interpretation of dynamic panel data regression, and the adjusted IPW). This naturally raise two additional questions: (a) Can causal inference be conducted using DPDEs without relying on sequential exchangeability? (b) Is sequential exchangeability feasible in observational data where treatments are endogenously determined by forward-looking individuals? We provide conditions under which the answers to both questions are indeed positive.
To answer (a), Propositions 4 and 5 delineate a case where our previous DPDEs identify well-defined causal estimands, and yet sequential exchangeability is generically violated without conditioning on further unobservables. Specifically, we focus on a parametric case where the data-generating process of the latent potential outcomes $Y_{it}(d^{t})$ follows a linear DPDM:
$Y_{it}(d^{t})=\beta^{\ast}d_{t}+\gamma^{\ast}Y_{it-1}(d^{t-1})+\theta_{t}^{\ast}+\alpha_{i}^{\ast}+\varepsilon_{it}^{\ast}(d_{t})\text{.}$ (3)
In this case, the observed outcomes $Y_{it}=Y_{it}(D_{i}^{t})$ are consistent with the reduced form DPDM in (1), where the coefficients coincide with constant causal effects (i.e., $\beta=\beta^{\ast},\gamma=\gamma^{\ast}$ as noted in Observation 2). A further assumption of conditional sequential exogeneity condition on the structural errors $\varepsilon_{it}^{\ast}(\cdot)$ guarantee consistency of our standard DPDEs but not sequential exchangeability, because the exogeneity between contemporaneous outcomes and treatments now only holds after conditioning on unobserved fixed effects. Hence, this allows for endogenous selection on such unobservables.
To investigate (b), we provide, in the context of observational data, conditions under which sequential exchangeability holds. This provides critical guidance for empiricists applying g-methods to observational data, especially when treatment choices are most likely dependent upon past outcomes. For example, when estimating the effect of a training program on earnings or employment outcomes, enrollment into the training program may depend on previous earnings and employment Ashenfelter (1978).
To this end, we consider sequential exchangeability in a model of dynamic choices that serves as an anchor for structurally motivating the conditional and unconditional versions of sequential exchangeability. A particular focus of the model is on the possibility of learning, which is compatible with the nature of sequential exchangeability assumptions that condition (like a learning decision-maker) on the history of observed outcomes. The structural model provides a coherent dynamic decision problem of an economic agent where preferences, beliefs and decision rules are specified and shows the kinds of restrictions that are required, especially on the distribution of prior beliefs, to yield the exogeneity assumptions we use. Thus, our work connects to recent work on identification of learning models in nonparametric settings (e.g., (Bunting, Diegert, and Maurel, 2022)). Also, we connect a conditional independence assumption in the learning model to common parallel trends assumptions and show how sequential random assignment of treatment can be compatible with sequential exogeneity as expected but not parallel trends.
Our work also relates to a broader agenda relating sequential exchangeability and parallel trends (e.g., Renson, Hudgens, Keil, Zivich, and Aiello (2023)) and the corresponding estimators, for example through “bracketing” relationships on the target parameter (Angrist and Pischke, 2008; Ding and Li, 2019). Additionally, Han (2021) studies the identification of dynamic treatment effects beyond sequential exchangeability, namely under a dynamic version of rank similarity. In complementary work, Kim (2023) studies identification of a dynamic panel data model when treatment is staggered and potential outcomes are non-negative, statically indexed, and take a multiplicative form; in this case, he obtains identification results for ratios of treatment effects under sequential exchangeability conditional on persistent heterogeneity.
The rest of the paper is organized as follows. Section 2 introduces the dynamic potential outcome framework. Section 3 studies the causal interpretation of DPDEs in the econometrics literature. Section 4 studies sequential exchangeability in the context of a model of dynamic choice with learning and illustrates the compatibility of parallel trends assumptions with conditional independence restrictions within the learning model and random assignment of treatments. Section 5 concludes.
2 A Dynamic Framework for Potential Outcomes
The researcher observes treatment $D_{it}$ with realized outcomes $Y_{it}$ for individual units $i$ in time periods $t=0,1,\dots,T$. There is a pre-treatment period $t=0$ with an initial realized outcome $Y_{i0}$ and $\Pr\{D_{i0}=0\}=1$. We take $Y_{i0}$ as given and unmodeled, thus avoiding the initial conditions problem (Heckman, 1981). Throughout what follows, we consider models in which potential outcomes in period $t$ are functions only of an individual’s own treatment history through period $t$.
Assumption 1
Potential outcomes for individual unit $i$ in period $t$ are indexed only by $i$’s own treatment through that time period, i.e., $Y_{it}(d_{i}^{t})$.
Assumption 1 rules out spillovers across individual units as well as anticipation of future treatments affecting present potential outcomes. Realized outcomes equal potential outcomes evaluated at realized treatments, $Y_{it}=Y_{it}(D_{i}^{t})$. We adopt notation $Y_{i}^{t}\equiv(Y_{i0},Y_{i1},\dots,Y_{it})$ for realized random (outcome) vectors, $Y_{i}^{t}(d^{t})=(Y_{i0},Y_{i1}(d^{1}),\dots,Y_{it}(d^{t}))$ for potential random vectors, and $Y_{i}^{t}(\cdot)$ for the collection of potential random vectors through period $t$. Throughout, we restrict to realized and potential treatment vectors $D_{i}^{t}$ and $d_{i}^{t}$ with degenerate initial realization $D_{i0}=d_{i0}=0$, which we suppress where convenient.
The main concern that motivates the inclusion of lagged outcomes as controls in a dynamic panel data model (DPDM) for observed outcomes in (1) is that the treatment is (mean) independent of potential outcomes only after conditioning on the observed history of past treatment and outcome. The assumption below formalizes such conditional independence. For subsequent references, we include both a full and mean version of the assumption. We drop individual subscripts $i$ to simplify notation.
Assumption 2 (Sequential Exchangeability)
Let $1\leq s\leq t\leq T$, and let $d^{s}$ and $d^{t}$ denote subvectors of the first $s$ and $t$ terms in $d^{T}\equiv(d_{1},\dots,d_{s},\dots,d_{t},\dots,d_{T})$ respectively.
  1. a.
    Under (full) sequential exchangeability,
    $(Y_{s}(d^{s}),\dots,Y_{T}(d^{T}))\perp D_{s}|Y^{s-1},D^{s-1}=d^{s-1}.$
  2. b.
    Under (mean) sequential exchangeability,
    $E[Y_{t}(d^{t})|Y^{s-1},D^{s-1}=d^{s-1},D_{s}=d_{s}^{\prime}]\quad\text{is invariant in $d_{s}^{\prime}$.}$
Assumption 2 in either form is based on the sequential exchangeability of Robins (1986) in the case where the covariate history consists only of lagged outcomes.222 Other names for Assumption 2b in the literature include “conditional exchangeability”, “sequential ignorability”, and “conditional parallel trends”.
Assumption 2 only restricts a stream of potential outcomes where the first $s-1$ treatment indices conform to the realized history of treatments $D^{t-1}=d^{t-1}$; it imposes conditional independence between present treatment for the entire stream of counterfactual potential outcomes in the future. Mean sequential exchangeability (Assumption 2b) relaxes full sequential exchangeability (Assumption 2a) in two ways. First, the joint independence of treatment with the vector of present and future potential outcomes is replaced with marginal independence of treatment with potential outcomes in each future period. Second, such marginal independence is further relaxed to marginal mean independence. Nonetheless, each version is consistent with the overarching motivation of random assignment of treatment conditional on the observed history.
It is well-known that sequential exchangeability is sufficient for identifying a variety of treatment effects (e.g, Robins and Hernán (2008)). Instead, our focus will be to show how this condition also leads to causal interpretations of alternative estimands in dynamic panel data models for observed outcomes.
Remark 1
A useful observation is that Assumption 2b, which is stated above in terms of the levels of potential outcomes $Y_{t}(d^{t})$, is equivalent to an analogous assumption on the trends in potential outcomes $Y_{t}(d^{t})-Y_{t-1}(d^{t-1})$. This provides insights on comparisons between Assumption 2b and the parallel trends assumption used in the difference-in-differences (DiD) literature. This equivalence result is stated in the next proposition.
Proposition 1
Assumption 2b is equivalent to: for any $t\geq 1$ and $d^{t}\in\{0,1\}^{t}$,
$E[Y_{t}(d^{t})-Y_{t-1}(d^{t-1})|Y^{s-1},D_{s}=d_{s}^{\prime},D^{s-1}=d^{s-1}]\quad\text{is invariant in $d_{s}^{\prime}$ for $t\geq s\geq 1$,}$
where $d^{s}$ is a vector consisting of the first $s$ components in $d^{t}$.
It is straightforward to illustrate Proposition 1 when $T=2$, in which case the alternative trend formulation is: for any $(d_{1},d_{2})\in\{0,1\}^{2}$,
$E[Y_{1}(d_{1})-Y_{0}|Y_{0},D_{1}]$ $=E[Y_{1}(d_{1})-Y_{0}|Y_{0}];$ (4)
$E[Y_{2}(d_{1},d_{2})-Y_{1}(d_{1})|Y_{0},D_{1}]$ $=E[Y_{2}(d_{1},d_{2})-Y_{1}(d_{1})|Y_{0}];$ (5)
$E[Y_{2}(d_{1},d_{2})-Y_{1}(d_{1})|Y_{0},Y_{1},D_{1}=d_{1},D_{2}]$ $=E[Y_{2}(d_{1},d_{2})-Y_{1}(d_{1})|Y_{0},Y_{1},D_{1}=d_{1}].$ (6)
It is clear from $T=2$ that in one way, sequential exchangeability is strong in that it is required to hold for all potential outcomes, both treated and untreated, whereas the parallel trends assumption only restricts the trends in untreated outcomes. On the other hand, sequential exchangeability is required to hold conditional on observed history — including past outcomes — while parallel trends do not condition on past outcomes.
3 Causal Inference via Dynamic Panel Data Methods
3.1 Dynamic Panel Data Models
Canonical dynamic panel data methods are rooted in linear models with lagged outcomes as a regressor. In our context, we consider causal interpretations of a dynamic panel data model (DPDM) for observed outcomes:
$Y_{it}=\beta D_{it}+\gamma Y_{it-1}+\theta_{t}+\alpha_{i}+\varepsilon_{it},$ (7)
where $\gamma,\beta$ are constant parameters, $\alpha_{i}$ are time-invariant individual fixed effects, $\theta_{t}$ are deterministic time trends, and $\varepsilon_{it}$ are time-varying idiosyncratic errors. Specifically, we investigate how the estimated coefficients can be related to the moments of heterogeneous treatment effects and trends in potential outcomes.
Equation (7) is natural for empiricists as a reduced-form model for a series of observed outcomes. For example, a natural interpretation is an extension of a static panel data model in which the researcher also wishes to control for lagged outcomes. A challenge for estimating (7) is the endogeneity of lagged outcomes on the right-hand side, due to the unobservable individual fixed effects $\alpha_{i}$. The ordinary least-squares (OLS) estimators and fixed-effect estimators using either “within” transformation or first-differences are biased and inconsistent when applied to (7) (e.g., (Baltagi, 2021)). Anderson and Hsiao (1982) and Arellano and Bond (1991) proposed consistent estimators for the coefficients in (7), using moments implied by instruments such as the lagged outcomes and covariates.
Next, we turn to the interpretation of such estimators. It is important to emphasize, however, that we do not interpret the DPDM for observed outcomes (7) as a causal model per se. Rather, we consider its implicit interpretation and estimation under non-parametric assumptions on the potential outcome process. This exercise is analogous to the interpretation of TWFE regressions under parallel trend assumptions, e.g., De Chaisemartin and d’Haultfoeuille (2020) and Goodman-Bacon (2021). In Section 3.4 we also consider an analogous parametric model to (7), but as a model of the potential outcome generating process.
3.2 Instrumented First-Difference Estimators
Our goal in this section is to obtain a causal interpretation of estimators for (7) that address the endogeneity of lagged outcomes. First-difference transformation of (7) implies:
$\Delta Y_{it}=\gamma\Delta Y_{it-1}+\beta\Delta D_{it}+\Delta\theta_{t}+\Delta\varepsilon_{it}\ \quad\text{for }2\leq t\leq T\text{,}$ (8)
where $\Delta$ denotes the first-difference operator, e.g., $\Delta Y_{it}\equiv Y_{it}-Y_{it-1}$. Anderson and Hsiao (1982) estimated the coefficients via IV regression, using $Y_{it-2}$ as instruments for $\Delta Y_{it-1}$ when $D_{it}$ is strictly exogenous. Arellano and Bond (1991) proposed a GMM estimator when $D_{it}$ is sequentially exogenous (a.k.a. pre-determined), using earlier lagged outcomes and treatments as instruments for $\Delta Y_{it-1}$ and $\Delta D_{it}$.
We focus henceforth on the causal interpretation of a two-stage least squares (2SLS) estimator for $\beta$ when $T=2$. In this case, the coefficients $\rho$ and $\gamma$ are just identified, with $(Y_{i0},D_{i1})$ being instruments for $(\Delta Y_{i1},\Delta D_{i2})$ in (8), and 2SLS and GMM estimators are numerically equivalent. As before, we drop individual subscripts $i$ for simplicity.
Define history-dependent trends in potential outcomes:
$\delta_{1}\equiv Y_{1}(0)-Y_{0};\quad\delta_{2}(d_{1})\equiv Y_{2}(d_{1},0)-Y_{1}(d_{1})$ (9)
and the treatment effects:
$\tau_{1}\equiv Y_{1}(1)-Y_{1}(0);\quad\tau_{2}(d_{1})\equiv Y_{2}(d_{1},1)-Y_{2}(d_{1},0).$ (10)
The trend of observed outcomes can then be written as:
$\Delta Y_{1}$ $=Y_{1}(D_{1})-Y_{0}=\delta_{1}+D_{1}\tau_{1}\text{;}$ (11)
$\Delta Y_{2}$ $=Y_{2}(D_{1},D_{2})-Y_{1}(D_{1})=\delta_{2}(D_{1})+D_{2}\tau_{2}(D_{1})\text{.}$ (12)
To fix ideas, suppose $Y_{0}$ takes discrete values. Let the first-stage regression be fully saturated, so that the first-stage fitted values of the trend in observed outcomes and treatments converge in probability to $E[\Delta Y_{1}|Y_{0},D_{1}]$ and $E[\Delta D_{2}|Y_{0},D_{1}]$ respectively.
The next lemma invokes the Frisch-Waugh-Lovell Theorem and the decomposition of $\Delta Y_{2}$ in (12) to express the probability limit of the 2SLS estimator for $\beta$ in terms of causal and non-causal components. Let $w(Y_{0},D_{1})$ be the projection errors in the linear projection of $E(\Delta D_{2}|Y_{0},D_{1})$ on $E(\Delta Y_{1}|Y_{0},D_{1})$. Assume $w(Y_{0},D_{1})$ is not degenerate at zero, which is an innocuous, empirically verifiable condition.
Lemma 1
Suppose Assumption 1 holds and $w(Y_{0},D_{1})$ is not degenerate at zero. Let the first-stage regressions be fully saturated. The 2SLS estimator of $\beta$ in (8), which uses $(Y_{0},D_{1})$ as instruments for $(\Delta Y_{1},\Delta D_{2})$, converges in probability to:
$\frac{E[w(Y_{0},D_{1})\Delta Y_{2}]}{E[w(Y_{0},D_{1})^{2}]}=E\left[\frac{w(Y_{0},D_{1})D_{2}}{E[w(Y_{0},D_{1})^{2}]}\times\tau_{2}(D_{1})\right]+E\left[\frac{w(Y_{0},D_{1})}{E[w(Y_{0},D_{1})^{2}]}\times\delta_{2}(D_{1})\right].$ (13)
Lemma 1 clarifies the basis and impediments to causal interpretations of the 2SLS estimator with no other conditions than the parametrization of potential outcomes (Assumption 1). It does not invoke sequential exchangeability in Assumption 2. The first term on the right in (13) is a weighted aggregate of treatment effects in period 2 among the treated ($D_{2}=1$), yet with no guarantee that the combination is convex or even positively weighted.333 This is analogous to a point by De Chaisemartin and d’Haultfoeuille (2020) and Goodman-Bacon (2021) in the causal interpretation of two-way fixed-effect (TWFE) regressions in static panel data models. The second term, if non-zero, is an additional confounder to causal interpretation in terms of treatment effects $\tau_{2}(d_{1})$.
To endow the 2SLS estimand in (13) with a clearer causal interpretation, we require further restrictions on heterogeneous treatment effects in addition to those on selection in Assumption 2.
Assumption 3 (Limited Heterogeneity of Trends and Treatment Effects)
For all $d_{1},d_{1}^{\prime}=0,1$:
$E[\delta_{1}|Y_{0}]$ $=E[\delta_{1}],$ (14)
$E[\delta_{2}(d_{1})|Y_{0}]$ $=E[\delta_{2}(d_{1})],$ (15)
$E[\tau_{1}|Y_{0}]$ $=E[\tau_{1}]\neq 0,$ (16)
$E[\tau_{2}(d_{1})|Y_{0},Y_{1}(d_{1}^{\prime})]$ $=E[\tau_{2}(d_{1})].$ (17)
Note first that each of the conditions in Assumption 3 would be satisfied in the special case where treatment effects $\tau_{t}(\cdot)$ and untreated trends $\delta_{t}(\cdot)$ are homogeneous in the population. Under Assumption 3, the next proposition shows that the 2SLS first-difference estimator recovers a convex combination of history-dependent treatment effects.
Proposition 2
Suppose Assumptions 1, 2a, 3 hold, and $w(Y_{0},D_{1})$ is not degenerate at zero. The 2SLS estimator of $\beta$ in (8) converges in probability to a convex combination of treatment effects:
$E\left[\frac{w(Y_{0},D_{1})^{2}}{E[w(Y_{0},D_{1})^{2}]}\times\tau_{2}(D_{1})\right],$ (18)
where the projection errors $w(Y_{0},D_{1})$ simplify to:
$w(Y_{0},D_{1})=E[D_{2}|Y_{0},D_{1}]-E[D_{2}|D_{1}].$ (19)
This proposition confirms that an intuitive causal interpretation of the 2SLS estimand for the coefficient of $\Delta D_{2}$ exists under sufficiently restrictive conditions. A proof is presented in Appendix A. Intuitively, the proof proceeds in the three steps. First, derive the expression in (19) for the projection errors. Next, eliminate the confounding (second) term in (13). Finally, establish the convex weights in (18).
A technical note worth mentioning is that only the mean version of sequential exchangeability (Assumption 2b) is invoked in the first and second steps of the proof; in the third step, the full version (Assumption 2a) is required in period $t=1$ to separate restrictions on selection from restrictions on heterogeneity.
Proposition 2 complements the existing literature. It is well-known that the average treatment effect $E[\tau(d_{1})]$ is identified for $d_{1}\in\{0,1\}$ under sequential exchangeability in Assumption (2); see, e.g., an application of the $g$-formula in (Robins and Hernán, 2008). In comparison, our contribution is to show how to use an alternative 2SLS approach to estimate a different causal effect parameter, which is a weighted aggregate of treatment effects where the weights can be recovered from observable moments.
3.3 An Adjusted Inverse-Propensity-Weighted Estimator
The result in Proposition 2 naturally raises a new question: can we modify our approach to construct new estimators for other averages of causal effects, without restricting the heterogeneity of treatment effects as in Assumption 3?
Our next result shows how to use an adjusted inverse-propensity-weighted (IPW) method to estimate an average treatment effect (ATE) upon transforming $\Delta Y_{2}$. The basis for this transformation is the same decomposition (12) invoked in the preceding subsection 3.2. In our dynamic potential outcome setting, treatment effects are also indexed by past treatment $\tau_{2}(d_{1})$. Thus, the ATE in period 2 that we target also averages over the observed past treatment $D_{1}$. This new (to the best of our knowledge) estimator does not invoke the restrictions on heterogeneous treatment effects and trends in Assumption 3 or the full version of sequential exchangeability of Assumption 2a; besides Assumptions 1 and 2b, it only requires a standard support assumption on treatment paths.
Assumption 4 (Full Support of Treatment)
$E[D_{2}|Y^{1},D_{1}]\in(0,1)$.
Under Assumption 4, define:
$\widetilde{\Delta Y_{2}}\equiv\frac{\Delta Y_{2}-E[\Delta Y_{2}|Y^{1},D_{1},D_{2}=0]}{E[D_{2}|Y^{1},D_{1}]}.$
The next proposition establishes that the mean of $\widetilde{\Delta Y_{2}}$ identifies the ATE in period 2, given the realized treatments in period 1.
Proposition 3
Suppose Assumptions 1, 2b and 4 hold. Then $E[\widetilde{\Delta Y_{2}}]=E[\tau_{2}(D_{1})]$.
There are two keys to the result. First, the mean of the trend $\delta_{2}(D_{1})$ in the decomposition of $\Delta Y_{2}$ in (12) can be identified and removed in the new estimand $E[\widetilde{\Delta Y_{2}}]$ under Assumption 2. Second, the use of inverse propensity weights allows us to attain a more interpretable causal parameter (ATE) without the weights in Proposition 2.
By Proposition 3, we can estimate the ATE, $E[\tau_{2}(D_{1})]$, using sample analogs. Let $\widehat{M}_{2i}$ and $\widehat{M}_{1i}$ be consistent (e.g., kernel-based, Nadaraya-Watson) estimators for
$M_{2i}\equiv E[\Delta Y_{2}|D_{2}=0,Y^{1}=(Y_{i0},Y_{i1}),D_{1}=D_{i1}]$
and
$M_{1i}\equiv E[D_{2}|Y^{1}=(Y_{i0},Y_{i1}),D_{1}=D_{i1}]$
evaluated at any realized $(Y_{i0},Y_{i1},D_{i1})$. The analog estimator for $\mu_{\tau_{2}}\equiv E[\tau_{2}(D_{1})]$ is:
$\widehat{\mu}_{\tau_{2}}\equiv\frac{1}{n}\sum\nolimits_{i=1}^{n}\left(\frac{\Delta Y_{i2}-\widehat{M}_{2i}}{\widehat{M}_{1i}}\right)\text{.}$ (20)
In Appendix B, we show this adjusted IPW estimator $\widehat{\mu}_{\tau_{2}}$ is consistent for $\mu_{\tau_{2}}$.
Robins (2000) introduced an IPW estimator for $E[Y_{t}(d^{t})]$ for any $d^{t}\in\{0,1\}^{t}$ in marginal structural models. Despite being motivated by his original IPW estimator, our adjusted IPW estimator targets a different causal parameter. Specifically, our estimand is $E[\tau_{2}(D_{1})]=E[Y_{2}(D_{1},1)-Y_{2}(D_{1},0)]$, while that of Robins’ IPW estimator is $E[Y_{2}(d_{1},d_{2})]$. Two further distinctions of our estimator are that it is based on the first-differenced outcome $\Delta Y_{2}$ and that it pools all observations in estimation. In addition to pooling in estimation, our target parameter provides a potentially useful summary causal effect in multi-period settings where the number of potential contrasts grows large. That is, $E[Y_{t}(D^{t-1},1)-Y_{t}(D^{t-1},0)]$ can be thought of as a weighted average of the $2^{t-1}$ average effects $E[Y_{t}(d^{t-1},1)-Y_{t}(d^{t-1},0)]$ that are in principle identified under sequential exchangeability.
Alternatively, we can construct a numerically equivalent 2SLS estimator for $E[\tau_{2}(D_{1})]$ after transforming $\widetilde{\Delta Y_{2}}$ using the weights $W(Y_{0},D_{1})$. To see this, let
$\Delta Y_{2}^{\dagger}\equiv\widetilde{\Delta Y_{2}}\times\frac{E[W^{2}]}{W},$
and run a transformed 2SLS regression on $\Delta Y_{2}^{\dagger}$. First, regress the endogenous regressors in (8), $\Delta Y_{1}$ and $\Delta D_{2}$, on a fully saturated set of instruments $(Y_{0},D_{1})$ to obtain the fitted values. Then regress $\Delta Y_{2}^{\dagger}$ on the fitted values (and an intercept). It is clear from (13) in Lemma 1 that, provided the projection errors $W(Y_{0},D_{1})$ are nonzero almost surely, such an estimand is equal to $E[\widetilde{\Delta Y_{2}}]$, and therefore $E[\tau_{2}(D_{1})]$ by Proposition 3. In practice, we can implement this transformed 2SLS by first estimating $\Delta Y_{2i}^{\dagger}$ as the product of each individual summand in $\widehat{\mu}_{\tau_{2}}$ and its estimated weights $\frac{1}{n}\sum_{j}\widehat{W}_{j}^{2}/\widehat{W}_{i}$, where $\widehat{W}_{i}$ is the sample analog estimator for the projection errors $w(Y_{0i},D_{1i})$.
3.4 Settings Without Sequential Exchangeability
Sequential exchangeability (Assumption 2) has played a central role in the preceding results of Subsections 3.2 and 3.3. It seems natural to ask whether this assumption is necessary for causal inference in our setting. Indeed, the identification of linear dynamic panel data models with persistent unobserved heterogeneity as in (1) suggests it is not necessary per se, at least if replaced with other (e.g., parametric) assumptions. The goal of this subsection is to establish this in our potential outcome framework and show that for a class of dynamic data-generating processes, standard GMM-based DPDEs recover causal estimands even though sequential exchangeability is generically violated without conditioning on unobservables.
Consider a specialized case in which the series of potential outcomes are structurally determined by a linear dynamic panel data model.
Assumption 5
Potential outcomes follow a linear, homogeneous, AR(1) model:
$Y_{it}(d^{t})=\beta^{*}d_{t}+\gamma^{*}Y_{it-1}(d^{t-1})+\theta_{t}^{*}+\alpha_{i}^{*}+\varepsilon_{it}^{*}(d_{t})\text{ for }t\geq 1,$ (21)
with initial condition $Y_{i0}(d^{0})\equiv Y_{i0}$, and where $\beta^{*},\gamma^{*}$, and $\theta_{t}^{*}$ are constant coefficients, $\alpha_{i}^{*}$ are individual-specific fixed effects, and $\varepsilon_{it}^{*}(d_{t})$ are time-varying noises in potential outcomes.
Unlike $\gamma$ in the dynamic panel data model for observed outcomes in (7), the coefficient $\gamma^{*}$ in this model for potential outcomes in (21) has an inherently causal interpretation: it measures how changes in present potential outcomes depend on counterfactual changes in lagged (past) potential outcomes, which depend recursively on the full counterfactual treatment history $d^{t-1}$. Note that the structure of Assumption 5 also implies the potential outcome parametrization of Assumption 1.
The following observation connects the linear dynamic panel data model for potential outcomes to the reduced-form model (7) for observed outcomes.
Observation 1
Suppose the data-generating process for potential outcomes satisfies Assumption 5, and the observed outcomes are $Y_{it}=Y_{it}(D_{it})$. Then the observed sequence $(D_{it},Y_{it})$ satisfies (7) with $(\gamma,\beta,\alpha_{i},\theta_{t})=(\gamma^{*},\beta^{*},\alpha_{i}^{*},\theta_{t}^{*})$ and $\varepsilon_{it}=\varepsilon_{it}^{*}(D_{it})$.
This connection is useful because conditions for consistent estimation of the reduced-form, realized outcome model (7) are well-established. We focus on the Arellano-Bond (1991) GMM estimator, which requires the following set of assumptions on the reduced-form errors $\varepsilon_{it}$:444 Condition (24) is implicit in the moment condition in Arellano and Bond (1991) (Eq. (2) in Sec. 2).
$E(\varepsilon_{is}\varepsilon_{it})=0\text{ for all }t\neq s;$ (22)
$E(D_{is}\varepsilon_{it})=0\text{ for all }s\leq t;$ (23)
$E(\Delta\varepsilon_{it}\alpha_{i})=E(\Delta\varepsilon_{it}Y_{i0})=0\text{ for }t\geq 2.$ (24)
In turn, a sufficient condition on the structural errors $\varepsilon_{it}^{*}(\cdot)$ combines sequential exogeneity and serial uncorrelation, conditional on individual fixed effects. For simplicity, we term this strong structural sequential exogeneity. Let $\varepsilon_{i}^{*t}(d^{t})\equiv(\varepsilon_{i1}^{*}(d_{1}),\dots,\varepsilon_{it}^{*}(d_{t}))$ for $t\geq 1$ and $\varepsilon_{i}^{*t}(\cdot)$ the collection of potential error vectors, with $\varepsilon^{*0}_{i}(\cdot)$ a vacuous null vector.
Assumption 6 (Strong Structural Sequential Exogeneity)
The potential error terms defined under Assumption 5 satisfy:
$E[\varepsilon_{it}^{*}(\cdot)|D_{i}^{t},\varepsilon^{*t-1}_{i}(\cdot),Y_{i0},\alpha_{i}^{*}]=0\text{ for all }t\geq 1.$ (25)
Sufficiency for the Arellano-Bond econometric assumptions is formalized in the following proposition:
Proposition 4
Under Assumptions 5 and 6, the Arellano-Bond econometric assumptions (22), (23), and (24) are satisfied.
The next proposition shows that, within the linear dynamic model for potential outcomes (Assumption 5), strong structural sequential exogeneity (Assumption 6) implies a “conditional” version of mean sequential exchangeability that controls for the fixed effects.
Proposition 5
Under Assumptions 5 and 6, mean sequential exchangeability holds conditional on the fixed effect:
$E[Y_{it}(d^{t})|D_{is}=d_{s}^{\prime},Y_{i}^{s-1},D_{i}^{s-1}=d^{s-1},\alpha_{i}^{*}]\quad\text{is invariant in $d_{s}^{\prime}$ for $t\geq s\geq 1$}$ (26)
for all $d^{t}\in\{0,1\}^{t}$, and where $d^{s-1}$ is a subvector of the first $s-1$ terms of $d^{t}$.
A corollary of Proposition 5 is that the mean sequential exchangeability in Assumption 2b would not hold in general, if the conditional mean in (26) is non-degenerate in individual fixed effect $\alpha_{i}^{*}$, and if the distribution of $\alpha_{i}^{*}$ conditional on history $Y_{i}^{s-1},D_{i}^{s}$ is not invariant in the most recent treatment $D_{is}$. In other words, the unconditional mean sequential exchangeability condition on potential outcomes would not hold in the context of the linear, structural model of Assumption 5 if both potential outcomes and contemporary treatments depend on individual fixed effects in non-trivial ways.
Summarizing, under Assumptions 5 and 6, standard GMM-based DPDEs recover causal estimands (Proposition 4) even though sequential exchangeability is generically violated without conditioning on unobservables (Proposition 5). It is important to reiterate that our results in Section 3.2 and 3.3 do not require such a specific structural model of potential outcomes or econometric restrictions on the errors.
4 Evaluating Sequential Exchangeability
In this section, we assess the plausibility of sequential exchangeability (SE) in Assumption 2, which has played a central role in our results and discussion in Section 3. This condition amounts to a form of “selection on observables”, as it posits (mean or full) independence between potential outcomes and contemporary treatment conditional on observed history of past outcomes and treatments.
Our objective is to provide guidelines for empirical researchers to decide when and how such restrictions on selection-on-observables can be justified in observational or experimental data. Naturally, justifying selection on observables in this dynamic setting would require restrictions on how the sequence of treatments are determined.
In Section 4.1 and 4.2, we showcase the nature of these restrictions within a structural model where treatments are chosen by rational, forward-looking individual units with learning, and compare them with the parallel trends condition in the DiD literature. In Section 4.3, we show that sequential exchangeability holds in an experimental design of sequentially randomized treatment (where the treatment in period $t$ is randomized conditioning on the history of past treatments and outcomes); in contrast, the parallel trends conditions in general would not hold in such a case.
4.1 A Structural Model of Treatment Choices with Learning
We focus on a structural model of dynamic treatment choices that enables individual decision-makers to be both forward-looking and capable of learning. Learning is an interesting possibility in dynamic choice which is potentially compatible with sequential exchangeability assumptions because of the latter’s dependence on the history of observed outcomes. Unlike Section 3.4, the model studied in this section refrains from any parametric specification for the potential outcomes; instead, the model emphasizes scenarios in which forward-looking individuals make rational treatment decisions, continually learning from previous treatments and outcomes. To simplify the notation in this section, we omit the individual subscripts $i$.
In each period $1\leq t\leq T$, the potential outcome is determined by:
$Y_{t}(d^{t})=\lambda_{t}(d^{t},Y^{t-1}(d^{t-1}),\alpha(d_{t}),\varepsilon_{t}(d_{t}))\text{,}$ (27)
where $Y^{s}(d^{s})=(Y_{0},Y_{1}(d^{1}),\dots Y_{s}(d^{s}))$ denotes a vector of counterfactual outcomes, with convention $Y^{0}(d^{0})=Y_{0}$. The initial condition $Y_{0}$ is allowed to be correlated with subsequent potential outcomes $Y_{t}(d^{t})$ for $t\geq 1$. We refer to $\alpha(\cdot)\equiv(\alpha(1),\alpha(0))$ as (stochastic) unobserved individual heterogeneity, or fixed effects, and $\varepsilon_{t}(\cdot)$ as potential, transitory noises affecting potential outcomes at time $t$. Individual decision makers do not observe $\alpha(\cdot)$ or $\varepsilon_{t}(\cdot)$, but know the deterministic function $\lambda_{t}(\cdot)$ as well as the distribution of $\varepsilon_{t}(\cdot)$ for all $t$555 Note that the structural dependence on the lagged outcome generalizes indexing of the full decision history because the outcome-generating function $\lambda_{t}(\cdot)$ is deterministic and depends only on present noise $\varepsilon_{t}(d_{t})$; in particular, the lagged outcome permits some dependence on the past noise vector $\varepsilon^{t-1}(d^{t-1})$.
Each individual holds a prior belief for $\alpha(\cdot)$ at $t=1$, and uses incremental information to update their beliefs for $\alpha(\cdot)$ at the beginning of each period $t\geq 2$. Specifically, the individual’s prior over $\alpha(\cdot)$ at the start of $t=1$ is $\Psi_{1}(\cdot;\psi_{1})$, which is fully characterized by a random vector $\psi_{1}$. Let $\psi_{1}\equiv\varphi_{1}(H_{0},\xi_{0})$, where $H_{0}=Y_{0}$ denotes the observed history and $\xi_{0}$the unobserved individual factors in period 0 that determine the initial prior in period 1. In period $t\geq 2$, the posterior for $\alpha(\cdot)$ is $\Psi_{t}(\cdot;\psi_{t})$, which is completely characterized by an updated vector $\psi_{t}=\varphi_{t}(\psi_{t-1},H_{t-1},\xi_{t-1})$, where $H_{t}\equiv(D_{t},Y_{t})$ is the incremental information (observed by the researcher) that affects the individual’s posterior, and $\xi_{t}(d^{t})$ is the incremental information (unobserved by the researcher) from choice $d_{t}$ in period $t$, with $\xi_{t}=\xi_{t}(D^{t})$. Let $H^{t}\equiv(H_{0},\dots,H_{t})$ and $\xi^{t}=(\xi_{0},\dots,\xi_{t})$ denote the realized vectors and $\xi^{t}(\cdot)$ the full potential vector through period $t$. Using recursive substitution and with a slight abuse of notation, we can write: [^6] With $\Psi_{t}$ and $\varphi_{t}$ indexed by $t$, we allow for non-conjugate beliefs. Also, $\varphi_{t}$ may depend on unknown parameters, such as the variance of $\varepsilon_{t}(\cdot)$ in a normal learning model.
$\psi_{t}=\varphi_{t}(H^{t-1},\xi^{t-1}).$
To reiterate, we assume that $\psi_{t}$ is a sufficient statistic for the posterior of $\alpha(\cdot)$ at time $t$ given the outcome-relevant history $(H^{t-1},\xi^{t-1})$ observed by the decision-maker.
Let $\mathcal{D}^{T}$ denote the set of possible sequences of choices; note that, in contrast to preceding sections, we do not assume that choices in each period are binary. Given a choice history $D^{t-1}=d^{t-1}$, an individual’s instantaneous, ex post utility from choosing $d_{t}$ at time $t$ is:
$u_{t}\left(d_{t},Y_{t}(d^{t}),X_{t}\right)+\eta_{t}(d_{t})$
where $X_{t}$ are additional states (apart from outcomes $Y_{t}$), and $\eta_{t}(\cdot)$ are potential shocks. Individuals observe $(X_{t},\eta_{t}(\cdot))$ and know the deterministic function $u_{t}(\cdot)$ for all $t\geq 1$. We consider the case where $X_{t}$ consists of persistent or serially correlated heterogeneities that are not reported in the data; conceptually, it is not hard to extend to the cases where some components in $X_{t}$are also reported in the sample. To allow the transition of states to depend on choice, it is useful to also define the potential state $X_{t}(d^{t-1})$ that would have occurred in period $t$ under choice history $d^{t-1}$; as usual, $X^{T}=X^{T}(D^{T-1})$.
Summarizing, in each period $t$, the individual observes $(\eta_{t}(\cdot),X_{t})$ before choosing $D_{t}$, and observes $(H_{t},\xi_{t})$ after choosing $D_{t}$. The individual knows the full joint distribution of potential random variables across periods, but not the realizations except as previously specified. For example, the individual does not generally observe the realization of $\alpha(\cdot)$ or $\varepsilon_{t}(\cdot)$ in the outcome-generating process. Define $\zeta^{t}(d^{t-1})\equiv(\xi^{t-1}(d^{t-1}),X^{t}(d^{t-1}),\eta^{t}(\cdot))$ as the potential accumulation of the individual’s information in period $t$ given potential choice history $d^{t-1}$, which is not reported in the data. With minor loss of generality, we consider choices $D_{t}$ that are a function of the individual’s observed and unobserved information:
$D_{t}=\phi_{t}(H^{t-1},\zeta^{t})$ (28)
Unlike time-homogeneous, Markovian choice rules in a typical dynamic discrete choice model, e.g., Rust (1987), the (rational) rule of treatment choices in this section needs to be neither Markovian nor time-homogeneous. This is because the individual unit needs to update its beliefs in each period using the full history of outcomes and endogenous choices in the past.
The key sufficient condition for sequential exchangeability is that the unobserved information for any counterfactual choice sequence is essentially uninformative about the outcome process, conditional on the initial observed outcome:
$\zeta^{T}(d^{T-1})\perp(\alpha(\cdot),\varepsilon^{T}(d^{T}))\mid Y_{0}\quad\text{for all $d^{T}\in\mathcal{D}^{T}$.}$ (29)
For clarity at the expense of some generality, this nests several more primitive assumptions based on the unobserved vector $\zeta^{t}=(\xi^{t-1},X^{t},\eta^{t}(\cdot))$. First, there is no meaningful unobserved heterogeneity in prior beliefs through $\xi_{0}$, though prior beliefs may vary systematically with $Y_{0}$. Second, there is no meaningful unobserved heterogeneity in incremental learning through $\xi_{t}$ for $t\geq 1$, though the belief parameter $\psi_{t}$ may evolve systematically based on the observed history $H^{t-1}$. Third, the state transition process $X^{T}$ is uninformative about the outcome process, so that there is no learning based on the unobserved evolution of states. Finally, independence of the shocks in instantaneous ex post utilities $\eta_{t}(\cdot)$ is standard. In summary, (29) permits systematic learning based only on observables. In that case, (mean) sequential exchangeability is satisfied.
Proposition 6
Suppose (29) holds in the structural choice model with learning. Then sequential exchangeability (Assumption 2) is satisfied.
The intuition underlying this result is straightforward. By definition (27), for any sequence of potential treatment $d^{t}$, the history of potential outcomes $Y^{t}(d^{t})$ is a function of initial condition $Y_{0}$, fixed effects $\alpha(\cdot)$ and noises $\varepsilon^{t}(d^{t})$ only. Meanwhile, conditional on an observed history $D^{s-1}=d^{s=1}$, $Y^{s-1}(d^{s-1})=y^{s-1}$, any event $D_{s}=d_{s}$ is defined over the sample space of choice unobservables $\zeta^{s}(d^{s-1})$, which is conditionally independent from the outcome unobservables, and consequently the sequence of potential outcomes. This means the confounders for the effect of $D_{s}$ on subsequent potential outcomes $Y_{t}(d^{s},\cdot)$ are also controlled for in this case under sequential exchangeability.
Our interpretation of this result is twofold. First, it highlights the limitations of sequential exchangeability in a dynamic choice context. In this sense, the result is more about specifying the channels of unobserved choice heterogeneity that must be ruled out, than about what is allowed. Thus, even as the choice model recovers the possibility of fixed effects in the outcome model without violating sequential exchangeability, it imposes significant restrictions on unobserved heterogeneity in choice.
At the same time, viewed in a more positive light, Proposition 6 provides sufficient conditions under which sequential exchangeability can be used to identify the causal effects when treatment choices are made by forward-looking individuals capable of learning. In this case, sequential exchangeability provides an alternative suite of methods for recovering treatment effects in the presence of learning under (limited) unobserved heterogeneity, complementary to the pure learning model of Bunting, Diegert, and Maurel (2022).
4.2 Comparison With Parallel Trends: Under Learning
We next compare sequential exchangeability and parallel trends within the context of the structural learning model in Section 4.1. It is most illustrative to make such a comparison in a simplified version of the model with no additional states $X_{t}(d^{t-1})$ and $T=2$. To reiterate, in such a simplified case, the potential outcomes are:
$Y_{1}(d_{1})$ $=$ $\lambda_{1}(d_{1},Y_{0},\alpha(d_{1}),\varepsilon_{1}(d_{1}))\text{;}$ (30)
$Y_{2}(d_{1},d_{2})$ $=$ $\lambda_{2}(d_{1},d_{2},Y_{0},Y_{1}(d_{1}),\alpha(d_{2}),\varepsilon_{2}(d_{2}))\text{,}$
and the rules of treatment choices are:
$D_{1}$ $=$ $\phi_{1}(Y_{0},\xi_{0},\eta_{1}(\cdot))\text{;}$ (31)
$D_{2}$ $=$ $\phi_{2}(Y_{0},Y_{1},D_{1},\xi^{1}(D_{1}),\eta_{2}(\cdot))\text{,}$
where $\xi^{1}(d_{1})\equiv(\xi_{0},\xi_{1}(d_{1}))$. [^7] Recall $D_{0}=0$ w.p.1 and is therefore dropped from $\xi_{0}(D_{0}),\xi_{1}(D_{0},D_{1})$ for simplicity.
In the current setting with dynamic treatment effects, the parallel trends assumption (conditional on initial outcomes $Y_{0}$) requires:
$E[Y_{2}(0,0)-Y_{1}(0)|D_{1},D_{2},Y_{0}]\text{ is invariant in }(D_{1},D_{2}).$ (32)
See, e.g., de Chaisemartin and D’Haultfoeuille (2024). The next proposition shows, under a conditional independence condition, sequential exchangeability holds whereas parallel trends would not hold in general.
Proposition 7
Consider the model Section 4.1 with no states $X_{t}(d^{t-1})$ and $T=2$. If
$(\alpha(\cdot),\varepsilon^{2}(d^{2}))\perp(\xi^{1}(d_{1}),\eta^{2}(\cdot))\mid Y_{0}\text{ for all }d^{2}\equiv(d_{1},d_{2})\text{,}$ (33)
then sequential exchangeability in Assumption 2b holds while $E[Y_{2}(0,0)-Y_{1}(0)|D_{1},D_{2},Y_{0}]$ is in general a non-degenerate function of $(D_{1},D_{2})$.
The conditional independence in (33) specializes (29) when there are no additional states $X_{t}(d^{t-1})$ and $T=2$. This condition posits the unobserved states influencing posterior formation are independent from the individual fixed effects, after controlling for the initial outcome $Y_{0}$; it does not further restrict the learning process.
For example, consider a simplistic case where $\varepsilon^{2}(d^{2}),\eta^{2}(\cdot)$ are idiosyncratic noises independent from all other variables, and where $\xi^{t}(d^{t})=\xi_{0}$. That is, the evolution of posterior beliefs only depends on an initial shifter of the prior $\xi_{0}$ and observed history $(D^{t-1},Y^{t-1})$, but no incremental, time-varying $\xi_{s}(d_{s})$ for $s\geq 1$. Then (33) holds when $\alpha(\cdot)$ is independent from the initial prior shifter $\xi_{0}$ given $Y_{0}$.
The result in Proposition 7 is intuitive. With forward-looking and learning, the treatment choices $D^{2}=(D_{1},D_{2})$ are a non-trivial function of a realized history of observed outcomes $Y_{0},Y_{1}(D_{1})$. Hence conditioning on different realizations of $D^{2}$ amounts to conditioning on different events defined over the sample space of $(Y_{0},\alpha(\cdot),\varepsilon_{1}(\cdot),\xi^{1}(\cdot),\eta^{2}(\cdot))$. Because $Y_{2}(0,0)-Y_{1}(0)$ is a function of $(Y_{0},\alpha(\cdot),\varepsilon^{2}(\cdot))$, its distribution (and mean) generally vary with the conditioning event defined by $D^{2}$. This non-degeneracy holds in general, even after controlling for the realization in $Y_{0}$.
4.3 Sequentially Randomized Treatments
We conclude this section with yet another intuitive and important case of treatment assignment mechanism, in which sequential exchangeability holds but parallel trends do not. This is the case where the treatment assignments are sequentially randomized conditional on earlier history of treatments and outcomes.
Suppose the treatments are sequentially randomized as follows:
$D_{1}\perp(Y_{1}(\cdot),Y_{2}(\cdot,\cdot))$$\mid$ $Y_{0},$ (34)
$D_{2}\perp Y_{2}(D_{1},\cdot)$$\mid$ $Y_{1},Y_{0},D_{1}\text{.}$ (35)
Such a treatment assignment mechanism is relevant especially in experimental designs where a researcher may be able to assign randomized treatment in each period after controlling for observed outcomes up to that period.
For simplicity, let the initial outcome $Y_{0}$ be degenerate at a constant value and suppress it in all conditioning events below. Mean sequential exchangeability in Assumption 2b holds under (34) and (35), because
$E[Y_{2}(d_{1},d_{2})|D_{2},Y_{1},D_{1}=d_{1}]$ $=$ $E[Y_{2}(d_{1},d_{2})|Y_{1},D_{1}=d_{1}]\text{,}$
$E[Y_{2}(d_{1},d_{2})|D_{1}]$ $=$ $E[Y_{2}(d_{1},d_{2})]\text{,}$
$E[Y_{1}(d_{1})|D_{1}]$ $=$ $E[Y_{1}(d_{1})]\text{,}$
for all $d_{1},d_{2}\in\{0,1\}$.
In contrast, the parallel trends condition in (32) would fail in general under (34) and (35). To see this, note:
$E[Y_{2}(0,0)-Y_{1}(0)|D_{2}=d_{2},D_{1}=0]$
[…]
where $\mu_{2}(s)\equiv E[Y_{2}(0,0)|Y_{1}(0)=s,D_{1}=0]$ is a function of $s$ that does not vary with $d_{2}$, and the second equality is due to (35). Besides, by construction,
[…]
where $p(s)\equiv\Pr\{D_{2}=1|Y_{1}(0)=s,D_{1}=0\}$ and $G(s)\equiv\Pr\{Y_{1}(0)\leq s|D_{1}=0\}=\Pr\{Y_{1}(0)\leq s\}$. Thus, when $p(\cdot)$ is non-degenerate and non-linear, $F_{Y_{1}(0)}(y|D_{2}=d_{2},D_{1}=0)$ generally varies with $d_{2}$. As a result, (32) would not hold in general.
Note that (32) holds in a special case where $p(\cdot)$ is degenerate, e.g., $D_{2}$ is randomly assigned independent of past $Y_{1}(\cdot)$. In such a case, $F_{Y_{1}(0)}(\cdot|D_{2}=d_{2},D_{1}=0)=G(\cdot)$ does not vary with $d_{2}$ and (32) holds.
5 Conclusion
In this note we considered the causal interpretation of dynamic panel data estimators in models of potential outcomes that can depend on past treatments. We derived sufficient conditions for instrumented first-difference estimators to be causally interpretable and proposed a new adjusted inverse probability of treatment estimator that uses the first-differenced outcome to recover a history-dependent average treatment effect.
An underlying assumption for (most of) these results was sequential exchangeability. In addition to investigating the plausibility of this assumption in a structural model of dynamic choice, we showed that this assumption is not necessary per se for dynamic panel data estimators to recover causal estimands in our setting. A natural avenue for future work is to consider (partial) identification when sequential exchangeability is relaxed or holds only conditional on a set of unobservables.
Appendix A: Proofs
Proof of Lemma 1. Let $\mathcal{Y}$ denote the (discrete) support of $Y_{0}$. Let $Z_{i}\equiv\{\mathbb{I}[(D_{i1},Y_{i0})=(d_{1},y_{0})]\}_{d_{1}\in\{0,1\},y_{0}\in\mathcal{Y}}$ be a row-vector of $2\times\#\mathcal{Y}$ (dummy) instrumental variables. The 2SLS estimator for the coefficient of $\Delta D_{i2}$ in (8) is constructed as follows. In the (fully saturated) first stage, obtain the fitted values of $\Delta Y_{i1}$ (and $\Delta D_{i2}$) from their regression on $Z_{i}$. These fitted values are $Z_{i}\hat{A}$ (and $Z_{i}\hat{B}$, respectively), where
$\hat{A}\equiv\left(\frac{1}{n}\sum\nolimits_{i}Z_{i}^{\prime}Z_{i}\right)^{-1}\left[\frac{1}{n}\sum\nolimits_{i}Z_{i}^{\prime}\left(\Delta Y_{i1}\right)\right]\text{ and }\hat{B}\equiv\left(\frac{1}{n}\sum\nolimits_{i}Z_{i}^{\prime}Z_{i}\right)^{-1}\left[\frac{1}{n}\sum\nolimits_{i}Z_{i}^{\prime}\left(\Delta D_{i2}\right)\right]\text{.}$
By the property of ordinary least squares estimators,
[…]
Formally, $A$ (and $B$) are $(2\times\#\mathcal{Y)}$-by-$1$ column vectors with each component being the conditional mean of $\Delta Y_{i1}$ (and $\Delta D_{i2}$) given $(D_{i1},Y_{i0})=(d_{1},y_{0})$ for some $d_{1}\in\{0,1\}$, $y_{0}\in\mathcal{Y}$. This means $Z_{i}A=E[\Delta Y_{i1}|D_{i1},Y_{i0}]$ and $Z_{i}B=E[\Delta D_{i2}|D_{i1},Y_{i0}]$ by construction.
In the second stage, regress $\Delta Y_{i2}$ on an intercept and the first-stage fitted values $Z_{i}\hat{A}$, $Z_{i}\hat{B}$. Our goal is to investigate the causal interpretation of the slope estimate for $Z_{i}\hat{B}$ in this second-stage regression. By the Frisch-Waugh-Lovell Theorem, this slope estimate is numerically equivalent to
$\frac{\frac{1}{n}\sum_{i}\hat{W}_{i}\Delta Y_{i2}}{\frac{1}{n}\sum_{i}\hat{W}_{i}^{2}}\text{,}$ (36)
where $\hat{W}_{i}$ is the residual from regressing $Z_{i}\hat{B}$ on $[1,Z_{i}\hat{A}]\equiv\hat{M}_{i}$. That is,
$\hat{W}_{i}$ $=$ $Z_{i}\hat{B}-\hat{M}_{i}\underset{\equiv\hat{C}}{\underbrace{\left(\frac{1}{n}\sum\nolimits_{i}\hat{M}_{i}^{\prime}\hat{M}_{i}\right)^{-1}\left(\frac{1}{n}\sum\nolimits_{i}\hat{M}_{i}^{\prime}Z_{i}\right)}}\hat{B}$ (37)
$=$ $\left(Z_{i}-\hat{M}_{i}\hat{C}\right)\hat{B}\text{,}$
where
[…]
We now turn to the denominator of the 2SLS estimator of slope coefficient of $\Delta D_{2}$ in (36):
$\frac{1}{n}\sum\nolimits_{i}\hat{W}_{i}^{2}=$ $\hat{B}^{\prime}\left[\frac{1}{n}\sum\nolimits_{i}(Z_{i}-\hat{M}_{i}\hat{C})^{\prime}(Z_{i}-\hat{M}_{i}\hat{C})\right]\hat{B}$
[…]
$=$ $E\left[(Z_{i}B-M_{i}CB)^{\prime}(Z_{i}B-M_{i}CB)\right],$
[…]
Next, note the numerator is
$\frac{1}{n}\sum\nolimits_{i}\hat{W}_{i}\left(\Delta Y_{i2}\right)=$ $\hat{B}^{\prime}\left[\frac{1}{n}\sum\nolimits_{i}(Z_{i}-\hat{M}_{i}\hat{C})^{\prime}(\Delta Y_{i2})\right]$
[…]
$=$ $E\left[(Z_{i}B-M_{i}CB)^{\prime}(\Delta Y_{i2})\right]=E[w(Y_{i0},D_{i1})\left(\Delta Y_{i2}\right)]\text{,}$
[…]
Proof of Proposition 1. Throughout the following proof, we fix $s\geq 1$. For $t=s$, the equivalence is immediate:
$E[Y_{s}(d^{s})|Y^{s-1},D^{s-1}=d^{s-1},D_{s}=d_{s}^{\prime}]\quad\text{is invariant in $d_{s}^{\prime}$}$
$\iff$ $E[Y_{s}(d^{s})|Y^{s-1}(d^{s-1}),D^{s-1}=d^{s-1},D_{s}=d_{s}^{\prime}]\quad\text{is invariant in $d_{s}^{\prime}$}$
$\iff$ $E[Y_{s}(d^{s})|Y^{s-1}(d^{s-1}),D^{s-1}=d^{s-1},D_{s}=d_{s}^{\prime}]-Y_{s-1}(d^{s-1})\quad\text{is invariant in $d_{s}^{\prime}$}$
$\iff$ $E[Y_{s}(d^{s})-Y_{s-1}(d^{s-1})|Y^{s-1}(d^{s-1}),D^{s-1}=d^{s-1},D_{s}=d_{s}^{\prime}]\quad\text{is invariant in $d_{s}^{\prime}$.}$
For $t>s$, Assumption 2 implies:
$E[Y_{t}(d^{t})|Y^{s-1},D^{s-1}=d^{s-1},D_{s}=d_{s}^{\prime}]\quad\text{is invariant in $d_{s}^{\prime}$}$
$E[Y_{t-1}(d^{t-1})|Y^{s-1},D^{s-1}=d^{s-1},D_{s}=d_{s}^{\prime}]\quad\text{is invariant in $d_{s}^{\prime}$}$
and therefore:
$E[Y_{t}(d^{t})-Y_{t-1}(d^{t-1})|Y^{s-1},D^{s-1}=d^{s-1},D_{s}=d_{s}^{\prime}]\quad\text{is invariant in $d_{s}^{\prime}$}.$
We prove the reverse direction of implication for $t>s$ inductively. Suppose we have shown that, for some $t\geq s$, the conditions on the trends $Y_{t}(d^{t})-Y_{t-1}(d^{t-1})$ in Proposition 1 imply:
$E[Y_{t}(d^{t})|Y^{s-1},D^{s-1}=d^{s-1},D_{s}=d_{s}^{\prime}]\quad\text{is invariant in $d_{s}^{\prime}$}.$
By the conditions on the trends, we also have:
$E[Y_{t+1}(d^{t+1})-Y_{t}(d^{t})|Y^{s-1},D^{s-1}=d^{s-1},D_{s}=d_{s}^{\prime}]\quad\text{is invariant in $d_{s}^{\prime}$.}$
Combining these two equalities then implies:
$E[Y_{t+1}(d^{t+1})|Y^{s-1},D^{s-1}=d^{s-1},D_{s}=d_{s}^{\prime}]\quad\text{is invariant in $d_{s}^{\prime}$.}$
Finally, recall we have already shown the equivalence (between the two sets of conditions on the levels and the trends of potential outcomes) for $t=s$ at the beginning of this proof. Therefore, the claim of the reverse direction of implication for $t>s$ holds.
For proofs of Propositions 2 and 3, it is useful to expand the definition of (mean) sequential exchangeability for $T=2$ to full sequential exchangeability, which is:
$(Y_{1}(d_{1}),Y_{2}(d_{1},d_{2}))\perp D_{1}|Y_{0},$ (38)
$Y_{2}(d_{1},d_{2})\perp D_{2}|Y_{0},Y_{1},D_{1}=d_{1},$ (39)
whereas mean sequential exchangeability is:
$E[Y_{1}(d_{1})|Y_{0},D_{1}]$ $=E[Y_{1}(d_{1})|Y_{0}];$ (40)
$E[Y_{2}(d_{1},d_{2})|Y_{0},D_{1}]$ $=E[Y_{2}(d_{1},d_{2})|Y_{0}];$ (41)
$E[Y_{2}(d_{1},d_{2})|Y_{0},Y_{1},D_{1}=d_{1},D_{2}]$ $=E[Y_{2}(d_{1},d_{2})|Y_{0},Y_{1},D_{1}=d_{1}]$ (42)
for all $d_{1},d_{2}\in\{0,1\}$.
Proof of Proposition 2. By Lemma 1, the estimator for the coefficient $\beta$ converges in probability to (13), recalled here for convenience:
$E\left[\frac{w(Y_{0},D_{1})D_{2}}{E[w(Y_{0},D_{1})^{2}]}\times\tau_{2}(D_{1})\right]+E\left[\frac{w(Y_{0},D_{1})}{E[w(Y_{0},D_{1})^{2}]}\times\delta_{2}(D_{1})\right]$
where $w(Y_{0},D_{1})$ is the projection error in the linear projection of $E(\Delta D_{2}|Y_{0},D_{1})$ on $E(\Delta Y_{1}|Y_{0},D_{1})$.
The first step of the proof is to derive a closed-form expression for the projection errors $w(Y_{0},D_{1})$. We have:
$E[\Delta Y_{1}|Y_{0},D_{1}]$ $=E[\delta_{1}+D_{1}\tau_{1}|Y_{0},D_{1}]$
$=E[\delta_{1}|Y_{0},D_{1}]+D_{1}E[\tau_{1}|Y_{0},D_{1}]$
$=E[\delta_{1}|Y_{0}]+D_{1}E[\tau_{1}|Y_{0}]$
$=E[\delta_{1}]+D_{1}E[\tau_{1}]$
where the first equality is due to the decomposition in (11), the second equality rearranges, the third equality follows from mean sequential exchangeability (Assumption 2b), specifically (40), and the fourth equality follows from limited heterogeneity (Assumption 3), specifically (14) and (16). Furthermore, by (16), this expression is strictly monotone in the realization $D_{1}$. Therefore, the linear projection of $E(\Delta D_{2}|Y_{0},D_{1})$ on $E(\Delta Y_{1}|D_{1},Y_{0})$ is identical to its mean conditional on $D_{1}$. This implies:
$w(Y_{0},D_{1})$ $=E(\Delta D_{2}|Y_{0},D_{1})-E[E(\Delta D_{2}|Y_{0},D_{1})|D_{1}]$
$=E(D_{2}|Y_{0},D_{1})-E(D_{2}|D_{1}).$ (43)
The second step of the proof is to eliminate the second bias term in (13) by establishing that its numerator:
$E[w(Y_{0},D_{1})\delta_{2}(D_{1})]$
equals zero. Applying the law of iterated expectations (LIE) over partitions $(Y_{0},D_{1})$, this numerator equals:
$E[w(Y_{0},D_{1})E[\delta_{2}(D_{1})|Y_{0},D_{1}]]$ (44)
Focusing on the inner expectation for treatment realization $d_{1}$,
$E[\delta_{2}(D_{1})|Y_{0},D_{1}=d_{1}]=E[\delta_{2}(d_{1})|Y_{0},D_{1}=d_{1}]=E[\delta_{2}(d_{1})|Y_{0}]=E[\delta_{2}(d_{1})]$ (45)
where the first equation plugs in realized treatment, the second follows from mean sequential exchangeability (Assumption 2b), specifically (40) and (41), and the third follows from limited heterogeneity (Assumption 3), specifically (15). Furthermore, we have:
$E[\delta_{2}(D_{1})|D_{1}=d_{1}]$ $=E[E[\delta_{2}(D_{1})|Y_{0},D_{1}=d_{1}]|D_{1}=d_{1}]$
$=E[E[\delta_{2}(d_{1})]|D_{1}=d_{1}]$
$=E[\delta_{2}(d_{1})]$
where the first equality follows from the LIE, the second from the previous equality (45), and the third from taking an expectation over a constant. Combining with (45), it follows that:
$E[\delta_{2}(D_{1})|Y_{0},D_{1}]=E[\delta_{2}(D_{1})|D_{1}].$ (46)
Returning to expression (44) of the second numerator:
$E[w(Y_{0},D_{1})E[\delta_{2}(D_{1})|Y_{0},D_{1}]]$ $=E[w(Y_{0},D_{1})E[\delta_{2}(D_{1})|D_{1}]]$
$=E[E[w(Y_{0},D_{1})|D_{1}]E[\delta_{2}(D_{1})|D_{1}]]$
$=0,$
where the first equality follows from (46), the second by the LIE over partitions $D_{1}$, and the third because $E[w(Y_{0},D_{1})|D_{1}]=0$ by definition of $w(Y_{0},D_{1})$.
The third step of the proof is to establish the convex weights over heterogeneous treatment effects in the first term of (13), with numerator $E[w(Y_{0},D_{1})D_{2}\tau_{2}(D_{1})]$. Applying the LIE over partitions $(Y_{0},Y_{1},D_{1},D_{2})$ and invoking mean sequential exchangeability (Assumption 2b), specifically (42), yields:
$E[w(Y_{0},D_{1})D_{2}\tau_{2}(D_{1})]$ $=E[w(Y_{0},D_{1})D_{2}E[\tau_{2}(D_{1})|Y_{0},Y_{1},D_{1},D_{2}]]$
$=E[w(Y_{0},D_{1})D_{2}E[\tau_{2}(D_{1})|Y_{0},Y_{1},D_{1}]]$ (47)
Focusing on the inner expectation for a treatment realization $D_{1}=d_{1}$,
$E[\tau_{2}(D_{1})|Y_{0},Y_{1},D_{1}=d_{1}]=E[\tau_{2}(d_{1})|Y_{0},Y_{1}(d_{1}),D_{1}=d_{1}]=E[\tau_{2}(d_{1})|Y_{0},Y_{1}(d_{1})]=E[\tau_{2}(d_{1})]$ (48)
where the first equality follows by definition of the observed variables at realized treatment, the second by the full sequential exchangeability (Assumption 2a) for period 1 in (38), specifically its implication that $D_{1}\perp Y_{2}(\cdot)|Y_{0},Y_{1}(d_{1})$ for any $d_{1}$, and the third by limited heterogeneity (Assumption 3), specifically (17). Similarly to the previous step, we also have:
$E[\tau_{2}(D_{1})|D_{1}=d_{1}]$ $=E[E[\tau_{2}(D_{1})|Y_{0},Y_{1},D_{1}=d_{1}]|D_{1}=d_{1}]$
$=E[E[\tau_{2}(d_{1})]|D_{1}=d_{1}]$
$=E[\tau_{2}(d_{1})]$
where the first equality follows from the LIE, the second from the previous equality (48), and the third from taking an expectation over a constant. Combining with (48), it follows that:
$E[\tau_{2}(D_{1})|Y_{0},Y_{1},D_{1}]=E[\tau_{2}(D_{1})|D_{1}].$ (49)
Proceeding with expression (47) for the first numerator:
$E[w(Y_{0},D_{1})D_{2}E[\tau_{2}(D_{1})|Y_{0},Y_{1},D_{1}]]$ $=E[w(Y_{0},D_{1})D_{2}E[\tau_{2}(D_{1})|D_{1}]]$
$=E[w(Y_{0},D_{1})E[D_{2}|D_{1},Y_{0}]E[\tau_{2}(D_{1})|D_{1}]]$
$=E[w(Y_{0},D_{1})(E[D_{2}|D_{1},Y_{0}]-E[D_{2}|D_{1}])E[\tau_{2}(D_{1})|D_{1}]]$
$=E[w(Y_{0},D_{1})^{2}E[\tau_{2}(D_{1})|D_{1}]]$
$=E[w(Y_{0},D_{1})^{2}E[\tau_{2}(D_{1})|Y_{0},Y_{1},D_{1}]]$
$=E[E[w(Y_{0},D_{1})^{2}\tau_{2}(D_{1})|Y_{0},Y_{1},D_{1}]]$
$=E[w(Y_{0},D_{1})^{2}\tau_{2}(D_{1})]$
where the first equality follows from (49), the second by the LIE over partition $(D_{1},Y_{0})$, the third by the LIE and definition of $w(Y_{0},D_{1})$ (which imply $E[w(Y_{0},D_{1})g(D_{1})]=0$ for any function $g(\cdot)$ with a sole argument $D_{1}$), the fourth by the previously derived projection errors (43), the fifth again by (49), the sixth by rearrangement, and the seventh again by the LIE.
Proof of Proposition 3. Plugging in the decomposition of $\Delta Y_{2}$ in (12), we get:
$E[\widetilde{\Delta Y_{2}}]=E\left[\frac{D_{2}\tau_{2}(D_{1})+\delta_{2}(D_{1})-E[\Delta Y_{2}|Y^{1},D_{1},D_{2}=0]}{E[D_{2}|Y^{1},D_{1}]}\right]$ (50)
Next, we show the second and third terms in the numerator cancel. Note:
$E[\Delta Y_{2}|Y^{1},D_{1},D_{2}=0]$ $=E[Y_{2}-Y_{1}|Y_{0},Y_{1},D_{1},D_{2}=0]$
$=E[Y_{2}(D_{1},0)|Y_{0},Y_{1},D_{1},D_{2}=0]-Y_{1}$
$=E[Y_{2}(D_{1},0)|Y_{0},Y_{1},D_{1}]-Y_{1}$
$=E[\delta_{2}(D_{1})|Y^{1},D_{1}],$
where the first equality follows by definition of $\Delta Y_{2}$, the second by plugging in observed variables and rearrangement, the third by mean sequential exchangeability (Assumption 2), specifically (42), and the fourth by rearrangement and the definition (9). Iterating expectations over the previous equality,
$E[E[\Delta Y_{2}|Y^{1},D_{1},D_{2}=0]]=E[E[\delta_{2}(D_{1})|Y^{1},D_{1}]]=E[\delta_{2}(D_{1})].$
Plugging into (50) and simplifying, we have:
$E[\widetilde{\Delta Y_{2}}]=E\left[\frac{D_{2}\tau_{2}(D_{1})}{E[D_{2}|Y^{1},D_{1}]}\right].$ (51)
We finish the proof by showing how $E[\widetilde{\Delta Y_{2}}]$ thus identifies the ATE:
$E[\widetilde{\Delta Y_{2}}]$ $=E[E[\widetilde{\Delta Y_{2}}|Y^{1},D_{1}]]$
$=E\left[\frac{E[D_{2}\tau_{2}(D_{1})|Y^{1},D_{1}]}{E[D_{2}|Y^{1},D_{1}]}\right]$
$=E\left[\frac{E[D_{2}E[\tau_{2}(D_{1})|Y^{1},D_{1},D_{2}]|Y^{1},D_{1}]}{E[D_{2}|Y^{1},D_{1}]}\right]$
$=E\left[\frac{E[D_{2}E[\tau_{2}(D_{1})|Y^{1},D_{1}]|Y^{1},D_{1}]}{E[D_{2}|Y^{1},D_{1}]}\right]$
$=E[E[\tau_{2}(D_{1})|Y^{1},D_{1}]]$
$=E[\tau_{2}(D_{1})],$
where the first equality follows by the law of iterated expectations (LIE) over $(Y^{1},D_{1})$, the second by (51), the third again by the LIE over $(Y^{1},D_{1},D_{2})$, the fourth by mean sequential exchangeability (Assumption 2), specifically (42), the fifth by rearrangement and cancellation of terms $E[D_{2}|Y^{1},D_{1}]$, and the last again by the LIE. (Note mean sequential exchangeability is invoked only for $D_{2}$ in (42).)
Proof of Proposition 4. For all $s\geq 1$,
$E(\varepsilon_{is}\alpha_{i})$ $=$ $E[\alpha_{i}E\left(\varepsilon_{is}|Y_{i0},\alpha_{i}\right)]$
$=$ $E\left\{\alpha_{i}\left[\sum\nolimits_{d^{s}}E\left(\varepsilon_{is}|D_{i}^{s}=d^{s},Y_{i0},\alpha_{i}\right)\times\Pr\{D_{i}^{s}=d^{s}|Y_{i0},\alpha_{i}\}\right]\right\}$
$=$ $E\left\{\alpha_{i}\left[\sum\nolimits_{d^{s}}E\left(\varepsilon_{is}^{\ast}(d_{s})|D_{i}^{s}=d^{s},Y_{i0},\alpha_{i}\right)\times\Pr\{D_{i}^{s}=d^{s}|Y_{i0},\alpha_{i}\}\right]\right\}=0\text{,}$
where the first and second equalities are due to the law of iterated expectation, the third due to the definition $\varepsilon_{it}=\varepsilon_{it}^{\ast}(D_{it})$, and the last due to the sequential exogeneity of $\varepsilon_{it}^{\ast}(\cdot)$ in (25). Hence $E(\Delta\varepsilon_{it}\alpha_{i})=0$ for all $t\geq 2$. Likewise, swapping the roles of $\alpha_{i}$ and $Y_{i0}$ and applying the same argument above shows that $E(\Delta\varepsilon_{it}Y_{i0})=0$.
Furthermore, note that for $s\leq t$, the law of iterated expectation implies
$E(D_{is}\varepsilon_{it})=E[D_{is}E(\varepsilon_{it}|D_{i}^{t},Y_{i0},\alpha_{i})]\text{,}$
where
$E\left[\varepsilon_{it}|D_{i}^{t}=d^{t}\equiv(d^{t-1},d_{t}),Y_{i0},\alpha_{i}\right]$
$=$ $E\left[\varepsilon_{it}^{\ast}(d_{t})|D_{i}^{t}=(d^{t-1},d_{t}),Y_{i0},\alpha_{i}\right]=0$
for all $d^{t}$, where the second equality holds due to (25). Thus (25) implies (28).
Next, without loss of generality, consider $s<t$ so that
$E(\varepsilon_{is}\varepsilon_{it})$ $=$ $E[E(\varepsilon_{is}\varepsilon_{it}|\varepsilon_{is},D_{i}^{s},Y_{i0},\alpha_{i})]$
$=$ $E\left\{\varepsilon_{is}^{\ast}(D_{is})E[\varepsilon_{it}|\varepsilon_{is}^{\ast}(D_{is}),D_{i}^{s},Y_{i0},\alpha_{i}]\right\}\text{.}$
By the law of iterated expectations,
$E\left[\varepsilon_{it}|\varepsilon_{is}^{\ast}(D_{is}),D_{i}^{s}=d^{s},Y_{i0},\alpha_{i}\right]$
$=$ $\sum_{d^{t}\in\omega_{t}(d^{s})}E\left[\varepsilon_{it}^{\ast}(d_{t})|D_{i}^{t}=d^{t},\varepsilon_{is}^{\ast}(d_{s}),Y_{i0},\alpha_{i}\right]\Pr\{D_{i}^{t}=d^{t}|\varepsilon_{is}^{\ast}(d_{s}),D_{i}^{s}=d^{s},Y_{i0},\alpha_{i}\}$
$=$ $0\text{,}$
where $\omega_{t}(d^{s})$ denotes the support of $D_{i}^{t}$ conditional on $D_{i}^{s}=d^{s}$, and the last equality holds because $E\left[\varepsilon_{it}^{\ast}(d_{t})|D_{i}^{t}=d^{t},\varepsilon_{is}^{\ast}(d_{s}),Y_{i0},\alpha_{i}\right]$ is zero under (25).
In what follows, it will be useful to collect the following relationships between potential outcomes and composite errors $\nu_{it}^{*}(d_{t})\equiv\alpha_{i}^{*}+\varepsilon_{it}^{*}(d_{t})$ in the linear dynamic model of potential outcomes (21).
Lemma 2
For any $d^{t}\in\{0,1\}^{t}$ and $t\geq s\geq 1$,
$\nu_{it}^{*}(d_{t})$ $=Y_{it}(d^{t})-\gamma^{*}Y_{it-1}(d^{t-1})-\beta^{*}d_{t}-\theta_{t}^{*}$ (52)
$Y_{it}(d^{t})$ $=(\gamma^{*})^{t-s+1}Y_{is-1}(d^{s-1})+\sum_{r=s}^{t}(\gamma^{*})^{t-r}\times[\beta^{*}d_{r}+\theta_{r}^{*}+\nu_{ir}^{*}(d_{r})]$ (53)
Proof of Lemma 2. The first equality (52) is immediate from rearranging the linear potential outcome model (21). The second equality (53) follows by induction on $t$ given $s$. For any $s\geq 1$, the base case $t=s$ repeats (21) and is thus immediate. Now suppose (53) holds for some index $t-1\geq s$:
$Y_{it-1}(d^{t-1})=(\gamma^{*})^{t-s}Y_{is-1}(d^{s-1})+\sum_{r=s}^{t-1}(\gamma^{*})^{t-1-r}\times[\beta^{*}d_{r}+\theta_{r}^{*}+\nu_{ir}^{*}(d_{r})]$
It suffices to show the formula also holds for $t$. For $t$, the linear potential outcome model states:
$Y_{it}(d^{t})=\gamma^{*}Y_{it-1}(d^{t-1})+\beta^{*}d_{t}+\theta_{t}^{*}+\nu_{it}^{*}(d_{t})$
Plugging in from the preceding equation:
$Y_{it}(d^{t})=$ $\gamma^{*}\left[(\gamma^{*})^{t-s}Y_{is-1}(d^{s-1})+\sum_{r=s}^{t-1}(\gamma^{*})^{t-1-r}\times[\beta^{*}d_{r}+\theta_{r}^{*}+\nu_{ir}^{*}(d_{r})]\right]$
$+\beta^{*}d_{t}+\theta_{t}^{*}+\nu_{it}^{*}(d_{t})$
Collecting terms yields the desired result.
Proof of Proposition 5. For $t\geq s\geq 1$, we have:
$E[Y_{it}(d^{t})|\alpha_{i}^{*},Y_{i}^{s-1},D_{i}^{s-1}=d^{s-1},D_{is}=d_{s}^{\prime}]$
$=$ $E[Y_{it}(d^{t})|\alpha_{i}^{*},Y_{i}^{s-1}(d^{s-1}),D_{i}^{s-1}=d^{s-1},D_{is}=d_{s}^{\prime}]$
$=$ $(\gamma^{*})^{t-s+1}Y_{is-1}(d^{s-1})+$
$\sum_{r=s}^{t}(\gamma^{*})^{t-r}\times[\beta^{*}d_{r}+\theta_{r}^{*}+\alpha_{i}^{*}+E[\varepsilon_{ir}^{*}(d_{r})|\alpha_{i}^{*},Y_{i}^{s-1}(d^{s-1}),D_{i}^{s-1}=d^{s-1},D_{is}=d_{s}^{\prime}]]$
where the first equality follows by substituting realized treatment, and the second by (53) of Lemma 2. It therefore suffices to show that:
$E[\varepsilon_{it}^{*}(d_{t})|\alpha_{i}^{*},Y_{i}^{s-1}(d^{s-1}),D_{i}^{s-1}=d^{s-1},D_{is}=d_{s}^{\prime}]$
is invariant in $d_{s}^{\prime}$, or a fortiori that the expectation equals zero, for $t\geq s$. Lemma 2 implies that the preceding expression is equal to:
$E[\varepsilon_{it}^{*}(d_{t})|\alpha_{i}^{*},Y_{i0},D_{i}^{s-1}=d^{s-1},D_{is}=d_{s}^{\prime},\nu_{i}^{*s-1}(d^{s-1})]$
because $\nu_{i}^{*s-1}(d^{s-1})$ can be recovered from $Y_{i}^{s-1}(d^{s-1})$ by (52), and conversely $Y_{i}^{s-1}(d^{s-1})$ can be recovered from $(Y_{i0},\nu_{i}^{*s-1}(d^{s-1}))$ by (53). Since $\varepsilon_{i}^{*s-1}(d^{s-1})=\alpha_{i}^{*}-\nu_{i}^{*s-1}(d^{s-1})$, this is equivalent to:
$E[\nu_{it}^{*}(d_{t})|\alpha_{i}^{*},Y_{i0},D_{i}^{s-1}=d^{s-1},D_{is}=d_{s}^{\prime},\varepsilon_{i}^{*s-1}(d^{s-1})]$
Finally, iterating expectations in the preceding expression and invoking (25) yields the desired result:
$E[E[\varepsilon_{it}^{*}(d_{t})|\alpha_{i}^{*},Y_{i0},D_{i}^{t},\varepsilon^{*t-1}(\cdot)]|\alpha_{i}^{*},Y_{i0},D_{i}^{s-1}=d^{s-1},D_{is}=d_{s}^{\prime},\varepsilon^{*s-1}(d^{s-1})]=0.$
Proof of Proposition 6.
Fix a sequence of treatment $d^{T}$. For simplicity of notation, expand a vector-valued representation of choice rule (28) as:
$D^{t}=\phi^{t}(D^{t-1},Y^{t-1}(D^{t-1}),\zeta^{t}(D^{t-1}))$
so that we can express the event:
$\{D^{t}=d^{t}\}=\{\phi^{t}(d^{t-1},Y^{t-1}(d^{t-1}),\zeta^{t}(d^{t-1}))=d^{t}\}$ (54)
Condition (29) implies:
$Y^{t}(d^{t})\perp\zeta^{t}(d^{t-1})\mid Y_{0}$
because $Y^{t}(d^{t})$ is a function of $(Y_{0},\alpha(\cdot),\varepsilon^{t}(d^{t}))$ by definition (27). By rules of conditional independence,
$(Y_{t}(d^{t}),\dots,Y_{T}(d^{T}))\perp\zeta^{t}(d^{t-1})\mid Y^{t-1}(d^{t-1})$
and furthermore, for $t\geq 2$,
$(Y_{t}(d^{t}),\dots,Y_{T}(d^{T}))\perp\zeta^{t}(d^{t-1})\mid Y^{t-1}(d^{t-1}),\,\phi^{t-1}(d^{t-2},Y^{t-2}(d^{t-2}),\zeta^{t-1}(d^{t-2}))=d^{t-1}$
because, having conditioned on $Y^{t-1}(d^{t-1})$, the second conditioning statement is a function of $\zeta^{t-1}(d^{t-2})$. By (54) and because $\zeta^{t}(D^{t-1})=\zeta^{t}$ and $Y^{t-1}(D^{t-1})=Y^{t-1}$, this reduces to:
$(Y_{t}(d^{t}),\dots,Y_{T}(d^{T}))\perp\zeta^{t}\mid Y^{t-1},\,D^{t-1}=d^{t-1}$
Conditioning on history $(D^{t-1},Y^{t-1})$, this implies that, for a deterministic function $\phi_{t}(\cdot)$:
$(Y_{t}(d^{t}),\dots,Y_{T}(d^{T}))\perp\phi_{t}(D^{t-1},Y^{t-1},\zeta^{t})\mid Y^{t-1},\,D^{t-1}=d^{t-1}$
which by definition (28) yields the desired result:
$(Y_{t}(d^{t}),\dots,Y_{T}(d^{T}))\perp D_{t}|Y^{t-1},D^{t-1}=d^{t-1}$
In the case where $t=1$, the conditioning event collapses to $Y_{0}$ and the argument proceeds analogously.
Proof of Proposition 7. Assumption 2b holds under (33) by the same argument as in the proof of Proposition 6. We adapt it here for the weaker case of mean sequential exchangeability in simpler notation. With potential outcomes and treatment rules in (30) and (31), we have:
$E[Y_{2}(d_{1},d_{2})|D_{2}=d_{2},Y_{1}=y_{1},D_{1}=d_{1},Y_{0}=y_{0}]$
$=$ $E\left[\lambda_{2}(d_{1},d_{2},y_{0},y_{1},\alpha(d_{2}),\varepsilon_{2}(d_{2}))\left|\begin{array}[]{c}\phi_{2}(y_{0},y_{1},d_{1},\xi^{1}(d_{1}),\eta_{2}(\cdot))=d_{2},\\ \lambda_{1}(d_{1},y_{0},\alpha(d_{1}),\varepsilon_{1}(d_{1}))=y_{1},\\ \phi_{1}(y_{0},\xi_{0},\eta_{1}(\cdot))=d_{1},\text{ \ }Y_{0}=y_{0}\end{array}\right.\right]$
$=$ $E[\lambda_{2}(d_{1},d_{2},y_{0},y_{1},\alpha(d_{2}),\varepsilon_{2}(d_{2}))|\lambda_{1}(d_{1},y_{0},\alpha(d_{1}),\varepsilon_{1}(d_{1}))=y_{1},Y_{0}=y_{0}]\text{,}$
where the last equality is due to (33) and the right-hand side does not vary with $d_{2}$.
Also note:
$E[Y_{2}(d_{1},d_{2})|D_{1}=d_{1},Y_{0}]$
$=$ $E[\lambda_{2}(0,0,y_{0},Y_{1}(d_{1}),\alpha(d_{2}),\varepsilon_{2}(d_{2}))|\phi_{1}(y_{0},\xi_{0},\eta_{1}(\cdot))=d_{1},Y_{0}=y_{0}]$
$=$ $E[\lambda_{2}(0,0,y_{0},Y_{1}(d_{1}),\alpha(d_{2}),\varepsilon_{2}(d_{2}))|Y_{0}=y_{0}]\text{,}$
where the last equality holds because of (33) and the fact that $Y_{1}(d_{1})$ is a function of $(Y_{0},\alpha(d_{1}),\varepsilon_{1}(d_{1}))$. Again, the right-hand side does not vary with $d_{1}$.
Furthermore,
$E[Y_{1}(d_{1})|D_{1}=d_{1},Y_{0}]$
$=$ $E[\lambda_{1}(d_{1},y_{0},\alpha(d_{1}),\varepsilon_{1}(d_{1}))|\phi_{1}(y_{0},\xi_{0},\eta_{1}(\cdot))=d_{1},Y_{0}=y_{0}]$
$=$ $E[\lambda_{1}(d_{1},y_{0},\alpha(d_{1}),\varepsilon_{1}(d_{1}))|Y_{0}=y_{0}]\text{,}$
where the last equality is due to (33) and the right-hand side does not vary with $d_{1}$. Thus, mean sequential exogeneity in Assumption 2b holds.
On the other hand, under (33),
$E[Y_{2}(0,0)-Y_{1}(0)|(D_{1},D_{2})=(d_{1},d_{2}),Y_{0}=y_{0}]$
$=$ $E\left[\lambda_{2}(0,0,y_{0},Y_{1}(0),\alpha(0),\varepsilon_{2}(0))-Y_{1}(0)\left|\begin{array}[]{c}\phi_{2}(y_{0},Y_{1}(d_{1}),d_{1},\xi^{1}(d_{1}),\eta_{2}(\cdot))=d_{2},\\ \phi_{1}(Y_{0},\xi_{0},\eta_{1}(\cdot))=d_{1},Y_{0}=y_{0}\end{array}\right.\right]$
where $Y_{2}(0,0)-Y_{1}(0)$ is only a deterministic function of $(Y_{0},\alpha(0),\varepsilon_{1}(0),\varepsilon_{2}(0))$. Under (33), the right-hand side does generally vary with $(d_{1},d_{2})$ even after controlling for $Y_{0}$. This is because the event “$(D_{1},D_{2})=(d_{1},d_{2})$” does restrict the joint support of $(Y_{0},\alpha(d_{1}),\varepsilon_{1}(d_{1}),\xi^{1}(d_{1}),\eta^{2}(\cdot))$. [^8] Note $Y_{1}(d_{1})=\lambda_{1}(d_{1},Y_{0},\alpha(d_{1}),\varepsilon_{1}(d_{1}))$ affects the definition of conditioning events due to its entry through $\phi_{2}(\cdot)$. Hence, even after conditioning on $Y_{0}=y_{0}$, the right-hand side is in general non-degenerate in $(d_{1},d_{2})$.
Appendix B: Consistency of Adjusted IPW Estimator
In this appendix, we establish the consistency of the adjusted IPW estimator $\widehat{\mu}_{\tau_{2}}$, which is defined in (20). Maintain the following conditions:
[…]
and
[…]
Condition (57) requires uniform convergence of preliminary estimators $\widehat{M}_{1i},\widehat{M}_{2i}$; it holds for many nonparametric (e.g. Nadaraya-Watson) estimators of conditional means under standard conditions. Condition (58) include standard, mild requirement that the sample moments of $M_{1i},M_{2i}$ are well-behaved.
Corollary B1
[…]
Proof of Corollary B1. By the Mean Value Theorem,
$\hat{\mu}_{\tau_{2}}=\frac{1}{n}\sum\nolimits_{i}\left[\frac{\Delta Y_{i2}-M_{2i}}{M_{1i}}+\left(\frac{-1}{\tilde{M}_{1i}}\right)\left(\hat{M}_{2i}-M_{2i}\right)+\left(\frac{\tilde{M}_{2i}}{\tilde{M}_{1i}^{2}}\right)(\hat{M}_{1i}-M_{1i})\right]\text{,}$
where $(\Delta Y_{i2}-M_{2i})/M_{1i}$ is an “infeasible” summand that involves the actual population moments $M_{1i},M_{2i}$, and $\tilde{M}_{ki}$ are intermediate values bounded between $\hat{M}_{ki}$ and $M_{ki}$ for $k=1,2$. By the triangular inequality,
$\left|\hat{\mu}_{\tau_{2}}-\frac{1}{n}\left(\sum\nolimits_{i}\frac{\Delta Y_{i2}-M_{2i}}{M_{1i}}\right)\right|$
$\leq$ $\frac{1}{n}\sum\nolimits_{i}\left|\tilde{M}_{1i}\right|^{-1}\left|\hat{M}_{2i}-M_{2i}\right|+\frac{1}{n}\sum\nolimits_{i}\left|\tilde{M}_{2i}\right|\left|\tilde{M}_{1i}\right|^{-2}\left|\hat{M}_{2i}-M_{2i}\right|$
$\leq$ $\left(\frac{1}{n}\sum\nolimits_{i}\left|\tilde{M}_{1i}\right|^{-1}\right)\left(\sup_{1\leq i^{\prime}\leq n}\left|\hat{M}_{2i^{\prime}}-M_{2i^{\prime}}\right|\right)+\left(\frac{1}{n}\sum\nolimits_{i}\left|\tilde{M}_{2i}\right|\left|\tilde{M}_{1i}\right|^{-2}\right)\left(\sup_{1\leq i^{\prime}\leq n}\left|\hat{M}_{1i}-M_{1i}\right|\right)\text{.}$
By (57), the two suprema of differences are $o_{p}(1)$, i.e., converge in probability to 00.
Next, we establish that $\frac{1}{n}\sum\nolimits_{i}\left|\tilde{M}_{1i}\right|^{-1}=O_{p}(1)$, i.e. is bounded in probability. For this, we want to show “$\forall\varepsilon>0$, $\exists\mathcal{C}_{\varepsilon}>0$ such that $\Pr\{\frac{1}{n}\sum\nolimits_{i}\left|\tilde{M}_{1i}\right|^{-1}>\mathcal{C}_{\varepsilon}\}<\varepsilon$ for $n$ large enough.” By construction,
$\Pr\{\frac{1}{n}\sum\nolimits_{i}\left|\tilde{M}_{1i}\right|^{-1}>\mathcal{C}_{\varepsilon}\}\leq\Pr\left\{\max\nolimits_{i\leq n}\left|\tilde{M}_{1i}\right|^{-1}>\mathcal{C}_{\varepsilon}\right\}$ (59)
$=$ $\Pr\left\{\min\nolimits_{i\leq n}\left|\tilde{M}_{1i}\right|<1/\mathcal{C}_{\varepsilon}\right\}\leq\Pr\left\{\min\nolimits_{i\leq n}M_{1i}-\tilde{e}_{1n}<1/\mathcal{C}_{\varepsilon}\right\}$
$=$ $\Pr\left\{\tilde{e}_{1n}>\min\nolimits_{i\leq n}M_{1i}-1/\mathcal{C}_{\varepsilon}\right\}\text{,}$
where $\tilde{e}_{kn}\equiv\sup_{i\leq n}\left|\tilde{M}_{ki}-M_{ki}\right|$ for $k=1,2$, and the second inequality is due to triangular inequality and positivity of $M_{i}$. By condition (i) in (58), $\min_{i}M_{i}-1/\mathcal{C}_{\varepsilon}$ is strictly positive when $\mathcal{C}_{\varepsilon}$ is large enough. The uniform convergence of $\hat{M}_{1i}$ in (57) then implies the right-hand side of (59) converges to zero as $n\rightarrow\infty$. This proves $\frac{1}{n}\sum\nolimits_{i}\left|\tilde{M}_{1i}\right|^{-1}=O_{p}(1)$.
Finally, we establish $\frac{1}{n}\sum\nolimits_{i}\left|\tilde{M}_{2i}\right|\left|\tilde{M}_{1i}\right|^{-2}$ is bounded in probability. Note:
$\frac{1}{n}\sum\nolimits_{i}\left|\tilde{M}_{2i}\right|\left|\tilde{M}_{1i}\right|^{-2}\leq\left(\frac{1}{n}\sum\nolimits_{i}\left|M_{2i}\right|+\tilde{e}_{2n}\right)\left(\max_{i^{\prime}}\left|\tilde{M}_{1i^{\prime}}\right|^{-2}\right)\text{,}$
due the triangular inequality and definition of $\tilde{e}_{2n}$. The second term on the right-hand side is $O_{p}(1)$ by the same argument above (with $\mathcal{C}_{\varepsilon}$ replaced by $\mathcal{C}_{\varepsilon}^{2}$). By condition (ii) in (58), the stochastic order of the first term on the right-hand side is $O_{p}(1)+o_{p}(1)=O_{p}(1)$. Thus the product of the two terms, and consequently $\frac{1}{n}\sum\nolimits_{i}\left|\tilde{M}_{2i}\right|\left|\tilde{M}_{1i}\right|^{-2}$, is bounded in probability.

### v2

@@
-Heterogeneous Intertemporal Treatment Effects via Dynamic Panel Data Models
+Heterogeneous Treatment Effects via Linear Dynamic
 Abstract
-We study the identification and estimation of heterogeneous, intertemporal treatment effects (TE) when potential outcomes depend on past treatments.
-First, applying a dynamic panel data model to observed outcomes, we show that instrument-based GMM estimators, such as Arellano and Bond (1991), converge to a non-convex (negatively weighted) aggregate of TE plus non-vanishing trends.
-We then provide restrictions on sequential exchangeability (SE) of treatment and TE heterogeneity that reduce the GMM estimand to a convex (positively weighted) aggregate of TE.
-Second, we introduce an adjusted inverse-propensity-weighted (IPW) estimator for a new notion of average treatment effect (ATE) over past observed treatments.
-Third, we show that when potential outcomes are generated by dynamic panel data models with homogeneous TE, such GMM estimators converge to causal parameters (even when SE is generically violated without conditioning on individual fixed effects).
-Finally, we motivate SE and compare it with parallel trends (PT) in various settings with observational data (when treatments are dynamic, rational choices under learning) or experimental data (when treatments are sequentially randomized).
+We study the identification of heterogeneous, intertemporal treatment effects (TE) when potential outcomes depend on past treatments.
+First, applying a dynamic panel data model to observed outcomes, we show that an instrumental variable (IV) version of the estimand in Arellano and Bond (1991) recovers a non-convex (negatively weighted) aggregate of TE plus non-vanishing trends.
+We then provide conditions on sequential exchangeability (SE) of treatment and on TE heterogeneity that reduce such an IV estimand to a convex (positively weighted) aggregate of TE.
+Second, even when SE is generically violated, such estimands identify causal parameters when potential outcomes are generated by dynamic panel data models with some homogeneity or mild selection assumptions.
+Finally, we motivate SE and compare it with parallel trends (PT) in various settings with experimental data (when treatments are sequentially randomized) and observational data (when treatments are dynamic, rational choices under learning).
 1 Introduction
-Consider a panel data setting in which a researcher observes individual outcomes $Y_{it}$ and treatments $D_{it}$ over multiple periods $t=1,...,T$.
-At any given time $t$, potential outcomes depend on a full sequence of potential treatments in the past, i.e., $Y_{it}(d^{t})$ where $d^{t}\equiv(d_{s})_{s=0}^{t}$ with $d_{s}\in\{0,1\}$.
+Consider a panel data setting in which a researcher observes individual outcomes $Y_{it}$ and treatments $D_{it}$ over multiple periods t=1,,Tt=1,...,T.
+At any given time $t$, potential outcomes depend on a full sequence of potential treatments in the past, i.e., $Y_{it}(d^{t})$ where $d^{t}\equiv(d_{s})_{s=1}^{t}$ with $d_{s}\in\{0,1\}$.
 The objective of causal inference in this context is to learn about certain moments of $Y_{it}(d^{t})$ using the joint distribution of $\{D_{it},Y_{it}\}_{t\leq T}$, where the observed outcomes are determined as $Y_{it}=Y_{it}(D_{i}^{t})$.
-This paper considers the question of causal inference through the lens of (linear) panel data models that also include a lagged outcome as a regressor, e.g.,
+We consider the question of causal inference through the lens of linear dynamic panel data models that also include a lagged outcome as a regressor, e.g.,
 $Y_{it}=\beta D_{it}+\gamma Y_{it-1}+\theta_{t}+\alpha_{i}+\varepsilon_{it}.$ (1)
 There are at least two reasons one might argue for the necessity of such a dynamic panel data model (DPDM).
-First, treatment may only be randomly assigned conditional on the past history.
-For example, when estimating the effect of a training program on earnings or employment outcomes, enrollment into the training program may depend on previous earnings and employment (Ashenfelter, 1978).
-In such cases, the lagged outcome is a necessary control for the exogeneity of treatment.
-Even though (1) appears to deviate from a static panel data model only by an additional control, the fact that this additional control is a lagged dependent variable raises concerns about the strict exogeneity assumption that is necessary for two-way fixed-effect (TWFE) regressions.
+First, treatments may only be randomly assigned conditional on the past observed history.
+For example, when estimating the effect of a training program on earnings or employment outcomes, enrollment in the program may depend on previous earnings or employment (Ashenfelter, 1978).
+In such cases, the lagged outcome is a necessary control for the exogeneity of the treatment $D_{it}$.
+Even though (1) appears to deviate from a static panel data model only by an additional control, the fact that this additional control is a lagged dependent variable raises concerns about the strict exogeneity condition on $(D_{it},Y_{it-1})$ that is necessary for two-way fixed-effect (TWFE) regressions.
 Second, outcomes may exhibit state-dependence.
@@
 A DPDM such as (1) allows researchers to disentangle the partial effects of $D_{it}$ from the state dependence in $Y_{it-1}$.
-One of our main questions is: when and how can we use existing dynamic panel data estimators (DPDE) of observed outcomes, particularly the GMM/2SLS estimators of Anderson and Hsiao (1982) and Arellano and Bond (1991), to shed light on causal effects?
+Section 2 introduces the potential outcomes framework underlying our analysis.
+We then formally state the sequential exchangeability assumption, an identifying condition extensively developed in the biostatistics literature (Robins, 1986), and we relate it to the parallel trends assumption that has been used in the econometrics literature.
+The sequential exchageability assumption requires that present potential outcomes are (mean) independent from contemporary treatment, once conditioned on a full history of observed outcomes and treatments.
+Section 3 investigates whether commonly used dynamic panel data estimators (DPDEs) for observed outcomes — most notably the first-differenced IV (and more generally, GMM) estimators of Anderson and Hsiao (1982) and Arellano and Bond (1991) — can be mapped to causal estimands of interest.
 These estimators use lagged regressors as instruments to consistently estimate $\beta,\gamma$ in the first-differenced transformation of (1):
@@
 where $\Delta$ denotes the first-difference operator.
-To answer this question, we begin by deriving a general causal decomposition of the aforementioned DPDEs (Lemma 1).
-Implicit in this decomposition, and more generally in our goal of causal inference, is the possibility that (1) and (2) are not structural models of the outcome-generating process, for example because of the possibility of treatment effect heterogeneity.
-This attempt to link such instrumented first-difference estimators of a reduced-form model to interpretations in terms of aggregated heterogeneous treatment effects (TE) is one contribution of this paper.
-The preceding exercise bears resemblance to, but differs qualitatively from, the work of De Chaisemartin and d’Haultfoeuille (2020) and Goodman-Bacon (2021) in decomposing and interpreting TWFE regressions.
-In a non-dynamic setting where potential outcomes are indexed only by contemporary treatment $Y_{it}(d_{t})$, they showed the TWFE estimand in static panel data models (a special case of (1) with $\gamma=0$) is a weighted average of heterogeneous TE $Y_{it}(1)-Y_{it}(0)$.
-[^1] The use of TWFE in static panel data models requires that $D_{it}$ be strict exogenous in order to identify $\beta$ and $\gamma$.
+To answer this question, we begin by deriving a general causal decomposition of the estimands of these DPDEs (Lemma 1).
+Implicit in this decomposition, and more generally in our goal of causal inference, is the possibility that (1) and (2) are not structural models of the outcome-generating process, for example, because of possible heterogeneity in treatment effects (TEs).
+Our analysis reveals a link between the instrumented first-difference estimators in a reduced-form model of observed outcomes (1) and the heterogeneous TEs on potential outcomes.
+This exercise bears resemblance to, but differs qualitatively from, the work of De Chaisemartin and d’Haultfoeuille (2020) and Goodman-Bacon (2021) in decomposing and interpreting TWFE regressions.
+In a non-dynamic setting where potential outcomes are indexed only by contemporary treatment $Y_{it}(d_{t})$, they showed the TWFE estimand in static panel data models, which is a special case of (1) with $\gamma=0$, is a (possibly negatively) weighted average of heterogeneous TE, $Y_{it}(1)-Y_{it}(0)$.222 The use of TWFE in static panel data models requires that $D_{it}$ be strict exogenous in order to identify $\beta$ and $\gamma$.
 However, the difference-in-differences approach in De Chaisemartin and d’Haultfoeuille (2020) does not impose that the true data-generating process satisfies the parametric, reduced-form $Y_{it}=\beta D_{it}+\theta_{t}+\alpha_{i}+\varepsilon_{it}$.
@@
 This literature has also considered the role and the complications of covariates (Heckman, Ichimura, and Todd, 1997; Abadie, 2005; Caetano and Callaway, 2023), as well as proposing alternative estimators under parallel trends (PT) and other assumptions on the outcome or treatment processes, like staggered designs (e.g.,De Chaisemartin and d’Haultfoeuille (2020); Callaway and Sant’Anna (2021); Sun and Abraham (2021)).
-In comparison, we face different challenges, because the treatment effects in our setting vary with the full sequence of past potential treatments, and the estimator in question uses instruments designed to address the failure of strict exogeneity of $D_{it}$ and $Y_{it-1}$ in the DPDM (1).
-In the context of papers imposing variants of parallel trends, the setting of de Chaisemartin and D’Haultfoeuille (2024) is most similar in that it also allows for dynamic potential outcomes and non-staggered designs; in contrast, however, our approach is based on a set of assumptions different than parallel trends.
-Building on our decomposition in Lemma 1, we derive sufficient conditions under which a 2SLS first-differenced estimator as in Anderson and Hsiao (1982) is guaranteed to be a positively weighted average of heterogeneous TE (Proposition 2).
-These conditions include the sequential exchangeability assumption of Robins (1986) — which posits that the present potential outcomes are (mean) independent from contemporary treatment, once conditioned on a full history of observed outcomes and treatments — as well as restrictions on the heterogeneity of TE.
-We extend and strengthen the result by introducing an adjusted inverse-(treament) propensity-weighted (IPW) estimator and show that a sample analogue is consistent for a particular average treatment effect.
-This new estimator dispenses with the restrictions on heterogeneous TE in Proposition 2; it only requires sequential exchangeability and some mild conditions on treatment assignment.
-(See Proposition 3.)
-By relying on an assumption of sequential exchangeability, our results relate closely to a broad suite of treatment effect estimators collectively referred to as $g$-methods that have been developed for causal inference in this setting [e.g., see Robins (1986, 1997) and Robins and Hernán (2008) for a survey].
+In comparison with these earlier works, we face different challenges, because the treatment effects in our setting vary with the full sequence of past potential treatments, and the estimator in question uses instruments designed to address the failure of strict exogeneity of observed $D_{it}$ and $Y_{it-1}$ in the DPDM (1).
+Among existing works that build on variants of parallel trends, the setting of de Chaisemartin and D’Haultfoeuille (2024) is most similar in that it also allows for intertemporal effects on potential outcomes in non-staggered designs; in contrast, however, our approach is based on sets of new assumptions different than parallel trends.
+Building on our decomposition in Lemma 1, we derive sufficient conditions under which an estimand using lagged variables as instruments in the first-differenced equation, as in Anderson and Hsiao (1982) and Arellano and Bond (1991), is guaranteed to be a positively weighted average of heterogeneous TE (Proposition 2).
+These conditions include the sequential exchangeability assumption of Robins (1986) as well as restrictions on the heterogeneity of TE.
+By invoking a condition of sequential exchangeability, our approach is related to, and complements, a broad suite of treatment effect estimators collectively referred to as $g$-methods that have been developed for causal inference in this setting [e.g., see Robins (1986, 1997) and Robins and Hernán (2008) for a survey].
 Most closely related, Robins (1997) proposes an IPW estimator that identifies $E[Y_{it}(d^{t})]$ using the observed outcomes $Y_{it}$ in stratum $\{D_{i}^{t}=d^{t}\}$; these potential outcomes can in turn be used to identify history-dependent treatment effects such as $E[Y_{it}(d^{t-1},1)-Y_{it}(d^{t-1},0)]$.
-In contrast, our adjusted IPW estimator identifies an aggregate treatment effect $E[Y_{it}(D_{i}^{t-1},1)-Y_{it}(D_{i}^{t-1},0)]$ using the first-differenced outcomes $\Delta Y_{it}$ among the entire sample.
-We conjecture that such an aggregation of treatment effects may be useful both from the standpoint of estimation and interpretation.
-In particular, our target parameter provides a potentially useful summary causal effect in multi-period settings where the number of potential contrasts grows large.
-The sequential exchangeability condition has played a central role in our results above (i.e., causal interpretation of dynamic panel data regression, and the adjusted IPW).
-This naturally raise two additional questions: (a) Can causal inference be conducted using DPDEs without relying on sequential exchangeability?
-(b) Is sequential exchangeability feasible in observational data where treatments are endogenously determined by forward-looking individuals?
+We connect our approach back to this existing estimator through an outcome-modifying first-differenced IV estimator (Proposition 3), which recovers the average treatment effect $E[Y_{it}(D^{t-1},1)-Y_{it}(D^{t-1},0)]$ while dispensing with substantive restrictions beyond sequential exchangeability in g-methods.
+The sequential exchangeability condition has played a central role in our results above (i.e., causal interpretation using DPDEs, and the outcome-modifying IV estimator).
+This naturally raises two follow-up questions: (a) Can causal inference be conducted using DPDEs without sequential exchangeability?
+(b) Can sequential exchangeability be justified in observational data where treatments are endogenously determined by forward-looking individuals?
 We provide conditions under which the answers to both questions are indeed positive.
-To answer (a), Propositions 4 and 5 delineate a case where our previous DPDEs identify well-defined causal estimands, and yet sequential exchangeability is generically violated without conditioning on further unobservables.
-Specifically, we focus on a parametric case where the data-generating process of the latent potential outcomes $Y_{it}(d^{t})$ follows a linear DPDM:
-$Y_{it}(d^{t})=\beta^{\ast}d_{t}+\gamma^{\ast}Y_{it-1}(d^{t-1})+\theta_{t}^{\ast}+\alpha_{i}^{\ast}+\varepsilon_{it}^{\ast}(d_{t})\text{.}$ (3)
-In this case, the observed outcomes $Y_{it}=Y_{it}(D_{i}^{t})$ are consistent with the reduced form DPDM in (1), where the coefficients coincide with constant causal effects (i.e., $\beta=\beta^{\ast},\gamma=\gamma^{\ast}$ as noted in Observation 2).
-A further assumption of conditional sequential exogeneity condition on the structural errors $\varepsilon_{it}^{\ast}(\cdot)$ guarantee consistency of our standard DPDEs but not sequential exchangeability, because the exogeneity between contemporaneous outcomes and treatments now only holds after conditioning on unobserved fixed effects.
-Hence, this allows for endogenous selection on such unobservables.
-To investigate (b), we provide, in the context of observational data, conditions under which sequential exchangeability holds.
-This provides critical guidance for empiricists applying g-methods to observational data, especially when treatment choices are most likely dependent upon past outcomes.
-For example, when estimating the effect of a training program on earnings or employment outcomes, enrollment into the training program may depend on previous earnings and employment Ashenfelter (1978).
-To this end, we consider sequential exchangeability in a model of dynamic choices that serves as an anchor for structurally motivating the conditional and unconditional versions of sequential exchangeability.
+To answer (a), Propositions 4, 5, and 6 delineate cases where sequential exchangeability is generically violated, and yet existing or modified DPDEs identify well-defined causal estimands.
+Specifically, we first focus on a parametric case where the data-generating process of the latent potential outcomes $Y_{it}(d^{t})$ follows a linear DPDM:
+$Y_{it}(d^{t})=\beta^{\ast}d_{t}+\gamma^{\ast}Y_{it-1}(d^{t-1})+\theta_{t}^{\ast}+\alpha_{i}^{\ast}+\varepsilon_{it}^{\ast}(d_{t})\text{.}$
+We show that in this case, the observed outcomes $Y_{it}=Y_{it}(D_{i}^{t})$ conform with a reduced-form DPDM in (1) in the following sense: (i) the coefficients in the DPDM coincide with constant causal effects, i.e., $\beta=\beta^{\ast},\gamma=\gamma^{\ast}$, and (ii) an assumption of conditional sequential exogeneity with respect to the structural errors $\varepsilon_{it}^{\ast}(\cdot)$ ensures the treatments $D_{it}$ are sequentially orthogonal to the implied reduced-form errors $\varepsilon_{it}\equiv\varepsilon^{*}_{it}(D_{i}^{t})$.
+Together, (i) and (ii) guarantee the existing DPDEs consistently estimate the constant causal effects, even when sequential exchangeability in Robins (1986) does not hold.333 The conditional sequential exogeneity of contemporaneous treatments and past outcomes in (ii) only holds after controlling for unobserved fixed effects.
+Hence, it allows for endogenous selection on such unobservables and violates the sequential exchangeability in Robins (1986).
+We extend the results for a more general model where only the untreated potential outcomes $Y_{it}(d^{t-1},0)$ take an auto-regressive form above while the intertemporal TEs $\tau_{it}(d^{t-1})\equiv Y_{it}(d^{t-1},1)-Y_{it}(d^{t-1},0)$ are time-varying and heterogeneous.
+We show that in this case, an IV regression conditional on observed history identifies a conditional average treatment effect on the treated, as well as time trends and the state dependence parameter $\gamma^{*}$.
+This result requires two mild assumptions: conditional serial uncorrelation of the structural errors of potential outcomes $\varepsilon^{*}_{t}(d^{t-1})$ (Assumption 6), and conditional mean independence of heterogeneous TE from past errors $\varepsilon_{is}^{*}(\cdot)\text{ for }s\leq t-1$ (Assumption 7).
+The latter is a weak restriction on the heterogeneous TEs in that it does not impose any “conditional unconfoundedness” condition.
+That is, it allows for conditional correlation between TEs and contemporary treatment choices.
+To investigate question (b), Section 4 provides, in the context of both experimental and observational data, conditions under which sequential exchangeability holds whereas a common alternative of parallel trends does not.
+For an experimental setting, we show sequentially randomized treatments are compatible with sequential exchangeability, but not compatible with parallel trends in general.
+We also provide critical guidance for empiricists to assess sequential exchangeability in observational data, especially when treatment choices are most likely dependent upon past outcomes.
+For example, when estimating the effect of a training program on earnings or employment outcomes, enrollment into the training program may depend on previous earnings and employment, as per Ashenfelter (1978).
+To this end, we consider sequential exchangeability in a model of dynamic choices that serves as a structural motivation.
 A particular focus of the model is on the possibility of learning, which is compatible with the nature of sequential exchangeability assumptions that condition (like a learning decision-maker) on the history of observed outcomes.
-The structural model provides a coherent dynamic decision problem of an economic agent where preferences, beliefs and decision rules are specified and shows the kinds of restrictions that are required, especially on the distribution of prior beliefs, to yield the exogeneity assumptions we use.
+The structural model provides a coherent dynamic decision problem of an economic agent where preferences, beliefs and decision rules are specified and shows the kinds of restrictions that are required, especially on the distribution of prior beliefs, to yield the sequential exchangeability conditions we use.
 Thus, our work connects to recent work on identification of learning models in nonparametric settings (e.g., (Bunting, Diegert, and Maurel, 2022)).
-Also, we connect a conditional independence assumption in the learning model to common parallel trends assumptions and show how sequential random assignment of treatment can be compatible with sequential exogeneity as expected but not parallel trends.
+This exercise is similar to the one done in Marx, Tamer, and Tang (2024) for a model under parallel trends.
 Our work also relates to a broader agenda relating sequential exchangeability and parallel trends (e.g., Renson, Hudgens, Keil, Zivich, and Aiello (2023)) and the corresponding estimators, for example through “bracketing” relationships on the target parameter (Angrist and Pischke, 2008; Ding and Li, 2019).
@@
 In complementary work, Kim (2023) studies identification of a dynamic panel data model when treatment is staggered and potential outcomes are non-negative, statically indexed, and take a multiplicative form; in this case, he obtains identification results for ratios of treatment effects under sequential exchangeability conditional on persistent heterogeneity.
+Finally, Klosin (2024) provides an interesting approach that corrects for the bias caused by ignoring dynamic feedback in static panel data regressions.
 The rest of the paper is organized as follows.
 Section 2 introduces the dynamic potential outcome framework.
-Section 3 studies the causal interpretation of DPDEs in the econometrics literature.
-Section 4 studies sequential exchangeability in the context of a model of dynamic choice with learning and illustrates the compatibility of parallel trends assumptions with conditional independence restrictions within the learning model and random assignment of treatments.
+Section 3 studies how to use DPDEs to make causal inference about intertemporal heterogeneous TEs in this framework, with and without the sequential exchangeability condition.
+Section 4 investigates when sequential exchangeability can be justified in an (observational) context of dynamic treatment choices with learning, and an (experimental) setting of sequential randomization of treatments.
 Section 5 concludes.
+Proofs are collected in Appendix A.
 2 A Dynamic Framework for Potential Outcomes
@@
 There is a pre-treatment period $t=0$ with an initial realized outcome $Y_{i0}$ and $\Pr\{D_{i0}=0\}=1$.
-We take $Y_{i0}$ as given and unmodeled, thus avoiding the initial conditions problem (Heckman, 1981).
-Throughout what follows, we consider models in which potential outcomes in period $t$ are functions only of an individual’s own treatment history through period $t$.
+We do not model how $Y_{i0}$ is determined prior to the sampling period, but allow it to be correlated with potential outcomes and treatment choices in our analysis.
+Potential outcomes in period $t\geq 1$ are functions only of an individual’s own treatment history through period $t$.
 Assumption 1
-Potential outcomes for individual unit $i$ in period $t$ are indexed only by $i$’s own treatment through that time period, i.e., $Y_{it}(d_{i}^{t})$.
+Potential outcomes for individual unit $i$ in period $t$ are indexed only by $i$’s own treatment through that time period, i.e., $Y_{it}(d_{i}^{t})$, with $d^{t}_{i}\equiv(d_{is})_{s=1}^{t}$.
 Assumption 1 rules out spillovers across individual units as well as anticipation of future treatments affecting present potential outcomes.
-Realized outcomes equal potential outcomes evaluated at realized treatments, $Y_{it}=Y_{it}(D_{i}^{t})$.
-We adopt notation $Y_{i}^{t}\equiv(Y_{i0},Y_{i1},\dots,Y_{it})$ for realized random (outcome) vectors, $Y_{i}^{t}(d^{t})=(Y_{i0},Y_{i1}(d^{1}),\dots,Y_{it}(d^{t}))$ for potential random vectors, and $Y_{i}^{t}(\cdot)$ for the collection of potential random vectors through period $t$.
+Observed outcomes equal potential outcomes evaluated at observed treatments, $Y_{it}=Y_{it}(D_{i}^{t})$.
+In this framework, the treatment effects $Y_{it}(d^{t})-Y_{it}(\widetilde{d}^{t})$ are intertemporal in that they are defined by the full history of past potential treatments.
+We adopt notation $Y_{i}^{t}\equiv(Y_{i0},Y_{i1},\dots,Y_{it})$ for the history of realized outcomes, $Y_{i}^{t}(d^{t})\equiv(Y_{i0},Y_{i1}(d^{1}),\dots,Y_{it}(d^{t}))$ for that of potential random outcomes, and $Y_{i}^{t}(\cdot)\equiv\{Y_{i}^{t}(d^{t}):d^{t}\in\{0,1\}^{t}\}$, i.e., the collection of potential random vectors through period $t$.
 Throughout, we restrict to realized and potential treatment vectors $D_{i}^{t}$ and $d_{i}^{t}$ with degenerate initial realization $D_{i0}=d_{i0}=0$, which we suppress where convenient.
-The main concern that motivates the inclusion of lagged outcomes as controls in a dynamic panel data model (DPDM) for observed outcomes in (1) is that the treatment is (mean) independent of potential outcomes only after conditioning on the observed history of past treatment and outcome.
+The main concern that motivates the inclusion of lagged outcomes as controls in a dynamic panel data model (DPDM) for observed outcomes in (1) is that the potential outcomes are (mean) independent from the treatments only after conditioning on the observed history of past treatments and outcomes.
 The assumption below formalizes such conditional independence.
@@
     $E[Y_{t}(d^{t})|Y^{s-1},D^{s-1}=d^{s-1},D_{s}=d_{s}^{\prime}]\quad\text{is invariant in $d_{s}^{\prime}$.}$
-Assumption 2 in either form is based on the sequential exchangeability of Robins (1986) in the case where the covariate history consists only of lagged outcomes.222 Other names for Assumption 2b in the literature include “conditional exchangeability”, “sequential ignorability”, and “conditional parallel trends”.
-Assumption 2 only restricts a stream of potential outcomes where the first $s-1$ treatment indices conform to the realized history of treatments $D^{t-1}=d^{t-1}$; it imposes conditional independence between present treatment for the entire stream of counterfactual potential outcomes in the future.
+Assumption 2 in either form is based on the sequential exchangeability condition in Robins (1986) when the covariate history consists only of lagged outcomes.444 Other names for Assumption 2b in the literature include “conditional exchangeability”, “sequential ignorability”, and “conditional parallel trends”.
+Assumption 2 only restricts a stream of potential outcomes where the first s1s-1 treatment indices conform to the realized history of treatments $D^{t-1}=d^{t-1}$; it imposes conditional independence between present treatment and the entire stream of counterfactual potential outcomes in the future.
 Mean sequential exchangeability (Assumption 2b) relaxes full sequential exchangeability (Assumption 2a) in two ways.
@@
 It is well-known that sequential exchangeability is sufficient for identifying a variety of treatment effects (e.g, Robins and Hernán (2008)).
-Instead, our focus will be to show how this condition also leads to causal interpretations of alternative estimands in dynamic panel data models for observed outcomes.
+Instead, our focus will be to show how this condition also leads to causal interpretations of alternative estimands in the dynamic panel data models for observed outcomes.
 Remark 1
@@
 It is straightforward to illustrate Proposition 1 when $T=2$, in which case the alternative trend formulation is: for any $(d_{1},d_{2})\in\{0,1\}^{2}$,
-$E[Y_{1}(d_{1})-Y_{0}|Y_{0},D_{1}]$ $=E[Y_{1}(d_{1})-Y_{0}|Y_{0}];$ (4)
-$E[Y_{2}(d_{1},d_{2})-Y_{1}(d_{1})|Y_{0},D_{1}]$ $=E[Y_{2}(d_{1},d_{2})-Y_{1}(d_{1})|Y_{0}];$ (5)
-$E[Y_{2}(d_{1},d_{2})-Y_{1}(d_{1})|Y_{0},Y_{1},D_{1}=d_{1},D_{2}]$ $=E[Y_{2}(d_{1},d_{2})-Y_{1}(d_{1})|Y_{0},Y_{1},D_{1}=d_{1}].$ (6)
+$E[Y_{1}(d_{1})-Y_{0}|Y_{0},D_{1}]$ $=E[Y_{1}(d_{1})-Y_{0}|Y_{0}];$
+$E[Y_{2}(d_{1},d_{2})-Y_{1}(d_{1})|Y_{0},D_{1}]$ $=E[Y_{2}(d_{1},d_{2})-Y_{1}(d_{1})|Y_{0}];$
+$E[Y_{2}(d_{1},d_{2})-Y_{1}(d_{1})|Y_{0},Y_{1},D_{1}=d_{1},D_{2}]$ $=E[Y_{2}(d_{1},d_{2})-Y_{1}(d_{1})|Y_{0},Y_{1},D_{1}=d_{1}].$
 It is clear from $T=2$ that in one way, sequential exchangeability is strong in that it is required to hold for all potential outcomes, both treated and untreated, whereas the parallel trends assumption only restricts the trends in untreated outcomes.
@@
 3 Causal Inference via Dynamic Panel Data Methods
-3.1 Dynamic Panel Data Models
-Canonical dynamic panel data methods are rooted in linear models with lagged outcomes as a regressor.
-In our context, we consider causal interpretations of a dynamic panel data model (DPDM) for observed outcomes:
+3.1 DPDM for Observed Outcomes
+We consider causal interpretations of an estimand motivated by an instrument-based method for estimating a dynamic panel data model (DPDM) for observed outcomes:
 $Y_{it}=\beta D_{it}+\gamma Y_{it-1}+\theta_{t}+\alpha_{i}+\varepsilon_{it},$ (3)
 where $\gamma,\beta$ are constant parameters, $\alpha_{i}$ are time-invariant individual fixed effects, $\theta_{t}$ are deterministic time trends, and $\varepsilon_{it}$ are time-varying idiosyncratic errors.
-Specifically, we investigate how the estimated coefficients can be related to the moments of heterogeneous treatment effects and trends in potential outcomes.
+We investigate how the estimated coefficients can be related to the moments of heterogeneous treatment effects and trends in potential outcomes.
 Equation (3) is natural for empiricists as a reduced-form model for a series of observed outcomes.
@@
 Anderson and Hsiao (1982) and Arellano and Bond (1991) proposed consistent estimators for the coefficients in (3), using moments implied by instruments such as the lagged outcomes and covariates.
-Next, we turn to the interpretation of such estimators.
-It is important to emphasize, however, that we do not interpret the DPDM for observed outcomes (7) as a causal model per se.
-Rather, we consider its implicit interpretation and estimation under non-parametric assumptions on the potential outcome process.
-This exercise is analogous to the interpretation of TWFE regressions under parallel trend assumptions, e.g., De Chaisemartin and d’Haultfoeuille (2020) and Goodman-Bacon (2021).
-In Section 3.4 we also consider an analogous parametric model to (7), but as a model of the potential outcome generating process.
-3.2 Instrumented First-Difference Estimators
-Our goal in this section is to obtain a causal interpretation of estimators for (7) that address the endogeneity of lagged outcomes. First-difference transformation of (7) implies:
-$\Delta Y_{it}=\gamma\Delta Y_{it-1}+\beta\Delta D_{it}+\Delta\theta_{t}+\Delta\varepsilon_{it}\ \quad\text{for }2\leq t\leq T\text{,}$ (8)
+Specifically, first-differencing (3) implies:
+$\Delta Y_{it}=\beta\Delta D_{it}+\gamma\Delta Y_{it-1}+\Delta\theta_{t}+\Delta\varepsilon_{it}\ \quad\text{for }2\leq t\leq T\text{,}$ (4)
 where $\Delta$ denotes the first-difference operator, e.g., $\Delta Y_{it}\equiv Y_{it}-Y_{it-1}$.
 Anderson and Hsiao (1982) estimated the coefficients via IV regression, using $Y_{it-2}$ as instruments for $\Delta Y_{it-1}$ when $D_{it}$ is strictly exogenous.
-Arellano and Bond (1991) proposed a GMM estimator when $D_{it}$ is sequentially exogenous (a.k.a.
-pre-determined), using earlier lagged outcomes and treatments as instruments for $\Delta Y_{it-1}$ and $\Delta D_{it}$.
-We focus henceforth on the causal interpretation of a two-stage least squares (2SLS) estimator for $\beta$ when $T=2$.
-In this case, the coefficients $\rho$ and $\gamma$ are just identified, with $(Y_{i0},D_{i1})$ being instruments for $(\Delta Y_{i1},\Delta D_{i2})$ in (8), and 2SLS and GMM estimators are numerically equivalent.
-As before, we drop individual subscripts $i$ for simplicity.
+Arellano and Bond (1991) proposed a GMM estimator when $D_{it}$ is sequentially exogenous (a.k.a., pre-determined), using earlier lagged outcomes and treatments as instruments for $\Delta Y_{it-1}$ and $\Delta D_{it}$.
+It is important to emphasize that we do not interpret the DPDM for observed outcomes (3) as a causal model per se.
+Rather, our goal is to obtain a causal interpretation of the limits to which a first-differenced instrument-based estimator converge.
+We refer to these limits as the Arellano-Bond (first-differenced) instrumental-variable (IV) estimands of DPDMs (or DPDEs).
+We show how these estimands relate to the distribution of heterogeneous treatment effects under non-parametric assumptions on the potential outcome process.
+This exercise is analogous to the interpretation of TWFE regressions under parallel trend assumptions, e.g., De Chaisemartin and d’Haultfoeuille (2020) and Goodman-Bacon (2021).
+3.2 Causal Interpretation of First-Differenced IV Estimands
+We focus henceforth on the causal interpretation of IV estimands of DPDM in (3) when $T=2$, which are defined as follows (we drop individual subscripts $i$ for simplicity):
+$\begin{pmatrix}\widetilde{\beta}\\ \widetilde{\gamma}\\ \widetilde{\Delta\theta_{2}}\end{pmatrix}\equiv E[Z^{\prime}X]^{-1}E[Z^{\prime}\Delta Y_{2}],$ (5)
+where
+$X$ $\equiv(\Delta D_{2},\Delta Y_{1},1)$
+$Z$ $\equiv E[X|Y_{0},D_{1}]$
+are each row vectors.
+If equation (3) is the actual data-generating process (DGP) for the observed outcomes, $\widetilde{\beta}$ is the probability limit of an IV estimator for $\beta$ with $(E[\Delta D_{2}|Y_{0},D_{1}],E[\Delta Y_{1}|Y_{0},D_{1}])$ instrumenting for $(\Delta D_{2},\Delta Y_{1})$.555 This IV estimator would also be numerically equivalent to a 2SLS estimator that uses $(D_{1},Y_{0})$ as instruments for $(\Delta D_{2},\Delta Y_{1})$, if the conditional expectations of $\Delta D_{2},\Delta Y_{1}$ are linear in $(Y_{0},D_{1})$.
+With such a choice of instruments, the model is just-identified, with GMM/2SLS estimators numerically identical to the IV estimator.
+In addition, if the DGP further satisfies the identifying assumptions in Arellano and Bond (1991), such as exogeneity of $D_{it}$ and serial uncorrelation of reduced-form errors $\varepsilon_{it}$, then $\widetilde{\beta}$ would indeed coincide with $\beta$ in (3).
+A natural question to ask is the following: are there any DGPs of potential outcomes and treatments that would lead to a model for observed outcomes that conform to the functional form in (3) and satisfy the identifying conditions in Arellano and Bond (1991) (so that $\widetilde{\beta}=\beta$)?
+The answer is positive; we present it in Section 3.3.1.
+For the rest of Section 3.2, we address a different question, namely, whether the IV estimand $\widetilde{\beta}$ above admits a causal interpretation in general.
+We explore conditions under which it does.
+While doing so, we maintain that the DGP of observed outcomes does not necessarily conform to (3) or the identifying conditions in Arellano and Bond (1991).
 Define history-dependent trends in potential outcomes:
@@
 and the treatment effects:
-$\tau_{1}\equiv Y_{1}(1)-Y_{1}(0);\quad\tau_{2}(d_{1})\equiv Y_{2}(d_{1},1)-Y_{2}(d_{1},0).$ (10)
-The trend of observed outcomes can then be written as:
+$\tau_{1}\equiv Y_{1}(1)-Y_{1}(0);\quad\tau_{2}(d_{1})\equiv Y_{2}(d_{1},1)-Y_{2}(d_{1},0).$
+The trend of observed outcomes are decomposed as:
 $\Delta Y_{1}$ $=Y_{1}(D_{1})-Y_{0}=\delta_{1}+D_{1}\tau_{1}\text{;}$ (7)
 $\Delta Y_{2}$ $=Y_{2}(D_{1},D_{2})-Y_{1}(D_{1})=\delta_{2}(D_{1})+D_{2}\tau_{2}(D_{1})\text{.}$ (8)
-To fix ideas, suppose $Y_{0}$ takes discrete values.
-Let the first-stage regression be fully saturated, so that the first-stage fitted values of the trend in observed outcomes and treatments converge in probability to $E[\Delta Y_{1}|Y_{0},D_{1}]$ and $E[\Delta D_{2}|Y_{0},D_{1}]$ respectively.
-The next lemma invokes the Frisch-Waugh-Lovell Theorem and the decomposition of $\Delta Y_{2}$ in (12) to express the probability limit of the 2SLS estimator for $\beta$ in terms of causal and non-causal components.
-Let $w(Y_{0},D_{1})$ be the projection errors in the linear projection of $E(\Delta D_{2}|Y_{0},D_{1})$ on $E(\Delta Y_{1}|Y_{0},D_{1})$.
+The next lemma invokes the Frisch-Waugh-Lovell Theorem and the decomposition of $\Delta Y_{2}$ in (8) to express the IV estimand $\widetilde{\beta}$ in (5) in terms of causal and non-causal components.
+Let $w(Y_{0},D_{1})$ be the errors in the linear projection of $Z_{1}\equiv E(\Delta D_{2}|Y_{0},D_{1})$ onto $Z_{-1}\equiv(E(\Delta Y_{1}|Y_{0},D_{1}),1)$:
+$w(Y_{0},D_{1})\equiv Z_{1}-Z_{-1}E[Z_{-1}^{\prime}Z_{-1}]^{-1}E[Z_{-1}^{\prime}Z_{1}].$ (9)
 Assume $w(Y_{0},D_{1})$ is not degenerate at zero, which is an innocuous, empirically verifiable condition.
 Lemma 1
-Suppose Assumption 1 holds and $w(Y_{0},D_{1})$ is not degenerate at zero.
-Let the first-stage regressions be fully saturated.
-The 2SLS estimator of $\beta$ in (8), which uses $(Y_{0},D_{1})$ as instruments for $(\Delta Y_{1},\Delta D_{2})$, converges in probability to:
+Suppose Assumption 1 holds, and $w(Y_{0},D_{1})$ as defined in (9) is not degenerate at zero. The IV estimand $\widetilde{\beta}$ in (5) is:
 $\frac{E[w(Y_{0},D_{1})\Delta Y_{2}]}{E[w(Y_{0},D_{1})^{2}]}=E\left[\frac{w(Y_{0},D_{1})D_{2}}{E[w(Y_{0},D_{1})^{2}]}\times\tau_{2}(D_{1})\right]+E\left[\frac{w(Y_{0},D_{1})}{E[w(Y_{0},D_{1})^{2}]}\times\delta_{2}(D_{1})\right].$ (10)
-Lemma 1 clarifies the basis and impediments to causal interpretations of the 2SLS estimator with no other conditions than the parametrization of potential outcomes (Assumption 1).
+Lemma 1 clarifies the basis and impediments to causal interpretations of the IV estimand $\widetilde{\beta}$ with no other conditions than Assumption 1.
 It does not invoke sequential exchangeability in Assumption 2.
-The first term on the right in (13) is a weighted aggregate of treatment effects in period 2 among the treated ($D_{2}=1$), yet with no guarantee that the combination is convex or even positively weighted.333 This is analogous to a point by De Chaisemartin and d’Haultfoeuille (2020) and Goodman-Bacon (2021) in the causal interpretation of two-way fixed-effect (TWFE) regressions in static panel data models.
-The second term, if non-zero, is an additional confounder to causal interpretation in terms of treatment effects $\tau_{2}(d_{1})$.
-To endow the 2SLS estimand in (13) with a clearer causal interpretation, we require further restrictions on heterogeneous treatment effects in addition to those on selection in Assumption 2.
+The first term on the right in (10) is a weighted aggregate of individually heterogeneous treatment effects $\tau_{2}(\cdot)$ in the second period among the treated ($D_{2}=1$), yet with no guarantee that the combination is convex or even positively weighted.666 This is analogous to a point by De Chaisemartin and d’Haultfoeuille (2020) and Goodman-Bacon (2021) in the causal interpretation of two-way fixed-effect (TWFE) regressions in static panel data models.
+The second term, if non-zero, is an additional confounder to causal interpretation in terms of treatment effects $\tau_{2}(\cdot)$.
+To endow this decomposition of the IV estimand $\widetilde{\beta}$ in (10) with a clearer causal interpretation, we require further restrictions on heterogeneous treatment effects in addition to those on selection in Assumption 2.
 Assumption 3 (Limited Heterogeneity of Trends and Treatment Effects)
@@
 $E[\tau_{2}(d_{1})|Y_{0},Y_{1}(d_{1}^{\prime})]$ $=E[\tau_{2}(d_{1})].$ (14)
-Note first that each of the conditions in Assumption 3 would be satisfied in the special case where treatment effects $\tau_{t}(\cdot)$ and untreated trends $\delta_{t}(\cdot)$ are homogeneous in the population.
-Under Assumption 3, the next proposition shows that the 2SLS first-difference estimator recovers a convex combination of history-dependent treatment effects.
+Assumption 3 requires the first moments of treatment effects $\tau_{t}(\cdot)$ and untreated trends $\delta_{t}(\cdot)$ to be invariant in the initial condition or earlier potential outcomes; obviously, it holds when these treatment effects and untreated trends are homogeneous in the population.
+The next proposition shows that under Assumption 3 the IV first-difference estimand recovers a convex combination of history-dependent treatment effects.
 Proposition 2
-Suppose Assumptions 1, 2a, 3 hold, and $w(Y_{0},D_{1})$ is not degenerate at zero. The 2SLS estimator of $\beta$ in (8) converges in probability to a convex combination of treatment effects:
+Suppose Assumptions 1, 2a, 3 hold, and $w(Y_{0},D_{1})$ is not degenerate at zero. Then the IV estimand $\widetilde{\beta}$ in (5) is a convex combination of treatment effects:
 $E\left[\frac{w(Y_{0},D_{1})^{2}}{E[w(Y_{0},D_{1})^{2}]}\times\tau_{2}(D_{1})\right],$ (15)
@@
 $w(Y_{0},D_{1})=E[D_{2}|Y_{0},D_{1}]-E[D_{2}|D_{1}].$ (16)
-This proposition confirms that an intuitive causal interpretation of the 2SLS estimand for the coefficient of $\Delta D_{2}$ exists under sufficiently restrictive conditions.
-A proof is presented in Appendix A.
-Intuitively, the proof proceeds in the three steps.
-First, derive the expression in (19) for the projection errors.
+This proposition confirms that an intuitive causal interpretation of the IV estimand for the coefficient of $\Delta D_{2}$ exists under additional conditions on individual heterogeneity.
+Intuitively, the proof (presented formally in Appendix A) proceeds in three steps.
+First, derive the expression of the projection errors in (16).
 Next, eliminate the confounding (second) term in (10).
 Finally, establish the convex weights in (15).
-A technical note worth mentioning is that only the mean version of sequential exchangeability (Assumption 2b) is invoked in the first and second steps of the proof; in the third step, the full version (Assumption 2a) is required in period $t=1$ to separate restrictions on selection from restrictions on heterogeneity.
+Only the mean version of sequential exchangeability (Assumption 2b) is invoked in the first and second steps; in the third step, the full version (Assumption 2a) is required in period $t=1$ to separate restrictions on selection from restrictions on heterogeneity.
 Proposition 2 complements the existing literature.
-It is well-known that the average treatment effect $E[\tau(d_{1})]$ is identified for $d_{1}\in\{0,1\}$ under sequential exchangeability in Assumption (2); see, e.g., an application of the $g$-formula in (Robins and Hernán, 2008).
-In comparison, our contribution is to show how to use an alternative 2SLS approach to estimate a different causal effect parameter, which is a weighted aggregate of treatment effects where the weights can be recovered from observable moments.
-3.3 An Adjusted Inverse-Propensity-Weighted Estimator
-The result in Proposition 2 naturally raises a new question: can we modify our approach to construct new estimators for other averages of causal effects, without restricting the heterogeneity of treatment effects as in Assumption 3?
-Our next result shows how to use an adjusted inverse-propensity-weighted (IPW) method to estimate an average treatment effect (ATE) upon transforming $\Delta Y_{2}$.
-The basis for this transformation is the same decomposition (12) invoked in the preceding subsection 3.2.
-In our dynamic potential outcome setting, treatment effects are also indexed by past treatment $\tau_{2}(d_{1})$.
-Thus, the ATE in period 2 that we target also averages over the observed past treatment $D_{1}$.
-This new (to the best of our knowledge) estimator does not invoke the restrictions on heterogeneous treatment effects and trends in Assumption 3 or the full version of sequential exchangeability of Assumption 2a; besides Assumptions 1 and 2b, it only requires a standard support assumption on treatment paths.
+It is well-known that the average treatment effect (ATE) $E[\tau_{2}(d_{1})]$ is identified for $d_{1}\in\{0,1\}$ under sequential exchangeability in Assumption (2); see, e.g., the $g$-formula in (Robins and Hernán, 2008).
+In comparison, in Proposition 2 we propose an alternative way to use the sequential exchangeability condition.
+By focusing only on causal interpretation of first-differenced IV estimators (instead of a more ambitious goal of estimating $E[\tau_{2}(d_{1})]$ for all $d_{1}$), we manage to obtain useful insights under additional mean restrictions on the heterogeneous treatment effects (Assumption 3).
+We conclude the section by noting that an outcome-modifying mean estimand identifies the average treatment effect $E[\tau_{2}(D_{1})]$.
+This is a different causal parameter than the one targeted in the $g$-formula.
+This result does not require the restrictions on heterogeneous treatment effects and trends in Assumption 3 or the full sequential exchangeability of Assumption 2a; besides Assumptions 1 and 2b, it only requires a standard support assumption on treatment paths below.
 Assumption 4 (Full Support of Treatment)
-$E[D_{2}|Y^{1},D_{1}]\in(0,1)$.
-Under Assumption 4, define:
+$E[D_{2}|Y^{1},D_{1}]\in(0,1)$ almost surely.
+Under Assumption 4, define:777 Note that $\widetilde{\Delta Y_{2}}$ is identical to an expression replacing terms $\Delta Y_{2}$ with $Y_{2}$.
+We preserve this redundant trend form in order to clarify the connection to our first-differenced IV estimand.
 $\widetilde{\Delta Y_{2}}\equiv\frac{\Delta Y_{2}-E[\Delta Y_{2}|Y^{1},D_{1},D_{2}=0]}{E[D_{2}|Y^{1},D_{1}].}$
+Relative to the decomposition (8), the transformation from $\Delta Y_{2}$ to $\widetilde{\Delta Y_{2}}$ in expectation removes the confounder $\delta_{2}(D_{1})$ through the term $E[\Delta Y_{2}|Y^{1},D_{1},D_{2}=0]$ and weights by the treated observations $E[D_{2}|Y^{1},D_{1}]$.
 The next proposition establishes that the mean of $\widetilde{\Delta Y_{2}}$ identifies the ATE in period 2, given the realized treatments in period 1.
@@
 Suppose Assumptions 1, 2b and 4 hold. Then $E[\widetilde{\Delta Y_{2}}]=E[\tau_{2}(D_{1})]$.
-There are two keys to the result.
-First, the mean of the trend $\delta_{2}(D_{1})$ in the decomposition of $\Delta Y_{2}$ in (12) can be identified and removed in the new estimand $E[\widetilde{\Delta Y_{2}}]$ under Assumption 2.
-Second, the use of inverse propensity weights allows us to attain a more interpretable causal parameter (ATE) without the weights in Proposition 2.
-By Proposition 3, we can estimate the ATE, $E[\tau_{2}(D_{1})]$, using sample analogs. Let $\widehat{M}_{2i}$ and $\widehat{M}_{1i}$ be consistent (e.g., kernel-based, Nadaraya-Watson) estimators for
-$M_{2i}\equiv E[\Delta Y_{2}|D_{2}=0,Y^{1}=(Y_{i0},Y_{i1}),D_{1}=D_{i1}]$
-and
-$M_{1i}\equiv E[D_{2}|Y^{1}=(Y_{i0},Y_{i1}),D_{1}=D_{i1}]$
-evaluated at any realized $(Y_{i0},Y_{i1},D_{i1})$. The analog estimator for $\mu_{\tau_{2}}\equiv E[\tau_{2}(D_{1})]$ is:
-$\widehat{\mu}_{\tau_{2}}\equiv\frac{1}{n}\sum\nolimits_{i=1}^{n}\left(\frac{\Delta Y_{i2}-\widehat{M}_{2i}}{\widehat{M}_{1i}}\right)\text{.}$ (20)
-In Appendix B, we show this adjusted IPW estimator $\widehat{\mu}_{\tau_{2}}$ is consistent for $\mu_{\tau_{2}}$.
-Robins (2000) introduced an IPW estimator for $E[Y_{t}(d^{t})]$ for any $d^{t}\in\{0,1\}^{t}$ in marginal structural models.
-Despite being motivated by his original IPW estimator, our adjusted IPW estimator targets a different causal parameter.
-Specifically, our estimand is $E[\tau_{2}(D_{1})]=E[Y_{2}(D_{1},1)-Y_{2}(D_{1},0)]$, while that of Robins’ IPW estimator is $E[Y_{2}(d_{1},d_{2})]$.
-Two further distinctions of our estimator are that it is based on the first-differenced outcome $\Delta Y_{2}$ and that it pools all observations in estimation.
-In addition to pooling in estimation, our target parameter provides a potentially useful summary causal effect in multi-period settings where the number of potential contrasts grows large.
-That is, $E[Y_{t}(D^{t-1},1)-Y_{t}(D^{t-1},0)]$ can be thought of as a weighted average of the $2^{t-1}$ average effects $E[Y_{t}(d^{t-1},1)-Y_{t}(d^{t-1},0)]$ that are in principle identified under sequential exchangeability.
-Alternatively, we can construct a numerically equivalent 2SLS estimator for $E[\tau_{2}(D_{1})]$ after transforming $\widetilde{\Delta Y_{2}}$ using the weights $W(Y_{0},D_{1})$. To see this, let
-$\Delta Y_{2}^{\dagger}\equiv\widetilde{\Delta Y_{2}}\times\frac{E[W^{2}]}{W},$
-and run a transformed 2SLS regression on $\Delta Y_{2}^{\dagger}$.
-First, regress the endogenous regressors in (8), $\Delta Y_{1}$ and $\Delta D_{2}$, on a fully saturated set of instruments $(Y_{0},D_{1})$ to obtain the fitted values.
-Then regress $\Delta Y_{2}^{\dagger}$ on the fitted values (and an intercept).
-It is clear from (13) in Lemma 1 that, provided the projection errors $W(Y_{0},D_{1})$ are nonzero almost surely, such an estimand is equal to $E[\widetilde{\Delta Y_{2}}]$, and therefore $E[\tau_{2}(D_{1})]$ by Proposition 3.
-In practice, we can implement this transformed 2SLS by first estimating $\Delta Y_{2i}^{\dagger}$ as the product of each individual summand in $\widehat{\mu}_{\tau_{2}}$ and its estimated weights $\frac{1}{n}\sum_{j}\widehat{W}_{j}^{2}/\widehat{W}_{i}$, where $\widehat{W}_{i}$ is the sample analog estimator for the projection errors $w(Y_{0i},D_{1i})$.
-3.4 Settings Without Sequential Exchangeability
-Sequential exchangeability (Assumption 2) has played a central role in the preceding results of Subsections 3.2 and 3.3.
-It seems natural to ask whether this assumption is necessary for causal inference in our setting.
-Indeed, the identification of linear dynamic panel data models with persistent unobserved heterogeneity as in (1) suggests it is not necessary per se, at least if replaced with other (e.g., parametric) assumptions.
-The goal of this subsection is to establish this in our potential outcome framework and show that for a class of dynamic data-generating processes, standard GMM-based DPDEs recover causal estimands even though sequential exchangeability is generically violated without conditioning on unobservables.
-Consider a specialized case in which the series of potential outcomes are structurally determined by a linear dynamic panel data model.
+Proposition 3 establishes the sample average of $\widetilde{\Delta Y_{2}}$ as a standalone estimator of the average treatment effect $\tau_{2}(D_{1})$ under the sequential exchangeability condition.
+A corollary of Lemma 1 and Proposition 3 is that an adjusted first-differenced IV estimand which replaces $\Delta Y_{2}$ with:
+$\Delta Y_{2}^{\dagger}\equiv\widetilde{\Delta Y_{2}}\times\frac{E[w(Y_{0},D_{1})^{2}]}{w(Y_{0},D_{1})}$
+also identifies $E[\tau_{2}(D_{1})]$. Unlike Proposition 2, this result holds even without Assumption 3 limiting heterogeneity.
+3.3 Causal Inference Without Sequential Exchangeability
+Sequential exchangeability (Assumption 2) played a central role in Section 3.2.
+Yet this condition does not hold generally in observational settings where selection into realized treatments depends on unobserved heterogeneity correlated with contemporary potential and past observed outcomes.
+(We provide a detailed discussion on this subject in Section 4.2.)
+It is therefore natural to ask whether we can recover causal parameters without sequential exchangeability.
+In this section, we explore other ways to use dynamic panel data models for causal inference where sequential exchangeability does not hold, and yet certain conditional average treatment effects remain (point) identifiable under auxiliary assumptions.
+3.3.1 AR(1) Potential Outcome Model
+Recall that the dynamic panel data model for observed outcomes in (1) is identified under the conditions of Arellano and Bond (1991), including sequential exogeneity of $D_{it}$ and the serial uncorrelation of the reduced-form errors $\varepsilon_{it}$.
+Our goal in Section 3.3.1 is to provide a set of sufficient assumptions on the distribution of potential outcomes and treatments, so that the implied DGP of observed outcomes conform to the functional form of DPDM in (3) and satisfy the conditions in Arellano and Bond (1991).
+The assumptions we introduce below include homogeneous treatment effects and (conditionally) serially uncorrelated structural errors in potential outcomes.
+Under such assumptions, the first-differenced IV estimand in (5) recovers the homogeneous treatment effects, even when sequential exchangeability does not hold.
+Consider a data-generating process in which the series of potential outcomes follow an AR(1) model with homogeneous causal effects.
+Assumption 5 (ARPO)
+Potential outcomes follow an AR(1) model:
+$Y_{it}(d^{t})=\beta^{*}d_{t}+\gamma^{*}Y_{it-1}(d^{t-1})+\theta_{t}^{*}+\alpha_{i}^{*}+\varepsilon_{it}^{*}(d^{t})\text{ for }t\geq 1,$ (17)
+with initial condition $Y_{i0}(d^{0})\equiv Y_{i0}$, [^8] Let $d^{0}$ denote the null vector so that $Y_{i0}(d^{0})\equiv Y_{i0}$ conforms with the notation $Y_{it}(d^{t})\equiv Y_{it}(d_{1},...,d_{t})$.
+where $\beta^{*},\gamma^{*}$, and $\theta_{t}^{*}$ are constants, $\alpha_{i}^{*}$ are individual fixed effects, and $\varepsilon_{it}^{*}(d^{t})$ are time- and history-varying noises satisfying:
+$E[\varepsilon_{it}^{*}(\cdot)|D_{i}^{t},Y_{i0},\varepsilon^{*t-1}_{i}(\cdot),\alpha_{i}^{*}]=0$ (18)
+for all $t\geq 1$, where $\varepsilon_{i}^{*t}(d^{t})\equiv(\varepsilon_{i1}^{*}(d^{1}),\dots,\varepsilon_{it}^{*}(d^{t}))$ for $t\geq 1$ and $\varepsilon_{i}^{*t}(\cdot)$ is the collection of potential error vectors with $\varepsilon^{*0}_{i}(\cdot)$ a vacuous null vector.
+Unlike $\gamma$ in the DPDM for observed outcomes in (3), the coefficient $\gamma^{*}$ for potential outcomes in (17) has a built-in causal interpretation: it measures how changes in present potential outcomes depend on counterfactual changes in lagged (past) potential outcomes, which depend recursively on the full counterfactual treatment history $d^{t-1}$.
+The condition (18) on the structural errors $\varepsilon_{it}^{*}(\cdot)$ combines sequential exogeneity and serial uncorrelation, conditional on individual fixed effects.
+Note Assumption 5 is also consistent with the condition on potential outcomes in Assumption 1.
+The observation below connects the structural model for potential outcomes in (17) to the reduced-form DPDM for observed outcomes in (3).
+Observation 1
+Suppose the data-generating process for potential outcomes satisfies Assumption 5, and the observed outcomes are $Y_{it}=Y_{it}(D_{i}^{t})$.
+Then the observed sequence $(D_{i}^{t},Y_{i}^{t})$ satisfies (3) with $(\gamma,\beta,\theta_{t},\alpha_{i})=(\gamma^{*},\beta^{*},\theta_{t}^{*},\alpha_{i}^{*})$ and $\varepsilon_{it}=\varepsilon_{it}^{*}(D_{i}^{t})$.
+Our next proposition shows Assumption 5 implies a “conditional” version of mean sequential exchangeability that controls for the fixed effects.
+Proposition 4
+Under Assumption 5, mean sequential exchangeability (Assumption 2b) holds conditional on the fixed effect:
+$E[Y_{it}(d^{t})|D_{is}=d_{s}^{\prime},Y_{i}^{s-1},D_{i}^{s-1}=d^{s-1},\alpha_{i}^{*}]\quad\text{is invariant in $d_{s}^{\prime}$ for $t\geq s\geq 1$}$ (19)
+for all $d^{t}\in\{0,1\}^{t}$, and where $d^{s-1}$ is a subvector of the first s1s-1 terms of $d^{t}$.
+Even under Assumption 5, the conditional mean in (19) could be non-degenerate in individual fixed effect $\alpha_{i}^{*}$, whose distribution conditional on history $Y_{i}^{s-1},D_{i}^{s}$ may depend on the most recent treatment $D_{is}$.
+In such cases, Assumption 2b does not hold in general, because both potential outcomes and contemporary treatments depend on individual fixed effects in non-trivial ways.
+Despite possible violation of sequential exchangeability, Assumption (5) is sufficient for consistent estimation of the structural parameters $(\beta^{*},\gamma^{*},\Delta\theta_{2}^{*})$ using the Arellano-Bond IV estimand in (5), as applied to a reduced-form DPDM for observed outcomes in (3).
+Proposition 5
+[…]
+Generalization to the identification of later trends $\Delta\theta^{*}_{t}$ for $t\geq 3$ is straightforward; it only requires using $X_{it}\equiv(\Delta D_{it},\Delta Y_{it-1},1)$ and $Z_{it}\equiv E(X_{it}|Y_{i}^{t-2},D_{i}^{t-1})$ in the first-differenced IV estimand.
+The intuition for Proposition 5 is that (18) implies the standard Arellano-Bond assumptions on reduced-form errors $\varepsilon_{it}$ are satisfied. The result then follows from Observation 1.
+In summary, under Assumptions 5, a first-differenced IV estimand recovers the causal estimands (Proposition 5), even when sequential exchangeability is violated without conditioning on unobservables (Proposition 4).
+Unlike Section 3.2, this result does require a structural model of potential outcomes in (17) as well as econometric restrictions on the errors in potential outcomes.
+3.3.2 AR(1) Untreated Potential Outcome Model
+In this subsection, we obtain stronger and more robust results for identifying causal effects, generalizing from the homogeneous treatment effects in the ARPO model (Assumption 5).
+Specifically, we let the treatment effects be heterogeneous, time-varying, and history-dependent:
+$\tau_{it}^{*}(d^{t-1})\equiv Y_{it}(d^{t-1},1)-Y_{it}(d^{t-1},0)$ (20)
+and impose the structure and sequential exogeneity only on untreated potential outcomes.
+We refer to this as an AR(1) untreated potential outcomes (ARUPO) model.
+Besides the heterogeneous treatment effects, we recycle coefficient labeling to avoid introducing new notation.
+Assumption 6 (ARUPO)
+Untreated potential outcomes follow an AR(1) model:
+$Y_{it}(d^{t-1},0)=\gamma^{*}Y_{it-1}(d^{t-1})+\alpha^{*}_{i}+\theta^{*}_{t}+\varepsilon^{*}_{it}(d^{t-1})\text{ for }t\geq 1,$ (21)
+with initial condition $Y_{i0}(d^{0})\equiv Y_{i0}$, where $\gamma^{*}$ and $\theta_{t}^{*}$ are constants, $\alpha_{i}^{*}$ are individual fixed effects, and $\varepsilon_{it}^{*}(d^{t-1})$ are time- and history-varying noises in potential outcomes satisfying:
+$E[\varepsilon^{*}_{it}(d^{t-1})|D_{i}^{t}=d^{t},Y_{i0},\tau_{i}^{*t-1}(d^{t-2}),\varepsilon_{i}^{*t-1}(\cdot),\alpha^{*}_{i}]=0,$ (22)
+for all $t\geq 1$ and $d^{t}\in\{0,1\}^{t}$.
+This condition allows $D_{it}$ to be a function of the most recent outcome $Y_{it-1}$ and the full past history $(D_{i}^{t-1},Y_{i}^{t-2})$.
+It posits the noises in potential outcomes $\varepsilon^{*}_{it}(\cdot)$ are uncorrelated with the sequence of past treatments up to $D_{i}^{t}$ and all latent elements that determine the past observed outcomes up to $Y_{i}^{t-1}$.
+The structure of the ARUPO model (Assumption 6) implies the observed outcomes are generated by a linear dynamic panel data model with a random coefficient for contemporary treatment:
+$Y_{it}\equiv Y_{it}(D_{i}^{t})=\tau_{it}^{*}D_{it}+\gamma^{*}Y_{it-1}+\alpha_{i}^{*}+\theta_{t}^{*}+\varepsilon_{it}^{*},$
+where $\tau_{it}^{*}$ and $\varepsilon_{it}^{*}$ are shorthand for $\tau^{*}_{it}(D_{i}^{t-1})$ and $\varepsilon^{*}_{it}(D_{i}^{t-1})$, respectively. First-differencing the observed outcomes yields:
+$\Delta Y_{it}=\tau_{it}^{*}D_{it}-\tau_{it-1}^{*}D_{it-1}+\gamma^{*}\Delta Y_{it-1}+\Delta\theta_{t}^{*}+\Delta\varepsilon_{it}^{*}\text{ for }t\geq 2.$ (23)
+Under Assumption 6,
+$E(\varepsilon_{it}^{*}|D_{i}^{t},Y_{i}^{t-1})=0\text{ for all }t\geq 1,$
+because the observed history $Y_{i}^{t-1}$ is a function of $(D_{i}^{t-1},Y_{i0},\tau_{i}^{*t-1}(D_{i}^{t-2}),\varepsilon_{i}^{*t-1}(D_{i}^{t-2}),\alpha^{*}_{i})$.
+This implies the observed history of treatment and outcomes is sequentially exogenous:
+$E(\Delta\varepsilon_{it}^{*}|H_{it-1})=0,\text{ where }H_{it-1}\equiv(D_{i}^{t-1},Y_{i}^{t-2}).$ (24)
+This is reminiscent of the moment condition implied by sequentially exogenous regressors in Arellano and Bond (1991).
+However, the ARUPO model is a generalization over the DPDM for observed outcomes: the first-difference of observed outcomes in (23) involves random coefficients (heterogeneous treatment effects) $\tau_{it}^{*}D_{it}-\tau_{it-1}^{*}D_{it-1}$ instead of the product of a constant coefficient and $\Delta D_{it}$.
+We first focus on identifying the average treatment effect of the first-time-treated defined for $t\geq 1$:
+$ATFT_{t}(y_{0})\equiv E[\tau_{it}^{*}|D_{it}=1,\mathcal{H}_{t-1}(y_{0})].$ (25)
+where for simplicity of notation we let:
+$\mathcal{H}_{t-1}(y_{0})\equiv\{D_{i}^{t-1}=0^{t-1},Y_{i0}=y_{0}\}$
+denote the event of an untreated history among those with initial outcome $y_{0}$.
+In Appendix B.1 we extend our approach to also identify similar causal parameters with a general history of past treatments $d^{t-1}\neq 0^{t-1}$.
+To begin, note that if the untreated model parameters $(\gamma^{*},\Delta\theta_{t}^{*})$ could be identified, then the $ATFT_{t}(y_{0})$ would also be identified through (23) and (24) upon taking conditional expectations and simplifying:
+$E[\Delta Y_{it}|\mathcal{H}_{it-1}(y_{0})]=E[D_{it}|\mathcal{H}_{it-1}(y_{0})]ATFT_{t}(y_{0})+\gamma^{*}E[\Delta Y_{it-1}|\mathcal{H}_{it-1}(y_{0})]+\Delta\theta_{t}^{*}.$ (26)
+Yet, it is not apparent how to identify these untreated model parameters in isolation without further assumptions.
+We show how to identify $ATFT_{t}(y_{0})$, using a conditional Arellano-Bond IV estimand for (23) under an additional condition on how the observed treatments relate to latent components in the ARUPO model (Assumption 6), in particular the heterogeneous treatment effects.
 Assumption 7
-Potential outcomes follow a linear, homogeneous, AR(1) model:
-$Y_{it}(d^{t})=\beta^{*}d_{t}+\gamma^{*}Y_{it-1}(d^{t-1})+\theta_{t}^{*}+\alpha_{i}^{*}+\varepsilon_{it}^{*}(d_{t})\text{ for }t\geq 1,$ (21)
-with initial condition $Y_{i0}(d^{0})\equiv Y_{i0}$, and where $\beta^{*},\gamma^{*}$, and $\theta_{t}^{*}$ are constant coefficients, $\alpha_{i}^{*}$ are individual-specific fixed effects, and $\varepsilon_{it}^{*}(d_{t})$ are time-varying noises in potential outcomes.
-Unlike $\gamma$ in the dynamic panel data model for observed outcomes in (7), the coefficient $\gamma^{*}$ in this model for potential outcomes in (21) has an inherently causal interpretation: it measures how changes in present potential outcomes depend on counterfactual changes in lagged (past) potential outcomes, which depend recursively on the full counterfactual treatment history $d^{t-1}$.
-Note that the structure of Assumption 5 also implies the potential outcome parametrization of Assumption 1.
-The following observation connects the linear dynamic panel data model for potential outcomes to the reduced-form model (7) for observed outcomes.
-Observation 1
-Suppose the data-generating process for potential outcomes satisfies Assumption 5, and the observed outcomes are $Y_{it}=Y_{it}(D_{it})$.
-Then the observed sequence $(D_{it},Y_{it})$ satisfies (7) with $(\gamma,\beta,\alpha_{i},\theta_{t})=(\gamma^{*},\beta^{*},\alpha_{i}^{*},\theta_{t}^{*})$ and $\varepsilon_{it}=\varepsilon_{it}^{*}(D_{it})$.
-This connection is useful because conditions for consistent estimation of the reduced-form, realized outcome model (7) are well-established.
-We focus on the Arellano-Bond (1991) GMM estimator, which requires the following set of assumptions on the reduced-form errors $\varepsilon_{it}$:444 Condition (24) is implicit in the moment condition in Arellano and Bond (1991) (Eq.
-(2) in Sec.
-2).
-$E(\varepsilon_{is}\varepsilon_{it})=0\text{ for all }t\neq s;$ (22)
-$E(D_{is}\varepsilon_{it})=0\text{ for all }s\leq t;$ (23)
-$E(\Delta\varepsilon_{it}\alpha_{i})=E(\Delta\varepsilon_{it}Y_{i0})=0\text{ for }t\geq 2.$ (24)
-In turn, a sufficient condition on the structural errors $\varepsilon_{it}^{*}(\cdot)$ combines sequential exogeneity and serial uncorrelation, conditional on individual fixed effects.
-For simplicity, we term this strong structural sequential exogeneity.
-Let $\varepsilon_{i}^{*t}(d^{t})\equiv(\varepsilon_{i1}^{*}(d_{1}),\dots,\varepsilon_{it}^{*}(d_{t}))$ for $t\geq 1$ and $\varepsilon_{i}^{*t}(\cdot)$ the collection of potential error vectors, with $\varepsilon^{*0}_{i}(\cdot)$ a vacuous null vector.
-Assumption 6 (Strong Structural Sequential Exogeneity)
-The potential error terms defined under Assumption 5 satisfy:
-$E[\varepsilon_{it}^{*}(\cdot)|D_{i}^{t},\varepsilon^{*t-1}_{i}(\cdot),Y_{i0},\alpha_{i}^{*}]=0\text{ for all }t\geq 1.$ (25)
-Sufficiency for the Arellano-Bond econometric assumptions is formalized in the following proposition:
+For any initial condition $y_{0}$,
+$E[\tau^{*}_{it}(0^{t-1})|D_{it},\mathcal{H}_{t-1}(y_{0}),\varepsilon_{i}^{*t-1}(\cdot),\alpha^{*}_{i}]=E[\tau^{*}_{it}(0^{t-1})|D_{it},\mathcal{H}_{t-1}(y_{0})].$
+Remarkably, this condition does not impose a notion of conditional unconfoundedness, because it does not rule out the correlation between the current treatment $D_{it}$ and the heterogeneous treatment effects $\tau^{*}_{it}(0^{t-1})$, which contribute to the contemporary potential outcome $Y_{it}(0^{t-1},1)$.
+Instead, it only states the fixed effect $\alpha^{*}_{i}$ and past noises $\varepsilon_{i}^{*t-1}(\cdot)$ do not affect the mean treatment effect conditional on the observed history.
+Of course, this assumption is also trivially satisfied in the ARPO model (Assumption 5) where the treatment effects are homogeneous.
+To define (a family of) conditional IV estimands, let $w_{t-1}(\cdot)$ be a chosen function of $H_{it-1}$ and write $W_{it-1}\equiv w_{t-1}(H_{it-1})$.
+Next, fix an initial condition value $Y_{0i}=y_{0}$ and consider an IV regression of $\Delta Y_{it}$ on $X_{it}\equiv(\Delta D_{it},\Delta Y_{it-1},1)$ conditional on the event $\mathcal{H}_{t-1}(y_{0})$, using the following instruments:
+$Z_{it}(y_{0})\equiv E[X_{it}|W_{it-1},\mathcal{H}_{t-1}(y_{0})].$
+These instruments are exogenous, because (24) implies $E[\Delta\varepsilon_{it}^{*}|Z_{it},\mathcal{H}_{t-1}(y_{0})]=0$. The estimand in such a (conditional) IV regression is:
+$\begin{pmatrix}\widetilde{\beta}_{t}(y_{0})\\ \widetilde{\gamma}_{t}(y_{0})\\ \widetilde{\Delta\theta_{t}}(y_{0})\end{pmatrix}\equiv E[Z_{it}^{\prime}(y_{0})X_{it}|\mathcal{H}_{t-1}(y_{0})]^{-1}E[Z_{it}^{\prime}(y_{0})\Delta Y_{it}|\mathcal{H}_{t-1}(y_{0})]$ (27)
+Our next result shows that under Assumptions 6 and 7, the IV regression in (27) recovers the causal parameter of interest, $ATFT_{t}(y_{0})$, as well as the (constant) state dependence parameter $\gamma^{*}$ and the time trends $\Delta\theta_{t}^{*}$.
 Proposition 6
-Under Assumptions 5 and 6, the Arellano-Bond econometric assumptions (22), (23), and (24) are satisfied.
-The next proposition shows that, within the linear dynamic model for potential outcomes (Assumption 5), strong structural sequential exogeneity (Assumption 6) implies a “conditional” version of mean sequential exchangeability that controls for the fixed effects.
-Proposition 5
-Under Assumptions 5 and 6, mean sequential exchangeability holds conditional on the fixed effect:
-$E[Y_{it}(d^{t})|D_{is}=d_{s}^{\prime},Y_{i}^{s-1},D_{i}^{s-1}=d^{s-1},\alpha_{i}^{*}]\quad\text{is invariant in $d_{s}^{\prime}$ for $t\geq s\geq 1$}$ (26)
-for all $d^{t}\in\{0,1\}^{t}$, and where $d^{s-1}$ is a subvector of the first $s-1$ terms of $d^{t}$.
-A corollary of Proposition 5 is that the mean sequential exchangeability in Assumption 2b would not hold in general, if the conditional mean in (26) is non-degenerate in individual fixed effect $\alpha_{i}^{*}$, and if the distribution of $\alpha_{i}^{*}$ conditional on history $Y_{i}^{s-1},D_{i}^{s}$ is not invariant in the most recent treatment $D_{is}$.
-In other words, the unconditional mean sequential exchangeability condition on potential outcomes would not hold in the context of the linear, structural model of Assumption 5 if both potential outcomes and contemporary treatments depend on individual fixed effects in non-trivial ways.
-Summarizing, under Assumptions 5 and 6, standard GMM-based DPDEs recover causal estimands (Proposition 4) even though sequential exchangeability is generically violated without conditioning on unobservables (Proposition 5).
-It is important to reiterate that our results in Section 3.2 and 3.3 do not require such a specific structural model of potential outcomes or econometric restrictions on the errors.
+Suppose Assumptions 6 and 7 hold, and $E[Z_{it}^{\prime}(y_{0})Z_{it}(y_{0})|\mathcal{H}_{t-1}(y_{0})]$ has full rank. Then the IV estimands (27) for $t\geq 3$ recover the causal terms:
+$\widetilde{\beta}_{t}(y_{0})$ $=ATFT_{t}(y_{0});$
+$\widetilde{\gamma}_{t}(y_{0})$ $=\gamma^{*};$
+$\widetilde{\Delta\theta_{t}}(y_{0})$ $=\Delta\theta_{t}^{*}.$
+Both $\gamma^{*}$ and $\Delta\theta_{t}^{*}$ are over-identified since the estimands are functions of $y_{0}$.
+The intuition for Proposition 6 is as follows.
+Under Assumption 7, the heterogeneous treatment effects $\tau^{*}_{it}$ are mean-independent from $Y_{i}^{t-2}$ conditional on $D_{it}$ and $\mathcal{H}_{t-1}(y_{0})$.
+This implies the conditional mean of $\Delta Y_{it}$ given $W_{it-1}$ and $\mathcal{H}_{t-1}(y_{0})$ is linear in the instruments $Z_{it}(y_{0})$, with slope coefficients $ATFT_{t}(y_{0})$, $\gamma^{*}$, and an intercept $\Delta\theta_{t}^{*}$.999 That is, an equality similar to (26) holds after further conditioning on $W_{it-1}$, which is excluded from $ATFT_{t}(y_{0})$ by Assumption 7.
+As a result, these parameters are recovered via an OLS regression of $\Delta Y_{it}$ on $Z_{it}(y_{0})$ conditional on $\mathcal{H}_{t-1}(y_{0})$.
+By construct, such an estimand also coincides with that of a conditional IV regression in (27), because $E[Z_{it}^{\prime}(y_{0})X_{it}(y_{0})|\mathcal{H}_{t-1}(y_{0})]=E[Z_{it}^{\prime}(y_{0})Z_{it}(y_{0})|\mathcal{H}_{t-1}(y_{0})]$ by the law of iterated expectations.
+In Appendix B.1 we generalize Assumption 7 and Proposition 6 to recover the average treatment effects beyond the first-treated:
+$ATT_{t}(d^{t-1},y_{0})\equiv E[\tau^{*}_{it}(d^{t-1})|D_{it}=1,D_{i}^{t-1}=d^{t-1},Y_{i0}=y_{0}]\text{ for }d^{t-1}\neq 0^{t-1}.$
+A natural further generalization of the ARUPO model in Assumption 6 would be to replace the constant $\gamma^{*}$ with random coefficients $\gamma_{i}^{*}$, which vary by individual but not over time; in turn, a natural extension of (18) conditions exogeneity on the extended vector of fixed effects.
+Even so, the analysis of Chamberlain (2022) suggests identification problems in this case of multiple random coefficients, with only partial identification of causal effects possible (e.g., see also Lee (2022)), even for the untreated outcome model in isolation.
+Therefore we do not consider this case further in this paper.
 4 Evaluating Sequential Exchangeability
-In this section, we assess the plausibility of sequential exchangeability (SE) in Assumption 2, which has played a central role in our results and discussion in Section 3.
-This condition amounts to a form of “selection on observables”, as it posits (mean or full) independence between potential outcomes and contemporary treatment conditional on observed history of past outcomes and treatments.
-Our objective is to provide guidelines for empirical researchers to decide when and how such restrictions on selection-on-observables can be justified in observational or experimental data.
-Naturally, justifying selection on observables in this dynamic setting would require restrictions on how the sequence of treatments are determined.
-In Section 4.1 and 4.2, we showcase the nature of these restrictions within a structural model where treatments are chosen by rational, forward-looking individual units with learning, and compare them with the parallel trends condition in the DiD literature.
-In Section 4.3, we show that sequential exchangeability holds in an experimental design of sequentially randomized treatment (where the treatment in period $t$ is randomized conditioning on the history of past treatments and outcomes); in contrast, the parallel trends conditions in general would not hold in such a case.
-4.1 A Structural Model of Treatment Choices with Learning
-We focus on a structural model of dynamic treatment choices that enables individual decision-makers to be both forward-looking and capable of learning.
-Learning is an interesting possibility in dynamic choice which is potentially compatible with sequential exchangeability assumptions because of the latter’s dependence on the history of observed outcomes.
-Unlike Section 3.4, the model studied in this section refrains from any parametric specification for the potential outcomes; instead, the model emphasizes scenarios in which forward-looking individuals make rational treatment decisions, continually learning from previous treatments and outcomes.
-To simplify the notation in this section, we omit the individual subscripts $i$.
-In each period $1\leq t\leq T$, the potential outcome is determined by:
-$Y_{t}(d^{t})=\lambda_{t}(d^{t},Y^{t-1}(d^{t-1}),\alpha(d_{t}),\varepsilon_{t}(d_{t}))\text{,}$ (27)
-where $Y^{s}(d^{s})=(Y_{0},Y_{1}(d^{1}),\dots Y_{s}(d^{s}))$ denotes a vector of counterfactual outcomes, with convention $Y^{0}(d^{0})=Y_{0}$.
-The initial condition $Y_{0}$ is allowed to be correlated with subsequent potential outcomes $Y_{t}(d^{t})$ for $t\geq 1$.
-We refer to $\alpha(\cdot)\equiv(\alpha(1),\alpha(0))$ as (stochastic) unobserved individual heterogeneity, or fixed effects, and $\varepsilon_{t}(\cdot)$ as potential, transitory noises affecting potential outcomes at time $t$.
-Individual decision makers do not observe $\alpha(\cdot)$ or $\varepsilon_{t}(\cdot)$, but know the deterministic function $\lambda_{t}(\cdot)$ as well as the distribution of $\varepsilon_{t}(\cdot)$ for all $t$555 Note that the structural dependence on the lagged outcome generalizes indexing of the full decision history because the outcome-generating function $\lambda_{t}(\cdot)$ is deterministic and depends only on present noise $\varepsilon_{t}(d_{t})$; in particular, the lagged outcome permits some dependence on the past noise vector $\varepsilon^{t-1}(d^{t-1})$.
-Each individual holds a prior belief for $\alpha(\cdot)$ at $t=1$, and uses incremental information to update their beliefs for $\alpha(\cdot)$ at the beginning of each period $t\geq 2$.
-Specifically, the individual’s prior over $\alpha(\cdot)$ at the start of $t=1$ is $\Psi_{1}(\cdot;\psi_{1})$, which is fully characterized by a random vector $\psi_{1}$.
-Let $\psi_{1}\equiv\varphi_{1}(H_{0},\xi_{0})$, where $H_{0}=Y_{0}$ denotes the observed history and $\xi_{0}$the unobserved individual factors in period 0 that determine the initial prior in period 1.
-In period $t\geq 2$, the posterior for $\alpha(\cdot)$ is $\Psi_{t}(\cdot;\psi_{t})$, which is completely characterized by an updated vector $\psi_{t}=\varphi_{t}(\psi_{t-1},H_{t-1},\xi_{t-1})$, where $H_{t}\equiv(D_{t},Y_{t})$ is the incremental information (observed by the researcher) that affects the individual’s posterior, and $\xi_{t}(d^{t})$ is the incremental information (unobserved by the researcher) from choice $d_{t}$ in period $t$, with $\xi_{t}=\xi_{t}(D^{t})$.
-Let $H^{t}\equiv(H_{0},\dots,H_{t})$ and $\xi^{t}=(\xi_{0},\dots,\xi_{t})$ denote the realized vectors and $\xi^{t}(\cdot)$ the full potential vector through period $t$.
-Using recursive substitution and with a slight abuse of notation, we can write: [^6] With $\Psi_{t}$ and $\varphi_{t}$ indexed by $t$, we allow for non-conjugate beliefs.
-Also, $\varphi_{t}$ may depend on unknown parameters, such as the variance of $\varepsilon_{t}(\cdot)$ in a normal learning model.
-$\psi_{t}=\varphi_{t}(H^{t-1},\xi^{t-1}).$
-To reiterate, we assume that $\psi_{t}$ is a sufficient statistic for the posterior of $\alpha(\cdot)$ at time $t$ given the outcome-relevant history $(H^{t-1},\xi^{t-1})$ observed by the decision-maker.
-Let $\mathcal{D}^{T}$ denote the set of possible sequences of choices; note that, in contrast to preceding sections, we do not assume that choices in each period are binary.
-Given a choice history $D^{t-1}=d^{t-1}$, an individual’s instantaneous, ex post utility from choosing $d_{t}$ at time $t$ is:
-$u_{t}\left(d_{t},Y_{t}(d^{t}),X_{t}\right)+\eta_{t}(d_{t})$
-where $X_{t}$ are additional states (apart from outcomes $Y_{t}$), and $\eta_{t}(\cdot)$ are potential shocks.
-Individuals observe $(X_{t},\eta_{t}(\cdot))$ and know the deterministic function $u_{t}(\cdot)$ for all $t\geq 1$.
-We consider the case where $X_{t}$ consists of persistent or serially correlated heterogeneities that are not reported in the data; conceptually, it is not hard to extend to the cases where some components in $X_{t}$are also reported in the sample.
-To allow the transition of states to depend on choice, it is useful to also define the potential state $X_{t}(d^{t-1})$ that would have occurred in period $t$ under choice history $d^{t-1}$; as usual, $X^{T}=X^{T}(D^{T-1})$.
-Summarizing, in each period $t$, the individual observes $(\eta_{t}(\cdot),X_{t})$ before choosing $D_{t}$, and observes $(H_{t},\xi_{t})$ after choosing $D_{t}$.
-The individual knows the full joint distribution of potential random variables across periods, but not the realizations except as previously specified.
-For example, the individual does not generally observe the realization of $\alpha(\cdot)$ or $\varepsilon_{t}(\cdot)$ in the outcome-generating process.
-Define $\zeta^{t}(d^{t-1})\equiv(\xi^{t-1}(d^{t-1}),X^{t}(d^{t-1}),\eta^{t}(\cdot))$ as the potential accumulation of the individual’s information in period $t$ given potential choice history $d^{t-1}$, which is not reported in the data.
-With minor loss of generality, we consider choices $D_{t}$ that are a function of the individual’s observed and unobserved information:
-$D_{t}=\phi_{t}(H^{t-1},\zeta^{t})$ (28)
-Unlike time-homogeneous, Markovian choice rules in a typical dynamic discrete choice model, e.g., Rust (1987), the (rational) rule of treatment choices in this section needs to be neither Markovian nor time-homogeneous.
-This is because the individual unit needs to update its beliefs in each period using the full history of outcomes and endogenous choices in the past.
-The key sufficient condition for sequential exchangeability is that the unobserved information for any counterfactual choice sequence is essentially uninformative about the outcome process, conditional on the initial observed outcome:
-$\zeta^{T}(d^{T-1})\perp(\alpha(\cdot),\varepsilon^{T}(d^{T}))\mid Y_{0}\quad\text{for all $d^{T}\in\mathcal{D}^{T}$.}$ (29)
-For clarity at the expense of some generality, this nests several more primitive assumptions based on the unobserved vector $\zeta^{t}=(\xi^{t-1},X^{t},\eta^{t}(\cdot))$.
-First, there is no meaningful unobserved heterogeneity in prior beliefs through $\xi_{0}$, though prior beliefs may vary systematically with $Y_{0}$.
-Second, there is no meaningful unobserved heterogeneity in incremental learning through $\xi_{t}$ for $t\geq 1$, though the belief parameter $\psi_{t}$ may evolve systematically based on the observed history $H^{t-1}$.
-Third, the state transition process $X^{T}$ is uninformative about the outcome process, so that there is no learning based on the unobserved evolution of states.
-Finally, independence of the shocks in instantaneous ex post utilities $\eta_{t}(\cdot)$ is standard.
-In summary, (29) permits systematic learning based only on observables.
-In that case, (mean) sequential exchangeability is satisfied.
-Proposition 6
-Suppose (29) holds in the structural choice model with learning. Then sequential exchangeability (Assumption 2) is satisfied.
-The intuition underlying this result is straightforward.
-By definition (27), for any sequence of potential treatment $d^{t}$, the history of potential outcomes $Y^{t}(d^{t})$ is a function of initial condition $Y_{0}$, fixed effects $\alpha(\cdot)$ and noises $\varepsilon^{t}(d^{t})$ only.
-Meanwhile, conditional on an observed history $D^{s-1}=d^{s=1}$, $Y^{s-1}(d^{s-1})=y^{s-1}$, any event $D_{s}=d_{s}$ is defined over the sample space of choice unobservables $\zeta^{s}(d^{s-1})$, which is conditionally independent from the outcome unobservables, and consequently the sequence of potential outcomes.
-This means the confounders for the effect of $D_{s}$ on subsequent potential outcomes $Y_{t}(d^{s},\cdot)$ are also controlled for in this case under sequential exchangeability.
-Our interpretation of this result is twofold.
-First, it highlights the limitations of sequential exchangeability in a dynamic choice context.
-In this sense, the result is more about specifying the channels of unobserved choice heterogeneity that must be ruled out, than about what is allowed.
-Thus, even as the choice model recovers the possibility of fixed effects in the outcome model without violating sequential exchangeability, it imposes significant restrictions on unobserved heterogeneity in choice.
-At the same time, viewed in a more positive light, Proposition 6 provides sufficient conditions under which sequential exchangeability can be used to identify the causal effects when treatment choices are made by forward-looking individuals capable of learning.
-In this case, sequential exchangeability provides an alternative suite of methods for recovering treatment effects in the presence of learning under (limited) unobserved heterogeneity, complementary to the pure learning model of Bunting, Diegert, and Maurel (2022).
-4.2 Comparison With Parallel Trends: Under Learning
-We next compare sequential exchangeability and parallel trends within the context of the structural learning model in Section 4.1.
-It is most illustrative to make such a comparison in a simplified version of the model with no additional states $X_{t}(d^{t-1})$ and $T=2$.
-To reiterate, in such a simplified case, the potential outcomes are:
-$Y_{1}(d_{1})$ $=$ $\lambda_{1}(d_{1},Y_{0},\alpha(d_{1}),\varepsilon_{1}(d_{1}))\text{;}$ (30)
-$Y_{2}(d_{1},d_{2})$ $=$ $\lambda_{2}(d_{1},d_{2},Y_{0},Y_{1}(d_{1}),\alpha(d_{2}),\varepsilon_{2}(d_{2}))\text{,}$
-and the rules of treatment choices are:
-$D_{1}$ $=$ $\phi_{1}(Y_{0},\xi_{0},\eta_{1}(\cdot))\text{;}$ (31)
-$D_{2}$ $=$ $\phi_{2}(Y_{0},Y_{1},D_{1},\xi^{1}(D_{1}),\eta_{2}(\cdot))\text{,}$
-where $\xi^{1}(d_{1})\equiv(\xi_{0},\xi_{1}(d_{1}))$. [^7] Recall $D_{0}=0$ w.p.1 and is therefore dropped from $\xi_{0}(D_{0}),\xi_{1}(D_{0},D_{1})$ for simplicity.
-In the current setting with dynamic treatment effects, the parallel trends assumption (conditional on initial outcomes $Y_{0}$) requires:
+So far, we have provided a causal interpretation of a first-differenced IV estimand under the sequential exchangeability (SE) condition in Section 3.2, and investigated two semiparametric structural models for which causal parameters, such as ATFT, are identified without SE in Section 3.3.
+Our work complements the popular difference-in-differences method, which estimates ATT under a parallel trends (PT) assumption.
+In the two-period case of our setting with heterogeneous intertemporal treatment effects, the PT assumption, conditional on initial outcomes $Y_{0}$, amounts to:
 $E[Y_{2}(0,0)-Y_{1}(0)|D_{1},D_{2},Y_{0}]\text{ is invariant in }(D_{1},D_{2}).$ (28)
 See, e.g., de Chaisemartin and D’Haultfoeuille (2024).
-The next proposition shows, under a conditional independence condition, sequential exchangeability holds whereas parallel trends would not hold in general.
-Proposition 7
-Consider the model Section 4.1 with no states $X_{t}(d^{t-1})$ and $T=2$. If
-$(\alpha(\cdot),\varepsilon^{2}(d^{2}))\perp(\xi^{1}(d_{1}),\eta^{2}(\cdot))\mid Y_{0}\text{ for all }d^{2}\equiv(d_{1},d_{2})\text{,}$ (33)
-then sequential exchangeability in Assumption 2b holds while $E[Y_{2}(0,0)-Y_{1}(0)|D_{1},D_{2},Y_{0}]$ is in general a non-degenerate function of $(D_{1},D_{2})$.
-The conditional independence in (33) specializes (29) when there are no additional states $X_{t}(d^{t-1})$ and $T=2$.
-This condition posits the unobserved states influencing posterior formation are independent from the individual fixed effects, after controlling for the initial outcome $Y_{0}$; it does not further restrict the learning process.
-For example, consider a simplistic case where $\varepsilon^{2}(d^{2}),\eta^{2}(\cdot)$ are idiosyncratic noises independent from all other variables, and where $\xi^{t}(d^{t})=\xi_{0}$.
-That is, the evolution of posterior beliefs only depends on an initial shifter of the prior $\xi_{0}$ and observed history $(D^{t-1},Y^{t-1})$, but no incremental, time-varying $\xi_{s}(d_{s})$ for $s\geq 1$.
-Then (33) holds when $\alpha(\cdot)$ is independent from the initial prior shifter $\xi_{0}$ given $Y_{0}$.
-The result in Proposition 7 is intuitive.
-With forward-looking and learning, the treatment choices $D^{2}=(D_{1},D_{2})$ are a non-trivial function of a realized history of observed outcomes $Y_{0},Y_{1}(D_{1})$.
-Hence conditioning on different realizations of $D^{2}$ amounts to conditioning on different events defined over the sample space of $(Y_{0},\alpha(\cdot),\varepsilon_{1}(\cdot),\xi^{1}(\cdot),\eta^{2}(\cdot))$.
-Because $Y_{2}(0,0)-Y_{1}(0)$ is a function of $(Y_{0},\alpha(\cdot),\varepsilon^{2}(\cdot))$, its distribution (and mean) generally vary with the conditioning event defined by $D^{2}$.
-This non-degeneracy holds in general, even after controlling for the realization in $Y_{0}$.
+In this section, we study which kinds of relations between treatment selection and potential outcomes are permitted or ruled out under SE and PT respectively.
+This provides a framework for empirical researchers to assess the circumstances under which SE and PT are justified in experimental or observational data.
+Since we no longer distinguish between homogeneous and heterogeneous individual-level parameters as in Subsection 3.3, we again suppress indexing by individual effects.
+In Section 4.1, we confirm that SE holds in an experimental design of sequentially randomized treatment (where the treatment in period $t$ is randomized conditioning on the history of past treatments and outcomes); in contrast (and perhaps surprisingly), PT would not generally hold in such a case.
+In Section 4.2, we turn to observational data and showcase the nature of SE restrictions within a structural model where treatments are chosen by rational, forward-looking individuals with learning; again we compare these SE restrictions with the PT conditions.
 4.1 Sequentially Randomized Treatments
-We conclude this section with yet another intuitive and important case of treatment assignment mechanism, in which sequential exchangeability holds but parallel trends do not.
-This is the case where the treatment assignments are sequentially randomized conditional on earlier history of treatments and outcomes.
-Suppose the treatments are sequentially randomized as follows:
+Suppose the treatments are sequentially randomized conditional on earlier history of realized treatments and outcomes. For $T=2$, this means:
 $D_{1}\perp(Y_{1}(\cdot),Y_{2}(\cdot,\cdot))$ $\mid$ $Y_{0},$ (29)
 $D_{2}\perp Y_{2}(D_{1},\cdot)$ $\mid$ $Y_{1},Y_{0},D_{1}\text{.}$ (30)
-Such a treatment assignment mechanism is relevant especially in experimental designs where a researcher may be able to assign randomized treatment in each period after controlling for observed outcomes up to that period.
-For simplicity, let the initial outcome $Y_{0}$ be degenerate at a constant value and suppress it in all conditioning events below.
-Mean sequential exchangeability in Assumption 2b holds under (34) and (35), because
+Such a treatment assignment mechanism is relevant especially in experimental designs where a researcher may be able to assign randomized treatments in each period after controlling for observed outcomes up to that period.
+We show that while the sequential randomization of treatments implies sequential exchangeability, it is generally not sufficient for parallel trends.
+For simplicity, let the initial condition $Y_{0}$ be degenerate (constant), and suppress it in all conditioning events.
+Proposition 7
+Sequential randomization of treatments in (29) and (30) implies the (mean) sequential exchangeability in Assumption 2b, but not the parallel trends in (28).
+The proof of Proposition 7 is intuitive and included in the text. Mean sequential exchangeability in Assumption 2b holds under (29) and (30) because:
 $E[Y_{2}(d_{1},d_{2})|D_{2},Y_{1},D_{1}=d_{1}]$ $=$ $E[Y_{2}(d_{1},d_{2})|Y_{1},D_{1}=d_{1}]\text{,}$
@@
 for all $d_{1},d_{2}\in\{0,1\}$.
-In contrast, the parallel trends condition in (32) would fail in general under (34) and (35). To see this, note:
+In contrast, the parallel trends condition in (28) is not implied under (29) and (30). To see this, note:
 $E[Y_{2}(0,0)-Y_{1}(0)|D_{2}=d_{2},D_{1}=0]$
-[…]
-[…]
-where $\mu_{2}(s)\equiv E[Y_{2}(0,0)|Y_{1}(0)=s,D_{1}=0]$ is a function of $s$ that does not vary with $d_{2}$, and the second equality is due to (35). Besides, by construction,
-[…]
-[…]
-where $p(s)\equiv\Pr\{D_{2}=1|Y_{1}(0)=s,D_{1}=0\}$ and $G(s)\equiv\Pr\{Y_{1}(0)\leq s|D_{1}=0\}=\Pr\{Y_{1}(0)\leq s\}$.
+$=$ $\int E[Y_{2}(0,0)-Y_{1}(0)|D_{2}=d_{2},Y_{1}(0)=y_{1},D_{1}=0]dF_{Y_{1}(0)}(y_{1}|D_{2}=d_{2},D_{1}=0)$
+$=$ $\int\left[\mu_{2}(y_{1})-y_{1}\right]dF_{Y_{1}(0)}(y_{1}|D_{2}=d_{2},D_{1}=0)\text{,}$
+where $\mu_{2}(d_{2},s)\equiv E[Y_{2}(0,0)|D_{2}=d_{2},Y_{1}(0)=s,D_{1}=0]$ does not vary with $d_{2}$ under (30), and hence the first argument is suppressed. Besides, by construction,
+$F_{Y_{1}(0)}(y|D_{2}=1,D_{1}=0)$ $=$ $\frac{\int_{-\infty}^{y}p(s)dG(s)}{\int_{-\infty}^{+\infty}p(s)dG(s)}\text{,}$
+$F_{Y_{1}(0)}(y|D_{2}=0,D_{1}=0)$ $=$ $\frac{\int_{-\infty}^{y}[1-p(s)]dG(s)}{\int_{-\infty}^{+\infty}[1-p(s)]dG(s)}\text{,}$
+where $p(s)\equiv\Pr\{D_{2}=1|Y_{1}(0)=s,D_{1}=0\}$, and $G(s)\equiv\Pr\{Y_{1}(0)\leq s|D_{1}=0\}=\Pr\{Y_{1}(0)\leq s\}$ under (29).
 Thus, when $p(\cdot)$ is non-degenerate and non-linear, $F_{Y_{1}(0)}(y|D_{2}=d_{2},D_{1}=0)$ generally varies with $d_{2}$.
-As a result, (32) would not hold in general.
-Note that (32) holds in a special case where $p(\cdot)$ is degenerate, e.g., $D_{2}$ is randomly assigned independent of past $Y_{1}(\cdot)$.
+Therefore, the parallel trends condition in (28) does not hold in general.
+Note that (28) does hold in a special case where $p(\cdot)$ is degenerate, e.g., $D_{2}$ is randomly assigned independent of past $Y_{1}(\cdot)$ conditional on $D_{1}=0$.
 In such a case, $F_{Y_{1}(0)}(\cdot|D_{2}=d_{2},D_{1}=0)=G(\cdot)$ does not vary with $d_{2}$ and (28) holds.
+This is consistent with results in Marx, Tamer, and Tang (2024) on failure of parallel trends in models where treatments are functions of past outcomes.
+4.2 Forward-Looking Treatment Choices with Learning
+We now study when the SE and PT conditions are justified in an observational setting, where individuals self-select into treatments through rational, forward-looking choices based on beliefs updated through learning.
+Consider a structural model with $T=2$, where the potential outcomes are:
+$Y_{1}(d_{1})$ $=$ $\lambda_{1}(d_{1},Y_{0},\alpha(d_{1}),\varepsilon_{1})\text{,}$ (31)
+$Y_{2}(d_{1},d_{2})$ $=$ $\lambda_{2}(d_{1},d_{2},Y_{0},Y_{1}(d_{1}),\alpha(d_{2}),\varepsilon_{2})\text{,}$
+with $\lambda_{t}(\cdot)$ being deterministic functions, $\alpha(\cdot)$ potential individual fixed effects that vary with contemporary counterfactual treatments $d_{t}$, and $\varepsilon_{t}$ idiosyncratic, time-varying structural errors in potential outcomes.
+The DGP for realized treatment choices are:
+$D_{1}$ $=$ $\phi_{1}(Y_{0},\xi_{0},\eta_{1}),$ (32)
+$D_{2}$ $=$ $\phi_{2}(Y^{1},D_{1},\xi_{1}(D_{1}),\eta_{2}),$
+where $Y^{1}\equiv(Y_{0},Y_{1})\equiv(Y_{0},Y_{1}(D_{1}))$, $\phi_{t}(\cdot)$ are deterministic functions, and $\xi_{0}$ and $\xi_{1}(d_{1})$ are potential individual fixed effects which may evolve over time, with $\xi_{1}(d_{1})$ possibly subsuming $\xi_{0}$ as a subvector, and $\eta_{t}$ are idiosyncratic, time-varying noises.
+[^10] Recall that $D_{0}=0$ w.p.1, and is dropped from $\xi_{0}(D_{0}),\xi_{1}(D_{0},D_{1})$ for simplicity.
+Our results below generalize to the cases with $T\geq 3$, or where the idiosyncratic noises $\varepsilon_{t}$ and $\eta_{t}$ are also indexed by contemporary counterfactual treatments $d_{t}$.
+Such generalizations do not require qualitatively different argument, and are omitted for conciseness.
+We provide an example of a structural model where the treatment rule in (32) arises endogenously from the rational choices by forward-looking individuals with learning.
+Example 1. (Dynamic treatment choices with learning.) Consider a model with two periods $T=2$. Each individual’s per-period, ex post payoff from choosing $D_{1}=d_{1}$ at $t=1$ is:
+$u_{1}(d_{1},Y_{1}(d_{1}))+\eta_{1}(d_{1}).$
+Moreover, given $D_{1}$, the instantaneous, ex post payoff from choosing $D_{2}=d_{2}$ at $t=2$ is:
+$u_{2}(d_{2},Y_{2}(D_{1},d_{2}))+\eta_{2}(d_{2}).$
+Potential outcomes $Y_{t}(\cdot)$ are determined by $(Y_{0},\varepsilon_{1},\varepsilon_{2})$ and individual fixed effects $\alpha\equiv(\alpha(0),\alpha(1))$ via deterministic functions $\lambda_{t}(\cdot)$ as defined in (31).
+Individuals know all functions $u_{t}(\cdot)$ and $\lambda_{t}(\cdot)$, but never observe the potential fixed effects $\alpha$ and idiosyncratic errors $\varepsilon_{t}$ in $Y_{t}(\cdot)$.
+In each period $t$, individuals observe $\eta_{1}\equiv(\eta_{1}(0),\eta_{1}(1))$, and know the distribution of the idiosyncratic errors $\varepsilon_{t}$.
+At the start of the first period $t=1$, each individual holds a prior belief for $\alpha$.
+This prior is characterized by a parameter vector $\kappa_{1}\equiv\varphi_{1}(Y_{0},\xi^{*})$, where $\xi^{*}$ is a persistent fixed effect known privately to the individual.
+Similarly, at the start of the next period $t=2$, the individual uses past observed outcomes $Y^{1}$ and the realized treatment $D_{1}$ to update the posterior for $\alpha$, which is characterized by $\kappa_{2}\equiv\varphi_{2}(Y^{1},D_{1},\xi^{*})$.
+In period $t=2$, a rational individual with learning chooses:
+$D_{2}=\operatorname*{argmax}_{d_{2}\in\{0,1\}}\underset{\mu_{2}(d_{2},Y^{1},D_{1},\xi^{*})}{\underbrace{\int u_{2}(d_{2},y_{2})dF_{Y_{2}(D_{1},d_{2})|Y^{1},D_{1},\xi^{*}}(y_{2})}}+\eta_{2}(d_{2}),$ (33)
+where $F_{Y_{2}(D_{1},d_{2})|Y^{1},D_{1},\xi^{*}}$ denotes the posterior for $Y_{2}(D_{1},d_{2})$ according to $\kappa_{2}$.111111 Formally, $F_{Y_{2}(D_{1},d)|Y^{1},D_{1},\xi^{*}}(y)=\int\left[\int 1\{\lambda_{2}(D_{1},d,Y^{1},a,e)\leq y\}dF_{\varepsilon_{2}}(e)\right]dF_{\alpha(d)|\kappa_{2}}(a)$, where $F_{\alpha(d)|\kappa_{2}}$ denotes the posterior for $\alpha(d)$ at the start of period $t=2$ given $(Y^{1},D_{1},\xi^{*})$.
+In period $t=1$, a rational, forward-looking individual chooses:
+$D_{1}=\operatorname*{argmax}_{d_{1}\in\{0,1\}}\int\left[u_{1}(d_{1},y_{1})+\rho V_{2}(Y_{0},y_{1},d_{1},\xi^{*})\right]dF_{Y_{1}(d_{1})|Y_{0},\xi^{*}}(y_{1})+\eta_{1}(d_{1}),$ (34)
+where $\rho$ is the time discount factor, $F_{Y_{1}(d)|Y_{0},\xi^{*}}$ is the prior of $Y_{1}(d)$ implied by $\kappa_{1}$, and:
+$V_{2}(Y_{0},Y_{1},D_{1},\xi^{*})\equiv\int\max_{d_{2}\in\{0,1\}}[\mu_{2}(d_{2},Y_{0},Y_{1},D_{1},\xi^{*})+\tilde{\eta}_{2}(d_{2})]F_{\eta_{2}}(\tilde{\eta}_{2}).$
+The solutions of (33) and (34) imply that rational treatment choices by forward-looking individuals with learning necessarily takes the form of (32) with persistent individual fixed effect in the prior and posterior: $\xi_{0}=\xi_{1}(\cdot)=\xi^{*}$.
+(End of Example 1.)
+The next proposition shows, in an observational context where potential outcomes and realized treatments are determined as in (31) and (32), a conditional independence condition is sufficient for sequential exchangeability, but does not imply parallel trends in general.
+Proposition 8
+Suppose the potential outcomes and realized treatments are determined as in (31) and (32), and:
+$(\alpha,\varepsilon_{1},\varepsilon_{2})\perp(\xi_{0},\xi_{1}(d_{1}),\eta_{1},\eta_{2})\mid Y_{0}\text{ for all }d_{1}\text{.}$ (35)
+Then sequential exchangeability in Assumption 2b holds while $E[Y_{2}(0,0)-Y_{1}(0)|D_{1},D_{2},Y_{0}]$ is in general non-degenerate in $(D_{1},D_{2})$.
+The conditional independence in (35) posits that the fixed effects and idiosyncratic errors influencing posteriors are independent from those determining the potential outcomes, after controlling for the initial outcome $Y_{0}$.
+It does not further restrict learning.
+For example, consider a simplistic case where $\varepsilon^{2},\eta^{2}$ are idiosyncratic noises independent from all other variables, and where $\xi_{1}(d_{1})=\xi_{0}$ almost surely.
+That is, the evolution of the posterior depends on an initial shifter of the prior $\xi_{0}$ and observed history $(D^{t-1},Y^{t-1})$.
+Then (35) holds when $\alpha(\cdot)$ is independent from $\xi_{0}$ given $Y_{0}$.
+The general failure of the PT condition in Proposition 8 is intuitive.
+With forward-looking and learning, the treatment choices $D^{2}=(D_{1},D_{2})$ are a non-trivial function of a realized history of observed outcomes $Y_{0},Y_{1}(D_{1})$.
+Hence conditioning on different realizations of $D^{2}$ amounts to conditioning on different events defined over the sample space of $(Y_{0},\alpha(\cdot),\varepsilon_{1},\xi_{0},\xi_{1}(\cdot),\eta^{2})$.
+Because $Y_{2}(0,0)-Y_{1}(0)$ is a function of $(Y_{0},\alpha(\cdot),\varepsilon^{2})$, its distribution (and mean) generally vary with the conditioning event defined by $D^{2}$.
+This non-degeneracy holds in general, even after controlling for the realization in $Y_{0}$.
 5 Conclusion
-In this note we considered the causal interpretation of dynamic panel data estimators in models of potential outcomes that can depend on past treatments.
-We derived sufficient conditions for instrumented first-difference estimators to be causally interpretable and proposed a new adjusted inverse probability of treatment estimator that uses the first-differenced outcome to recover a history-dependent average treatment effect.
-An underlying assumption for (most of) these results was sequential exchangeability.
-In addition to investigating the plausibility of this assumption in a structural model of dynamic choice, we showed that this assumption is not necessary per se for dynamic panel data estimators to recover causal estimands in our setting.
+We studied the causal interpretation of dynamic panel data estimators in models where potential outcomes depend on past treatments.
+We derived sufficient conditions for instrumented first-difference estimators to be causally interpretable.
+A motivation and underlying assumption for one set of these results was sequential exchangeability.
+Besides investigating the plausibility of this assumption in both experimental and observational settings (with a structural model of dynamic treatment choices), we showed how to conduct causal inference using dynamic panel data estimators when this assumption fails.
 A natural avenue for future work is to consider (partial) identification when sequential exchangeability is relaxed or holds only conditional on a set of unobservables.
-Appendix A: Proofs
-Proof of Lemma 1.
-Let $\mathcal{Y}$ denote the (discrete) support of $Y_{0}$.
-Let $Z_{i}\equiv\{\mathbb{I}[(D_{i1},Y_{i0})=(d_{1},y_{0})]\}_{d_{1}\in\{0,1\},y_{0}\in\mathcal{Y}}$ be a row-vector of $2\times\#\mathcal{Y}$ (dummy) instrumental variables.
-The 2SLS estimator for the coefficient of $\Delta D_{i2}$ in (8) is constructed as follows.
-In the (fully saturated) first stage, obtain the fitted values of $\Delta Y_{i1}$ (and $\Delta D_{i2}$) from their regression on $Z_{i}$.
-These fitted values are $Z_{i}\hat{A}$ (and $Z_{i}\hat{B}$, respectively), where
-$\hat{A}\equiv\left(\frac{1}{n}\sum\nolimits_{i}Z_{i}^{\prime}Z_{i}\right)^{-1}\left[\frac{1}{n}\sum\nolimits_{i}Z_{i}^{\prime}\left(\Delta Y_{i1}\right)\right]\text{ and }\hat{B}\equiv\left(\frac{1}{n}\sum\nolimits_{i}Z_{i}^{\prime}Z_{i}\right)^{-1}\left[\frac{1}{n}\sum\nolimits_{i}Z_{i}^{\prime}\left(\Delta D_{i2}\right)\right]\text{.}$
-By the property of ordinary least squares estimators,
-[…]
-[…]
-Formally, $A$ (and $B$) are $(2\times\#\mathcal{Y)}$-by-$1$ column vectors with each component being the conditional mean of $\Delta Y_{i1}$ (and $\Delta D_{i2}$) given $(D_{i1},Y_{i0})=(d_{1},y_{0})$ for some $d_{1}\in\{0,1\}$, $y_{0}\in\mathcal{Y}$.
-This means $Z_{i}A=E[\Delta Y_{i1}|D_{i1},Y_{i0}]$ and $Z_{i}B=E[\Delta D_{i2}|D_{i1},Y_{i0}]$ by construction.
-In the second stage, regress $\Delta Y_{i2}$ on an intercept and the first-stage fitted values $Z_{i}\hat{A}$, $Z_{i}\hat{B}$.
-Our goal is to investigate the causal interpretation of the slope estimate for $Z_{i}\hat{B}$ in this second-stage regression.
-By the Frisch-Waugh-Lovell Theorem, this slope estimate is numerically equivalent to
-$\frac{\frac{1}{n}\sum_{i}\hat{W}_{i}\Delta Y_{i2}}{\frac{1}{n}\sum_{i}\hat{W}_{i}^{2}}\text{,}$ (36)
-where $\hat{W}_{i}$ is the residual from regressing $Z_{i}\hat{B}$ on $[1,Z_{i}\hat{A}]\equiv\hat{M}_{i}$. That is,
-$\hat{W}_{i}$ $=$ $Z_{i}\hat{B}-\hat{M}_{i}\underset{\equiv\hat{C}}{\underbrace{\left(\frac{1}{n}\sum\nolimits_{i}\hat{M}_{i}^{\prime}\hat{M}_{i}\right)^{-1}\left(\frac{1}{n}\sum\nolimits_{i}\hat{M}_{i}^{\prime}Z_{i}\right)}}\hat{B}$ (37)
-$=$ $\left(Z_{i}-\hat{M}_{i}\hat{C}\right)\hat{B}\text{,}$
-where
-[…]
-[…]
-We now turn to the denominator of the 2SLS estimator of slope coefficient of $\Delta D_{2}$ in (36):
-$\frac{1}{n}\sum\nolimits_{i}\hat{W}_{i}^{2}=$ $\hat{B}^{\prime}\left[\frac{1}{n}\sum\nolimits_{i}(Z_{i}-\hat{M}_{i}\hat{C})^{\prime}(Z_{i}-\hat{M}_{i}\hat{C})\right]\hat{B}$
-[…]
-$=$ $E\left[(Z_{i}B-M_{i}CB)^{\prime}(Z_{i}B-M_{i}CB)\right],$
-[…]
-Next, note the numerator is
-$\frac{1}{n}\sum\nolimits_{i}\hat{W}_{i}\left(\Delta Y_{i2}\right)=$ $\hat{B}^{\prime}\left[\frac{1}{n}\sum\nolimits_{i}(Z_{i}-\hat{M}_{i}\hat{C})^{\prime}(\Delta Y_{i2})\right]$
-[…]
-$=$ $E\left[(Z_{i}B-M_{i}CB)^{\prime}(\Delta Y_{i2})\right]=E[w(Y_{i0},D_{i1})\left(\Delta Y_{i2}\right)]\text{,}$
-[…]
-Proof of Proposition 1. Throughout the following proof, we fix $s\geq 1$. For $t=s$, the equivalence is immediate:
+Appendix A Proofs
+Proof of Proposition 1. We fix $s\geq 1$ throughout the following proof. For $t=s$, the equivalence is immediate:
 $E[Y_{s}(d^{s})|Y^{s-1},D^{s-1}=d^{s-1},D_{s}=d_{s}^{\prime}]\quad\text{is invariant in $d_{s}^{\prime}$}$
@@
 $\iff$ $E[Y_{s}(d^{s})-Y_{s-1}(d^{s-1})|Y^{s-1}(d^{s-1}),D^{s-1}=d^{s-1},D_{s}=d_{s}^{\prime}]\quad\text{is invariant in $d_{s}^{\prime}$.}$
-For $t>s$, Assumption 2 implies:
+For $t>s$, Assumption 2b implies:
 $E[Y_{t}(d^{t})|Y^{s-1},D^{s-1}=d^{s-1},D_{s}=d_{s}^{\prime}]\quad\text{is invariant in $d_{s}^{\prime}$}$
@@
 $E[Y_{t+1}(d^{t+1})|Y^{s-1},D^{s-1}=d^{s-1},D_{s}=d_{s}^{\prime}]\quad\text{is invariant in $d_{s}^{\prime}$.}$
-Finally, recall we have already shown the equivalence (between the two sets of conditions on the levels and the trends of potential outcomes) for $t=s$ at the beginning of this proof.
-Therefore, the claim of the reverse direction of implication for $t>s$ holds.
+Finally, recall we already showed the equivalence (between the two sets of conditions on the levels and the trends of potential outcomes) for $t=s$ at the start of this proof.
+Hence by induction the claim of the reverse direction of implication for $t>s$ holds.
+[…]
+$E[Z^{\prime}\Delta X]^{-1}E[Z^{\prime}\Delta Y_{2}]$
+Note that $Z$ contains conditional expectations of each element of $\Delta X$. By the law of iterated expectations, this estimand is equal to an OLS estimand:
+$E[Z^{\prime}Z]^{-1}E[Z^{\prime}\Delta Y_{2}].$
+By the formula for partitioned regressions, a.k.a. a population version of the Frisch-Waugh-Lovell Theorem, the estimand for the slope coefficient of the first component in $Z$ above is equal to
+$\frac{E[w(Y_{0},D_{1})\Delta Y_{2}]}{E[w(Y_{0},D_{1})^{2}]},$
+where $w(Y_{0},D_{1})$, defined in (9), is the projection error from the linear projection of $E(\Delta D_{2}|Y_{0},D_{1})$ onto $E(\Delta Y_{1}|Y_{0},D_{1})$ and a constant intercept.
+Plugging in the decomposition (8) of $\Delta Y_{2}$ yields the desired expression in (10).
 For proofs of Propositions 2 and 3, it is useful to expand the definition of (mean) sequential exchangeability for $T=2$ to full sequential exchangeability, which is:
@@
 for all $d_{1},d_{2}\in\{0,1\}$.
-Proof of Proposition 2. By Lemma 1, the estimator for the coefficient $\beta$ converges in probability to (13), recalled here for convenience:
-$E\left[\frac{w(Y_{0},D_{1})D_{2}}{E[w(Y_{0},D_{1})^{2}]}\times\tau_{2}(D_{1})\right]+E\left[\frac{w(Y_{0},D_{1})}{E[w(Y_{0},D_{1})^{2}]}\times\delta_{2}(D_{1})\right]$
-where $w(Y_{0},D_{1})$ is the projection error in the linear projection of $E(\Delta D_{2}|Y_{0},D_{1})$ on $E(\Delta Y_{1}|Y_{0},D_{1})$.
-The first step of the proof is to derive a closed-form expression for the projection errors $w(Y_{0},D_{1})$. We have:
+Proof of Proposition 2. By Lemma 1, the (first-differenced) IV estimand $\widetilde{\beta}$ is:
+$E\left[\frac{w(Y_{0},D_{1})D_{2}}{E[w(Y_{0},D_{1})^{2}]}\times\tau_{2}(D_{1})\right]+E\left[\frac{w(Y_{0},D_{1})}{E[w(Y_{0},D_{1})^{2}]}\times\delta_{2}(D_{1})\right],$ (41)
+where $w(Y_{0},D_{1})$ is the error in the linear projection of $E(\Delta D_{2}|Y_{0},D_{1})$ on $E(\Delta Y_{1}|Y_{0},D_{1})$.
+First, we derive a closed-form expression for the projection errors $w(Y_{0},D_{1})$. We have:
 $E[\Delta Y_{1}|Y_{0},D_{1}]$ $=E[\delta_{1}+D_{1}\tau_{1}|Y_{0},D_{1}]$
@@
 where the first equality is due to the decomposition in (7), the second equality rearranges, the third equality follows from mean sequential exchangeability (Assumption 2b), specifically (38), and the fourth equality follows from limited heterogeneity (Assumption 3), specifically (11) and (13).
-Furthermore, by (16), this expression is strictly monotone in the realization $D_{1}$.
+Furthermore, by (13), this right-hand side above is strictly monotone in $D_{1}$.
 Therefore, the linear projection of $E(\Delta D_{2}|Y_{0},D_{1})$ on $E(\Delta Y_{1}|D_{1},Y_{0})$ is identical to its mean conditional on $D_{1}$.
@@
 $=E(D_{2}|Y_{0},D_{1})-E(D_{2}|D_{1}).$ (42)
-The second step of the proof is to eliminate the second bias term in (13) by establishing that its numerator:
+Next, we eliminate the second (bias) term in (41) by showing its numerator:
 $E[w(Y_{0},D_{1})\delta_{2}(D_{1})]$
-equals zero. Applying the law of iterated expectations (LIE) over partitions $(Y_{0},D_{1})$, this numerator equals:
-$E[w(Y_{0},D_{1})E[\delta_{2}(D_{1})|Y_{0},D_{1}]]$ (44)
-Focusing on the inner expectation for treatment realization $d_{1}$,
+is zero. Applying the law of iterated expectations (LIE) over partitions $(Y_{0},D_{1})$, we write this numerator as:
+$E[w(Y_{0},D_{1})E[\delta_{2}(D_{1})|Y_{0},D_{1}]].$ (43)
+The inner expectation conditional on a realized treatment $D_{1}=d_{1}$ equals:
 $E[\delta_{2}(D_{1})|Y_{0},D_{1}=d_{1}]=E[\delta_{2}(d_{1})|Y_{0},D_{1}=d_{1}]=E[\delta_{2}(d_{1})|Y_{0}]=E[\delta_{2}(d_{1})]$ (44)
@@
 $=E[E[\delta_{2}(d_{1})]|D_{1}=d_{1}]$
-$=E[\delta_{2}(d_{1})]$
+$=E[\delta_{2}(d_{1})],$
 where the first equality follows from the LIE, the second from the previous equality (44), and the third from taking an expectation over a constant. Combining with (44), it follows that:
@@
 where the first equality follows from (45), the second by the LIE over partitions $D_{1}$, and the third because $E[w(Y_{0},D_{1})|D_{1}]=0$ by definition of $w(Y_{0},D_{1})$.
-The third step of the proof is to establish the convex weights over heterogeneous treatment effects in the first term of (13), with numerator $E[w(Y_{0},D_{1})D_{2}\tau_{2}(D_{1})]$.
+The last step of the proof is to establish the convex weights over heterogeneous treatment effects in the first term of (41), with numerator $E[w(Y_{0},D_{1})D_{2}\tau_{2}(D_{1})]$.
 Applying the LIE over partitions $(Y_{0},Y_{1},D_{1},D_{2})$ and invoking mean sequential exchangeability (Assumption 2b), specifically (40), yields:
@@
 where the first equality follows by definition of the observed variables at realized treatment, the second by the full sequential exchangeability (Assumption 2a) for period 1 in (36), specifically its implication that $D_{1}\perp Y_{2}(\cdot)|Y_{0},Y_{1}(d_{1})$ for any $d_{1}$, and the third by limited heterogeneity (Assumption 3), specifically (14).
-Similarly to the previous step, we also have:
+Similarly to the previous step, we have:
 $E[\tau_{2}(D_{1})|D_{1}=d_{1}]$ $=E[E[\tau_{2}(D_{1})|Y_{0},Y_{1},D_{1}=d_{1}]|D_{1}=d_{1}]$
 $=E[E[\tau_{2}(d_{1})]|D_{1}=d_{1}]$
-$=E[\tau_{2}(d_{1})]$
-where the first equality follows from the LIE, the second from the previous equality (48), and the third from taking an expectation over a constant. Combining with (48), it follows that:
+$=E[\tau_{2}(d_{1})],$
+where the first equality follows from the LIE, the second from (47), and the third from taking an expectation over a constant. Combining with (47), it follows that:
 $E[\tau_{2}(D_{1})|Y_{0},Y_{1},D_{1}]=E[\tau_{2}(D_{1})|D_{1}].$ (48)
-Proceeding with expression (47) for the first numerator:
+Thus, the right-hand side of (46) is:
 $E[w(Y_{0},D_{1})D_{2}E[\tau_{2}(D_{1})|Y_{0},Y_{1},D_{1}]]$ $=E[w(Y_{0},D_{1})D_{2}E[\tau_{2}(D_{1})|D_{1}]]$
@@
 $=E[w(Y_{0},D_{1})^{2}\tau_{2}(D_{1})]$
-where the first equality follows from (49), the second by the LIE over partition $(D_{1},Y_{0})$, the third by the LIE and definition of $w(Y_{0},D_{1})$ (which imply $E[w(Y_{0},D_{1})g(D_{1})]=0$ for any function $g(\cdot)$ with a sole argument $D_{1}$), the fourth by the previously derived projection errors (43), the fifth again by (49), the sixth by rearrangement, and the seventh again by the LIE.
+where the first equality follows from (48), the second by the LIE over partition $(D_{1},Y_{0})$, the third by the LIE and definition of $w(Y_{0},D_{1})$ (which implies $E[w(Y_{0},D_{1})g(D_{1})]=0$ for any function $g(\cdot)$ with a sole argument $D_{1}$), the fourth by the previously derived projection errors (42), the fifth again by (48), the sixth by rearrangement, and the seventh again by the LIE.
 Proof of Proposition 3. Plugging in the decomposition of $\Delta Y_{2}$ in (8), we get:
-$E[\widetilde{\Delta Y_{2}}]=E\left[\frac{D_{2}\tau_{2}(D_{1})+\delta_{2}(D_{1})-E[\Delta Y_{2}|Y^{1},D_{1},D_{2}=0]}{E[D_{2}|Y^{1},D_{1}]}\right]$ (50)
+$E[\widetilde{\Delta Y_{2}}]=E\left[\frac{D_{2}\tau_{2}(D_{1})+\delta_{2}(D_{1})-E[\Delta Y_{2}|Y^{1},D_{1},D_{2}=0]}{E[D_{2}|Y^{1},D_{1}]}\right].$ (49)
 Next, we show the second and third terms in the numerator cancel. Note:
@@
 $=E[\delta_{2}(D_{1})|Y^{1},D_{1}],$
-where the first equality follows by definition of $\Delta Y_{2}$, the second by plugging in observed variables and rearrangement, the third by mean sequential exchangeability (Assumption 2), specifically (42), and the fourth by rearrangement and the definition (9).
+where the first equality follows by definition of $\Delta Y_{2}$, the second by plugging in observed variables and rearrangement, the third by mean sequential exchangeability (Assumption 2b), specifically (40), and the fourth by rearrangement and the definition of $\delta_{2}$ in (6).
 Iterating expectations over the previous equality,
@@
 $=E[\tau_{2}(D_{1})],$
-where the first equality follows by the law of iterated expectations (LIE) over $(Y^{1},D_{1})$, the second by (51), the third again by the LIE over $(Y^{1},D_{1},D_{2})$, the fourth by mean sequential exchangeability (Assumption 2), specifically (42), the fifth by rearrangement and cancellation of terms $E[D_{2}|Y^{1},D_{1}]$, and the last again by the LIE.
+where the first equality follows by the law of iterated expectations (LIE) over $(Y^{1},D_{1})$, the second by (50), the third again by the LIE over $(Y^{1},D_{1},D_{2})$, the fourth by mean sequential exchangeability (Assumption 2b), specifically (40), the fifth by rearrangement and cancellation of terms $E[D_{2}|Y^{1},D_{1}]$, and the last again by the LIE.
 (Note mean sequential exchangeability is invoked only for $D_{2}$ in (40).)
-Proof of Proposition 4. For all $s\geq 1$,
-$E(\varepsilon_{is}\alpha_{i})$ $=$ $E[\alpha_{i}E\left(\varepsilon_{is}|Y_{i0},\alpha_{i}\right)]$
-$=$ $E\left\{\alpha_{i}\left[\sum\nolimits_{d^{s}}E\left(\varepsilon_{is}|D_{i}^{s}=d^{s},Y_{i0},\alpha_{i}\right)\times\Pr\{D_{i}^{s}=d^{s}|Y_{i0},\alpha_{i}\}\right]\right\}$
-$=$ $E\left\{\alpha_{i}\left[\sum\nolimits_{d^{s}}E\left(\varepsilon_{is}^{\ast}(d_{s})|D_{i}^{s}=d^{s},Y_{i0},\alpha_{i}\right)\times\Pr\{D_{i}^{s}=d^{s}|Y_{i0},\alpha_{i}\}\right]\right\}=0\text{,}$
-where the first and second equalities are due to the law of iterated expectation, the third due to the definition $\varepsilon_{it}=\varepsilon_{it}^{\ast}(D_{it})$, and the last due to the sequential exogeneity of $\varepsilon_{it}^{\ast}(\cdot)$ in (25).
-Hence $E(\Delta\varepsilon_{it}\alpha_{i})=0$ for all $t\geq 2$.
-Likewise, swapping the roles of $\alpha_{i}$ and $Y_{i0}$ and applying the same argument above shows that $E(\Delta\varepsilon_{it}Y_{i0})=0$.
-Furthermore, note that for $s\leq t$, the law of iterated expectation implies
-$E(D_{is}\varepsilon_{it})=E[D_{is}E(\varepsilon_{it}|D_{i}^{t},Y_{i0},\alpha_{i})]\text{,}$
-where
-$E\left[\varepsilon_{it}|D_{i}^{t}=d^{t}\equiv(d^{t-1},d_{t}),Y_{i0},\alpha_{i}\right]$
-$=$ $E\left[\varepsilon_{it}^{\ast}(d_{t})|D_{i}^{t}=(d^{t-1},d_{t}),Y_{i0},\alpha_{i}\right]=0$
-for all $d^{t}$, where the second equality holds due to (25). Thus (25) implies (28).
-Next, without loss of generality, consider $s<t$ so that
-$E(\varepsilon_{is}\varepsilon_{it})$ $=$ $E[E(\varepsilon_{is}\varepsilon_{it}|\varepsilon_{is},D_{i}^{s},Y_{i0},\alpha_{i})]$
-$=$ $E\left\{\varepsilon_{is}^{\ast}(D_{is})E[\varepsilon_{it}|\varepsilon_{is}^{\ast}(D_{is}),D_{i}^{s},Y_{i0},\alpha_{i}]\right\}\text{.}$
-By the law of iterated expectations,
-$E\left[\varepsilon_{it}|\varepsilon_{is}^{\ast}(D_{is}),D_{i}^{s}=d^{s},Y_{i0},\alpha_{i}\right]$
-$=$ $\sum_{d^{t}\in\omega_{t}(d^{s})}E\left[\varepsilon_{it}^{\ast}(d_{t})|D_{i}^{t}=d^{t},\varepsilon_{is}^{\ast}(d_{s}),Y_{i0},\alpha_{i}\right]\Pr\{D_{i}^{t}=d^{t}|\varepsilon_{is}^{\ast}(d_{s}),D_{i}^{s}=d^{s},Y_{i0},\alpha_{i}\}$
-$=$ $0\text{,}$
-where $\omega_{t}(d^{s})$ denotes the support of $D_{i}^{t}$ conditional on $D_{i}^{s}=d^{s}$, and the last equality holds because $E\left[\varepsilon_{it}^{\ast}(d_{t})|D_{i}^{t}=d^{t},\varepsilon_{is}^{\ast}(d_{s}),Y_{i0},\alpha_{i}\right]$ is zero under (25).
-In what follows, it will be useful to collect the following relationships between potential outcomes and composite errors $\nu_{it}^{*}(d_{t})\equiv\alpha_{i}^{*}+\varepsilon_{it}^{*}(d_{t})$ in the linear dynamic model of potential outcomes (21).
+In what follows, it will be useful to collect the following relationships between potential outcomes and composite errors $\nu_{it}^{*}(d^{t})\equiv\alpha_{i}^{*}+\varepsilon_{it}^{*}(d^{t})$ in the linear dynamic model of potential outcomes (17).
 Lemma 2
-For any $d^{t}\in\{0,1\}^{t}$ and $t\geq s\geq 1$,
-$\nu_{it}^{*}(d_{t})$ $=Y_{it}(d^{t})-\gamma^{*}Y_{it-1}(d^{t-1})-\beta^{*}d_{t}-\theta_{t}^{*}$ (52)
-$Y_{it}(d^{t})$ $=(\gamma^{*})^{t-s+1}Y_{is-1}(d^{s-1})+\sum_{r=s}^{t}(\gamma^{*})^{t-r}\times[\beta^{*}d_{r}+\theta_{r}^{*}+\nu_{ir}^{*}(d_{r})]$ (53)
+Suppose the potential outcomes follow an AR(1) model in (17). For any $d^{t}\in\{0,1\}^{t}$ and $t\geq s\geq 1$,
+$\nu_{it}^{*}(d^{t})$ $=Y_{it}(d^{t})-\gamma^{*}Y_{it-1}(d^{t-1})-\beta^{*}d_{t}-\theta_{t}^{*},$ (51)
+$Y_{it}(d^{t})$ $=(\gamma^{*})^{t-s+1}Y_{is-1}(d^{s-1})+\sum_{r=s}^{t}(\gamma^{*})^{t-r}\times[\beta^{*}d_{r}+\theta_{r}^{*}+\nu_{ir}^{*}(d^{r})].$ (52)
 Proof of Lemma 2.
@@
 Now suppose (52) holds for some index $t-1\geq s$:
-$Y_{it-1}(d^{t-1})=(\gamma^{*})^{t-s}Y_{is-1}(d^{s-1})+\sum_{r=s}^{t-1}(\gamma^{*})^{t-1-r}\times[\beta^{*}d_{r}+\theta_{r}^{*}+\nu_{ir}^{*}(d_{r})]$
+$Y_{it-1}(d^{t-1})=(\gamma^{*})^{t-s}Y_{is-1}(d^{s-1})+\sum_{r=s}^{t-1}(\gamma^{*})^{t-1-r}\times[\beta^{*}d_{r}+\theta_{r}^{*}+\nu_{ir}^{*}(d^{r})]$
 It suffices to show the formula also holds for $t$. For $t$, the linear potential outcome model states:
-$Y_{it}(d^{t})=\gamma^{*}Y_{it-1}(d^{t-1})+\beta^{*}d_{t}+\theta_{t}^{*}+\nu_{it}^{*}(d_{t})$
+$Y_{it}(d^{t})=\gamma^{*}Y_{it-1}(d^{t-1})+\beta^{*}d_{t}+\theta_{t}^{*}+\nu_{it}^{*}(d^{t})$
 Plugging in from the preceding equation:
-$Y_{it}(d^{t})=$ $\gamma^{*}\left[(\gamma^{*})^{t-s}Y_{is-1}(d^{s-1})+\sum_{r=s}^{t-1}(\gamma^{*})^{t-1-r}\times[\beta^{*}d_{r}+\theta_{r}^{*}+\nu_{ir}^{*}(d_{r})]\right]$
-$+\beta^{*}d_{t}+\theta_{t}^{*}+\nu_{it}^{*}(d_{t})$
+$Y_{it}(d^{t})=$ $\gamma^{*}\left[(\gamma^{*})^{t-s}Y_{is-1}(d^{s-1})+\sum_{r=s}^{t-1}(\gamma^{*})^{t-1-r}\times[\beta^{*}d_{r}+\theta_{r}^{*}+\nu_{ir}^{*}(d^{r})]\right]$
+$+\beta^{*}d_{t}+\theta_{t}^{*}+\nu_{it}^{*}(d^{t})$
 Collecting terms yields the desired result.
@@
 $=$ $(\gamma^{*})^{t-s+1}Y_{is-1}(d^{s-1})+$
-$\sum_{r=s}^{t}(\gamma^{*})^{t-r}\times[\beta^{*}d_{r}+\theta_{r}^{*}+\alpha_{i}^{*}+E[\varepsilon_{ir}^{*}(d_{r})|\alpha_{i}^{*},Y_{i}^{s-1}(d^{s-1}),D_{i}^{s-1}=d^{s-1},D_{is}=d_{s}^{\prime}]]$
+$\sum_{r=s}^{t}(\gamma^{*})^{t-r}\times[\beta^{*}d_{r}+\theta_{r}^{*}+\alpha_{i}^{*}+E[\varepsilon_{ir}^{*}(d^{r})|\alpha_{i}^{*},Y_{i}^{s-1}(d^{s-1}),D_{i}^{s-1}=d^{s-1},D_{is}=d_{s}^{\prime}]]$
 where the first equality follows by substituting realized treatment, and the second by (52) of Lemma 2. It therefore suffices to show that:
-$E[\varepsilon_{it}^{*}(d_{t})|\alpha_{i}^{*},Y_{i}^{s-1}(d^{s-1}),D_{i}^{s-1}=d^{s-1},D_{is}=d_{s}^{\prime}]$
+$E[\varepsilon_{it}^{*}(d^{t})|\alpha_{i}^{*},Y_{i}^{s-1}(d^{s-1}),D_{i}^{s-1}=d^{s-1},D_{is}=d_{s}^{\prime}]$
 is invariant in $d_{s}^{\prime}$, or a fortiori that the expectation equals zero, for $t\geq s$. Lemma 2 implies that the preceding expression is equal to:
-$E[\varepsilon_{it}^{*}(d_{t})|\alpha_{i}^{*},Y_{i0},D_{i}^{s-1}=d^{s-1},D_{is}=d_{s}^{\prime},\nu_{i}^{*s-1}(d^{s-1})]$
+$E[\varepsilon_{it}^{*}(d^{t})|\alpha_{i}^{*},Y_{i0},D_{i}^{s-1}=d^{s-1},D_{is}=d_{s}^{\prime},\nu_{i}^{*s-1}(d^{s-1})]$
 because $\nu_{i}^{*s-1}(d^{s-1})$ can be recovered from $Y_{i}^{s-1}(d^{s-1})$ by (51), and conversely $Y_{i}^{s-1}(d^{s-1})$ can be recovered from $(Y_{i0},\nu_{i}^{*s-1}(d^{s-1}))$ by (52).
 Since $\varepsilon_{i}^{*s-1}(d^{s-1})=\alpha_{i}^{*}-\nu_{i}^{*s-1}(d^{s-1})$, this is equivalent to:
-$E[\nu_{it}^{*}(d_{t})|\alpha_{i}^{*},Y_{i0},D_{i}^{s-1}=d^{s-1},D_{is}=d_{s}^{\prime},\varepsilon_{i}^{*s-1}(d^{s-1})]$
+$E[\varepsilon_{it}^{*}(d^{t})|\alpha_{i}^{*},Y_{i0},D_{i}^{s-1}=d^{s-1},D_{is}=d_{s}^{\prime},\varepsilon_{i}^{*s-1}(d^{s-1})]$
 Finally, iterating expectations in the preceding expression and invoking (18) yields the desired result:
-$E[E[\varepsilon_{it}^{*}(d_{t})|\alpha_{i}^{*},Y_{i0},D_{i}^{t},\varepsilon^{*t-1}(\cdot)]|\alpha_{i}^{*},Y_{i0},D_{i}^{s-1}=d^{s-1},D_{is}=d_{s}^{\prime},\varepsilon^{*s-1}(d^{s-1})]=0.$
+$E[E[\varepsilon_{it}^{*}(d^{t})|\alpha_{i}^{*},Y_{i0},D_{i}^{t},\varepsilon^{*t-1}(\cdot)]|\alpha_{i}^{*},Y_{i0},D_{i}^{s-1}=d^{s-1},D_{is}=d_{s}^{\prime},\varepsilon^{*s-1}(d^{s-1})]=0.$
 Proof of Proposition 5.
-Fix a sequence of treatment $d^{T}$. For simplicity of notation, expand a vector-valued representation of choice rule (28) as:
-$D^{t}=\phi^{t}(D^{t-1},Y^{t-1}(D^{t-1}),\zeta^{t}(D^{t-1}))$
-so that we can express the event:
-$\{D^{t}=d^{t}\}=\{\phi^{t}(d^{t-1},Y^{t-1}(d^{t-1}),\zeta^{t}(d^{t-1}))=d^{t}\}$ (54)
-Condition (29) implies:
-$Y^{t}(d^{t})\perp\zeta^{t}(d^{t-1})\mid Y_{0}$
-because $Y^{t}(d^{t})$ is a function of $(Y_{0},\alpha(\cdot),\varepsilon^{t}(d^{t}))$ by definition (27). By rules of conditional independence,
-$(Y_{t}(d^{t}),\dots,Y_{T}(d^{T}))\perp\zeta^{t}(d^{t-1})\mid Y^{t-1}(d^{t-1})$
-and furthermore, for $t\geq 2$,
-$(Y_{t}(d^{t}),\dots,Y_{T}(d^{T}))\perp\zeta^{t}(d^{t-1})\mid Y^{t-1}(d^{t-1}),\,\phi^{t-1}(d^{t-2},Y^{t-2}(d^{t-2}),\zeta^{t-1}(d^{t-2}))=d^{t-1}$
-because, having conditioned on $Y^{t-1}(d^{t-1})$, the second conditioning statement is a function of $\zeta^{t-1}(d^{t-2})$.
-By (54) and because $\zeta^{t}(D^{t-1})=\zeta^{t}$ and $Y^{t-1}(D^{t-1})=Y^{t-1}$, this reduces to:
-$(Y_{t}(d^{t}),\dots,Y_{T}(d^{T}))\perp\zeta^{t}\mid Y^{t-1},\,D^{t-1}=d^{t-1}$
-Conditioning on history $(D^{t-1},Y^{t-1})$, this implies that, for a deterministic function $\phi_{t}(\cdot)$:
-$(Y_{t}(d^{t}),\dots,Y_{T}(d^{T}))\perp\phi_{t}(D^{t-1},Y^{t-1},\zeta^{t})\mid Y^{t-1},\,D^{t-1}=d^{t-1}$
-which by definition (28) yields the desired result:
-$(Y_{t}(d^{t}),\dots,Y_{T}(d^{T}))\perp D_{t}|Y^{t-1},D^{t-1}=d^{t-1}$
-In the case where $t=1$, the conditioning event collapses to $Y_{0}$ and the argument proceeds analogously.
-Proof of Proposition 7.
-Assumption 2b holds under (33) by the same argument as in the proof of Proposition 6.
-We adapt it here for the weaker case of mean sequential exchangeability in simpler notation.
-With potential outcomes and treatment rules in (30) and (31), we have:
+Per Observation 1, when the DGP of potential outcomes satisfies Assumption 5, the implied sequence of observables $(D_{i}^{t},Y_{i}^{t})$ satisfies the reduced-form model of observed outcomes in (3) with $(\gamma,\beta,\theta_{t},\alpha_{i})=(\gamma^{*},\beta^{*},\theta_{t}^{*},\alpha_{i}^{*})$ and reduced-form errors $\varepsilon_{it}=\varepsilon_{it}^{*}(D_{i}^{t})$.
+First-differencing of (3) yields
+$\Delta Y_{it}=\beta^{*}\Delta D_{it}+\gamma^{*}\Delta Y_{it-1}+\Delta\theta_{t}^{*}+\Delta\varepsilon_{it},$
+where $E(\varepsilon_{it}|D_{i}^{t},Y_{i}^{t-1})=0$ because of the condition (18) in Assumption 5 and the fact that $Y_{i}^{t-1}$ is a function of $\{D_{i}^{t-1},Y_{i0},\varepsilon_{i}^{*t-1}(\cdot),\alpha_{i}\}$.
+Consequently, $E(\Delta\varepsilon_{it}|D_{i}^{t-1},Y_{i}^{t-2})=0$ for all $t\geq 2$.
+Using $E(\Delta D_{it}|D_{i}^{t-1},Y_{i}^{t-2})$ and $E(\Delta Y_{it-1}|D_{i}^{t-1},Y_{i}^{t-2})$ as instruments for $\Delta D_{it}$ and $\Delta Y_{it-1}$ in first-differenced instrumental-variable regression proves the result.
+Proof of Proposition 6. Under Assumption 6, we use the first-differenced observed equation (23) to write the conditional mean of $\Delta Y_{it}$ given $W_{it-1},\mathcal{H}_{t-1}(y_{0})$ as:
+$E[\Delta Y_{it}|W_{it-1},\mathcal{H}_{t-1}(y_{0})]=E[\tau^{*}_{it}D_{it}|W_{it-1},\mathcal{H}_{t-1}(y_{0})]+\gamma^{*}E[\Delta Y_{it-1}|W_{it-1},\mathcal{H}_{t-1}(y_{0})]+\Delta\theta_{t}^{*},$ (53)
+because $E(\Delta\varepsilon^{*}_{it}|H_{it-1})=0$ and $W_{it-1}$ is a function of $H_{it-1}$. Moreover,
+$E[\tau^{*}_{it}D_{it}|W_{it-1},\mathcal{H}_{t-1}(y_{0})]$ $=$ $E[\tau^{*}_{it}(D_{i}^{t-1})|D_{it}=1,W_{it-1},\mathcal{H}_{t-1}(y_{0})]\times E[D_{it}|W_{it-1},\mathcal{H}_{t-1}(y_{0})]$
+$=$ $\underset{ATFT_{t}(y_{0})}{\underbrace{E[\tau^{*}_{it}(D_{i}^{t-1})|D_{it}=1,\mathcal{H}_{t-1}(y_{0})]}}\times E[\Delta D_{it}|W_{it-1},\mathcal{H}_{t-1}(y_{0})],$
+where the first equality holds by the law of iterated expectation, and the second holds under Assumption 7 because $W_{it-1}$ is a function of $\{\varepsilon_{i}^{*t-1}(\cdot),\alpha^{*}_{i}\}$ and $D_{i}^{t-1}=0$ once conditional on $\mathcal{H}_{t-1}(y_{0})$.
+This implies the mean of $\Delta Y_{it}$ conditional on $W_{it-1}$ and the event $\mathcal{H}_{t-1}(y_{0})$ is linear in $E[Z_{it}^{\prime}(y_{0})|\mathcal{H}_{t-1}(y_{0})]$, with the coefficients being $(ATFT_{t}(y_{0}),\gamma^{*},\Delta\theta_{t}^{*})$.
+Therefore,
+$(ATFT_{t}(y_{0}),\gamma^{*},\Delta\theta_{t}^{*})^{\prime}$ $=$ $E[Z_{it}^{\prime}Z_{it}|\mathcal{H}_{t-1}(y_{0})]^{-1}E[Z_{it}^{\prime}\Delta Y_{it}|\mathcal{H}_{t-1}(y_{0})]$
+$=$ $E[Z_{it}^{\prime}X_{it}|\mathcal{H}_{t-1}(y_{0})]^{-1}E[Z_{it}^{\prime}\Delta Y_{it}|\mathcal{H}_{t-1}(y_{0})],$
+where we suppress the dependence of instruments $Z_{it}(y_{0})$ on the initial condition $y_{0}$ for simplicity, and the second equality holds because $E[Z_{it}^{\prime}X_{it}|\mathcal{H}_{t-1}(y_{0})]=E[Z_{it}^{\prime}Z_{it}|\mathcal{H}_{t-1}(y_{0})]$ by the law of iterated expectation.
+Proof of Proposition 8. With potential outcomes and treatment rules in (31) and (32), we have:
 $E[Y_{2}(d_{1},d_{2})|D_{2}=d_{2},Y_{1}=y_{1},D_{1}=d_{1},Y_{0}=y_{0}]$
-$=$ $E\left[\lambda_{2}(d_{1},d_{2},y_{0},y_{1},\alpha(d_{2}),\varepsilon_{2}(d_{2}))\left|\begin{array}[]{c}\phi_{2}(y_{0},y_{1},d_{1},\xi^{1}(d_{1}),\eta_{2}(\cdot))=d_{2},\\ \lambda_{1}(d_{1},y_{0},\alpha(d_{1}),\varepsilon_{1}(d_{1}))=y_{1},\\ \phi_{1}(y_{0},\xi_{0},\eta_{1}(\cdot))=d_{1},\text{ \ }Y_{0}=y_{0}\end{array}\right.\right]$
-$=$ $E[\lambda_{2}(d_{1},d_{2},y_{0},y_{1},\alpha(d_{2}),\varepsilon_{2}(d_{2}))|\lambda_{1}(d_{1},y_{0},\alpha(d_{1}),\varepsilon_{1}(d_{1}))=y_{1},Y_{0}=y_{0}]\text{,}$
+[…]
+$=$ $E[\lambda_{2}(d_{1},d_{2},y_{0},y_{1},\alpha(d_{2}),\varepsilon_{2})|\lambda_{1}(d_{1},y_{0},\alpha(d_{1}),\varepsilon_{1})=y_{1},Y_{0}=y_{0}]\text{,}$
 where the last equality is due to (35) and the right-hand side does not vary with $d_{2}$.
@@
 $E[Y_{2}(d_{1},d_{2})|D_{1}=d_{1},Y_{0}]$
-$=$ $E[\lambda_{2}(0,0,y_{0},Y_{1}(d_{1}),\alpha(d_{2}),\varepsilon_{2}(d_{2}))|\phi_{1}(y_{0},\xi_{0},\eta_{1}(\cdot))=d_{1},Y_{0}=y_{0}]$
-$=$ $E[\lambda_{2}(0,0,y_{0},Y_{1}(d_{1}),\alpha(d_{2}),\varepsilon_{2}(d_{2}))|Y_{0}=y_{0}]\text{,}$
-where the last equality holds because of (33) and the fact that $Y_{1}(d_{1})$ is a function of $(Y_{0},\alpha(d_{1}),\varepsilon_{1}(d_{1}))$. Again, the right-hand side does not vary with $d_{1}$.
+$=$ $E[\lambda_{2}(0,0,y_{0},Y_{1}(d_{1}),\alpha(d_{2}),\varepsilon_{2})|\phi_{1}(y_{0},\xi_{0},\eta_{1})=d_{1},Y_{0}=y_{0}]$
+$=$ $E[\lambda_{2}(0,0,y_{0},Y_{1}(d_{1}),\alpha(d_{2}),\varepsilon_{2}|Y_{0}=y_{0}]\text{,}$
+where the last equality holds because of (35) and the fact that $Y_{1}(d_{1})$ is a function of $(Y_{0},\alpha(d_{1}),\varepsilon_{1})$. Again, the right-hand side does not vary with $d_{1}$.
 Furthermore,
 $E[Y_{1}(d_{1})|D_{1}=d_{1},Y_{0}]$
-$=$ $E[\lambda_{1}(d_{1},y_{0},\alpha(d_{1}),\varepsilon_{1}(d_{1}))|\phi_{1}(y_{0},\xi_{0},\eta_{1}(\cdot))=d_{1},Y_{0}=y_{0}]$
-$=$ $E[\lambda_{1}(d_{1},y_{0},\alpha(d_{1}),\varepsilon_{1}(d_{1}))|Y_{0}=y_{0}]\text{,}$
+$=$ $E[\lambda_{1}(d_{1},y_{0},\alpha(d_{1}),\varepsilon_{1})|\phi_{1}(y_{0},\xi_{0},\eta_{1})=d_{1},Y_{0}=y_{0}]$
+$=$ $E[\lambda_{1}(d_{1},y_{0},\alpha(d_{1}),\varepsilon_{1})|Y_{0}=y_{0}]\text{,}$
 where the last equality is due to (35) and the right-hand side does not vary with $d_{1}$. Thus, mean sequential exogeneity in Assumption 2b holds.
@@
 $E[Y_{2}(0,0)-Y_{1}(0)|(D_{1},D_{2})=(d_{1},d_{2}),Y_{0}=y_{0}]$
-$=$ $E\left[\lambda_{2}(0,0,y_{0},Y_{1}(0),\alpha(0),\varepsilon_{2}(0))-Y_{1}(0)\left|\begin{array}[]{c}\phi_{2}(y_{0},Y_{1}(d_{1}),d_{1},\xi^{1}(d_{1}),\eta_{2}(\cdot))=d_{2},\\ \phi_{1}(Y_{0},\xi_{0},\eta_{1}(\cdot))=d_{1},Y_{0}=y_{0}\end{array}\right.\right]$
-where $Y_{2}(0,0)-Y_{1}(0)$ is only a deterministic function of $(Y_{0},\alpha(0),\varepsilon_{1}(0),\varepsilon_{2}(0))$.
-Under (33), the right-hand side does generally vary with $(d_{1},d_{2})$ even after controlling for $Y_{0}$.
-This is because the event “$(D_{1},D_{2})=(d_{1},d_{2})$” does restrict the joint support of $(Y_{0},\alpha(d_{1}),\varepsilon_{1}(d_{1}),\xi^{1}(d_{1}),\eta^{2}(\cdot))$.
+[…]
+where $Y_{2}(0,0)-Y_{1}(0)$ is a deterministic function of $(Y_{0},\alpha(0),\varepsilon_{1},\varepsilon_{2})$.
+Under (35), the right-hand side does generally vary with $(d_{1},d_{2})$ even conditional on $Y_{0}$.
+This is because the event $(D_{1},D_{2})=(d_{1},d_{2})$” does restrict the joint support of $(Y_{0},\alpha(d_{1}),\varepsilon_{1},\xi_{1}(d_{1}),\eta^{2})$.
 [^12] Note $Y_{1}(d_{1})=\lambda_{1}(d_{1},Y_{0},\alpha(d_{1}),\varepsilon_{1}(d_{1}))$ affects the definition of conditioning events due to its entry through $\phi_{2}(\cdot)$.
 Hence, even after conditioning on $Y_{0}=y_{0}$, the right-hand side is in general non-degenerate in $(d_{1},d_{2})$.
-Appendix B: Consistency of Adjusted IPW Estimator
-In this appendix, we establish the consistency of the adjusted IPW estimator $\widehat{\mu}_{\tau_{2}}$, which is defined in (20). Maintain the following conditions:
-[…]
+Appendix B Generalizations
+B.1 Identification of Treatment Effects for ARUPO
+In this appendix we extend the definition and identification of treatment effects to general sequences of past treatments $d^{t-1}$. That is, for any $d^{t-1}\in\{0,1\}^{t-1}$, generalize (25) to:
+$ATT_{t}(d^{t-1},y_{0})\equiv E[\tau_{it}^{*}(d^{t-1})|D_{it}=1,\mathcal{H}_{t-1}(d^{t-1},y_{0})]$
+where $\mathcal{H}_{t-1}(d^{t-1},y_{0})\equiv\{D_{i}^{t-1}=d^{t-1},Y_{i0}=y_{0}\}$. We maintain the ARUPO model (Assumption 6), but strengthen Assumption 7 on treatment effects as follows.
+Assumption 8
+For any initial condition $y_{0}$ and $t\geq 1$,
+$E[\tau_{it}^{*}(d^{t-1})|D_{it},\mathcal{H}_{t-1}(d^{t-1},y_{0}),\tau_{i}^{*t-1}(d^{t-2}),\alpha_{i}^{*},\varepsilon_{i}^{*t-1}]=E[\tau_{it}(d^{t-1})|D_{it},\mathcal{H}_{t-1}(d^{t-1},y_{0})],$
+where $d^{t-1}\equiv(d^{t-2},d_{t-1})$.
+Assumption 8 resembles Assumption 7 except that it also posits the (conditional) serial uncorrelation of past treatment effects. Under Assumption 8,
+$E[\tau_{it}(d^{t-1})|D_{it},D_{i}^{t-1}=d^{t-1},Y_{i}^{t-1}]=E[\tau_{it}(d^{t-1})|D_{it},D_{i}^{t-1}=d^{t-1},Y_{i0}],$ (56)
+because $Y_{i}^{t-1}$ is a function of $\{Y_{i0},\tau_{i}^{*t-1}(d^{t-2}),\varepsilon_{i}^{*t-1},\alpha_{i}^{*}\}$ conditional on $D_{i}^{t-1}=d^{t-1}$.
+Extending our IV estimand to the generalized history, we (re)define:
+$Z_{it}(d^{t-1},y_{0})\equiv E[X_{it}|W_{it-1},\mathcal{H}_{t-1}(d^{t-1},y_{0})]$
 and
-[…]
-Condition (57) requires uniform convergence of preliminary estimators $\widehat{M}_{1i},\widehat{M}_{2i}$; it holds for many nonparametric (e.g.
-Nadaraya-Watson) estimators of conditional means under standard conditions.
-Condition (58) include standard, mild requirement that the sample moments of $M_{1i},M_{2i}$ are well-behaved.
-Corollary B1
-[…]
-Proof of Corollary B1. By the Mean Value Theorem,
-$\hat{\mu}_{\tau_{2}}=\frac{1}{n}\sum\nolimits_{i}\left[\frac{\Delta Y_{i2}-M_{2i}}{M_{1i}}+\left(\frac{-1}{\tilde{M}_{1i}}\right)\left(\hat{M}_{2i}-M_{2i}\right)+\left(\frac{\tilde{M}_{2i}}{\tilde{M}_{1i}^{2}}\right)(\hat{M}_{1i}-M_{1i})\right]\text{,}$
-where $(\Delta Y_{i2}-M_{2i})/M_{1i}$ is an “infeasible” summand that involves the actual population moments $M_{1i},M_{2i}$, and $\tilde{M}_{ki}$ are intermediate values bounded between $\hat{M}_{ki}$ and $M_{ki}$ for $k=1,2$.
-By the triangular inequality,
-$\left|\hat{\mu}_{\tau_{2}}-\frac{1}{n}\left(\sum\nolimits_{i}\frac{\Delta Y_{i2}-M_{2i}}{M_{1i}}\right)\right|$
-$\leq$ $\frac{1}{n}\sum\nolimits_{i}\left|\tilde{M}_{1i}\right|^{-1}\left|\hat{M}_{2i}-M_{2i}\right|+\frac{1}{n}\sum\nolimits_{i}\left|\tilde{M}_{2i}\right|\left|\tilde{M}_{1i}\right|^{-2}\left|\hat{M}_{2i}-M_{2i}\right|$
-$\leq$ $\left(\frac{1}{n}\sum\nolimits_{i}\left|\tilde{M}_{1i}\right|^{-1}\right)\left(\sup_{1\leq i^{\prime}\leq n}\left|\hat{M}_{2i^{\prime}}-M_{2i^{\prime}}\right|\right)+\left(\frac{1}{n}\sum\nolimits_{i}\left|\tilde{M}_{2i}\right|\left|\tilde{M}_{1i}\right|^{-2}\right)\left(\sup_{1\leq i^{\prime}\leq n}\left|\hat{M}_{1i}-M_{1i}\right|\right)\text{.}$
-By (57), the two suprema of differences are $o_{p}(1)$, i.e., converge in probability to 00.
-Next, we establish that $\frac{1}{n}\sum\nolimits_{i}\left|\tilde{M}_{1i}\right|^{-1}=O_{p}(1)$, i.e.
-is bounded in probability.
-For this, we want to show “$\forall\varepsilon>0$, $\exists\mathcal{C}_{\varepsilon}>0$ such that $\Pr\{\frac{1}{n}\sum\nolimits_{i}\left|\tilde{M}_{1i}\right|^{-1}>\mathcal{C}_{\varepsilon}\}<\varepsilon$ for $n$ large enough.” By construction,
-$\Pr\{\frac{1}{n}\sum\nolimits_{i}\left|\tilde{M}_{1i}\right|^{-1}>\mathcal{C}_{\varepsilon}\}\leq\Pr\left\{\max\nolimits_{i\leq n}\left|\tilde{M}_{1i}\right|^{-1}>\mathcal{C}_{\varepsilon}\right\}$ (59)
-$=$ $\Pr\left\{\min\nolimits_{i\leq n}\left|\tilde{M}_{1i}\right|<1/\mathcal{C}_{\varepsilon}\right\}\leq\Pr\left\{\min\nolimits_{i\leq n}M_{1i}-\tilde{e}_{1n}<1/\mathcal{C}_{\varepsilon}\right\}$
-$=$ $\Pr\left\{\tilde{e}_{1n}>\min\nolimits_{i\leq n}M_{1i}-1/\mathcal{C}_{\varepsilon}\right\}\text{,}$
-where $\tilde{e}_{kn}\equiv\sup_{i\leq n}\left|\tilde{M}_{ki}-M_{ki}\right|$ for $k=1,2$, and the second inequality is due to triangular inequality and positivity of $M_{i}$.
-By condition (i) in (58), $\min_{i}M_{i}-1/\mathcal{C}_{\varepsilon}$ is strictly positive when $\mathcal{C}_{\varepsilon}$ is large enough.
-The uniform convergence of $\hat{M}_{1i}$ in (57) then implies the right-hand side of (59) converges to zero as $n\rightarrow\infty$.
-This proves $\frac{1}{n}\sum\nolimits_{i}\left|\tilde{M}_{1i}\right|^{-1}=O_{p}(1)$.
-Finally, we establish $\frac{1}{n}\sum\nolimits_{i}\left|\tilde{M}_{2i}\right|\left|\tilde{M}_{1i}\right|^{-2}$ is bounded in probability. Note:
-$\frac{1}{n}\sum\nolimits_{i}\left|\tilde{M}_{2i}\right|\left|\tilde{M}_{1i}\right|^{-2}\leq\left(\frac{1}{n}\sum\nolimits_{i}\left|M_{2i}\right|+\tilde{e}_{2n}\right)\left(\max_{i^{\prime}}\left|\tilde{M}_{1i^{\prime}}\right|^{-2}\right)\text{,}$
-due the triangular inequality and definition of $\tilde{e}_{2n}$.
-The second term on the right-hand side is $O_{p}(1)$ by the same argument above (with $\mathcal{C}_{\varepsilon}$ replaced by $\mathcal{C}_{\varepsilon}^{2}$).
-By condition (ii) in (58), the stochastic order of the first term on the right-hand side is $O_{p}(1)+o_{p}(1)=O_{p}(1)$.
-Thus the product of the two terms, and consequently $\frac{1}{n}\sum\nolimits_{i}\left|\tilde{M}_{2i}\right|\left|\tilde{M}_{1i}\right|^{-2}$, is bounded in probability.
+$\begin{pmatrix}\widetilde{\beta}(d^{t-1},y_{0})\\ \widetilde{\gamma}(d^{t-1},y_{0})\\ \widetilde{\Delta\theta_{t}}(d^{t-1},y_{0})\end{pmatrix}\equiv E[Z_{it}^{\prime}(d^{t-1},y_{0})X_{it}|\mathcal{H}_{t-1}(d^{t-1},y_{0})]^{-1}E[Z_{it}^{\prime}(d^{t-1},y_{0})\Delta Y_{it}|\mathcal{H}_{t-1}(d^{t-1},y_{0})].$ (57)
+The next result is analogous to Proposition 6 under Assumption 6 and 7, only with the conditioning event and the causal parameter defined as $\mathcal{H}_{t-1}(d^{t-1},y_{0})$ and $ATT_{t}(d^{t-1},y_{0})$.
+Proposition 9
+Suppose Assumptions 6 and 8 hold, and:
+$E[Z_{it}^{\prime}(d^{t-1},y_{0})Z_{it}(d^{t-1},y_{0})|\mathcal{H}_{t-1}(d^{t-1},y_{0})]$
+has full rank. Then the IV slope estimands (57) for $t\geq 3$ recover the causal terms:
+$\widetilde{\beta}(d^{t-1},y_{0})$ $=ATT_{t}(d^{t-1},y_{0});$
+$\widetilde{\gamma}(d^{t-1},y_{0})$ $=\gamma^{*},$
+and the intercept estimand recovers:
+$\widetilde{\Delta\theta_{t}}(d^{t-1},y_{0})=\begin{cases}\Delta\theta_{t}^{*}&\text{when $d_{t-1}=0$};\\ \Delta\theta_{t}^{*}+ATT_{t}(d^{t-1},y_{0})-ATT_{t-1}(d^{t-2},y_{0})&\text{when $d_{t-1}=1$.}\end{cases}$
+Proof of Proposition 9. Since this proof broadly resembles the proof of Proposition 9, we only note the points of departure and extension.
+First, consider the case with $d_{t-1}=0$ so that $\Delta D_{it}=D_{it}$ conditional on $\mathcal{H}_{t-1}(d^{t-1},y_{0})$.
+The proof of these analogous identification results in Proposition 6 remains mostly the same, with the following changes.
+First, it conditions on $\mathcal{H}_{t-1}(d^{t-1},y_{0})$ instead of the previous $\mathcal{H}_{t-1}(y_{0})$.
+Second, $W_{it-1}$ is now a function of $\{\varepsilon_{i}^{*t-1}(\cdot),\alpha_{i}^{*},\tau_{t-2}^{*}(d^{t-2})\}$ once conditional on $\mathcal{H}_{t-1}(d^{t-1},y_{0})$, which requires Assumption 8 to conclude that:
+$E[\tau_{it}^{*}(D_{i}^{t-1})|D_{it}=1,W_{it-1},\mathcal{H}_{t-1}(d^{t-1},y_{0})]=ATT_{t}(d^{t-1},y_{0}).$
+In the other case, $d_{t-1}=1$ so that $\Delta D_{it}=D_{it}-1$ conditional on $\mathcal{H}_{t-1}(d^{t-1},y_{0})$.
+Relative to Proposition 6, the conditional mean of the differenced outcome (analogous to (53)) now includes an additional term:
+$E[\tau_{it-1}^{*}(D_{i}^{t-2})|W_{it-1},\mathcal{H}_{t-1}(d^{t-1},y_{0})]=-ATT_{t-1}(d^{t-2},y_{0}),$
+where the equality holds under Assumption 8 because $W_{it-1}$ is a function of $\varepsilon_{i}^{*t-1}(\cdot)$, $\alpha_{i}^{*}$ and $\tau_{t-2}^{*}(d^{t-2})$ once conditional on $\mathcal{H}_{t-1}(d^{t-1},y_{0})$.
+Plugging in and rearranging yields the linear conditional expectation expression:
+$E[\Delta Y_{it}|W_{it-1},\mathcal{H}_{t-1}(d^{t-1},y_{0})]=$ $ATT_{t}(d^{t-1},y_{0})E[\Delta D_{it}|W_{it-1},\mathcal{H}_{t-1}(d^{t-1},y_{0})]+$
+$\gamma^{*}E[\Delta Y_{it-1}|W_{it-1},\mathcal{H}_{t-1}(y_{0})]+\Delta\theta_{t}^{*}+$
+$ATT_{t}(d^{t-1},y_{0})-ATT_{t-1}(d^{t-2},y_{0}).$
+The argument for interpreting the OLS estimand as an IV estimand follows as before.
+Note that in this case, the estimand for the constant term includes an additional intertemporal difference in treatment effects:
+$\widetilde{\Delta\theta_{t}}=\Delta\theta_{t}^{*}+ATT_{t}(d^{t-1},y_{0})-ATT_{t-1}(d^{t-2},y_{0}).$